### v1

# Constrained and Ordered Level Planarity Parameterized by the Number of Levels

Abstract
The problem Level Planarity asks for a crossing-free drawing of a graph in the plane such that vertices are placed at prescribed y-coordinates (called levels) and such that every edge is realized as a y-monotone curve. In the variant Constrained Level Planarity (CLP), each level $y$ is equipped with a partial order $\prec_{y}$ on its vertices and in the desired drawing the left-to-right order of vertices on level $y$ has to be a linear extension of $\prec_{y}$. Ordered Level Planarity (OLP) corresponds to the special case of CLP where the given partial orders $\prec_{y}$ are total orders. Previous results by Brückner and Rutter [SODA 2017] and Klemz and Rote [ACM Trans. Alg. 2019] state that both CLP and OLP are NP-hard even in severely restricted cases. In particular, they remain NP-hard even when restricted to instances whose width (the maximum number of vertices that may share a common level) is at most two. In this paper, we focus on the other dimension: we study the parameterized complexity of CLP and OLP with respect to the height (the number of levels).
We show that OLP parameterized by the height is complete with respect to the complexity class XNLP, which was first studied by Elberfeld, Stockhusen, and Tantau [Algorithmica 2015] (under a different name) and recently made more prominent by Bodlaender, Groenland, Nederlof, and Swennenhuis [FOCS 2021]. It contains all parameterized problems that can be solved nondeterministically in time $f(k)\cdot n^{\mathcal{O}(1)}$ and space $f(k)\cdot\log n$ (where $f$ is a computable function, $n$ is the input size, and $k$ is the parameter). If a problem is XNLP-complete, it lies in XP, but is W[$t$]-hard for every $t$.
In contrast to the fact that OLP parameterized by the height lies in XP, it turns out that CLP is NP-hard even when restricted to instances of height 4. We complement this result by showing that CLP can be solved in polynomial time for instances of height at most 3.
keywords:
Parameterized Complexity, Graph Drawing, XNLP, XP, W[t]-hard, Level Planarity, Planar Poset Diagram, Computational Geometry
1 Introduction
In an upward drawing of a directed graph, every edge $e=(u,v)$ is realized as a y-monotone curve that goes upwards from $u$ to $v$, i.e., the y-coordinate strictly increases when traversing $e$ from $u$ towards $v$. Also known as poset diagrams, these drawings provide a natural way to visualize a partial order on a set of items. The classical problem Upward Planarity asks whether a given directed graph admits a drawing that is both upward and planar (i.e., crossing-free). It is known to be NP-hard [17], but becomes solvable in polynomial time if the y-coordinate of each vertex is prescribed [10, 19, 22]. In contrast, when both the y-coordinate and the x-coordinate of each vertex is prescribed, the problem is yet again NP-hard [24]. The paper at hand is concerned with the parameterized complexity of (a generalization of) the latter variant of Upward Planarity, the parameter being the number of levels. Next, we define these problems more precisely, adopting the notation and terminology used in [24].
Level planarity.
A level graph $\mathcal{G}=(G,\gamma)$ is a directed graph $G=(V,E)$ together with a level assignment, which is a surjective map $\gamma\colon V\rightarrow\{1,2,\dots,h\}$ where $\gamma(u)<\gamma(v)$ for every edge $(u,v)\in E$. The vertex set $V_{i}=\{v\mid\gamma(v)=i\}$ is called the $i$-th level of $\mathcal{G}$. The width of level $V_{i}$ is $|V_{i}|$. The levelwidth of $\mathcal{G}$ is the maximum width of any level in $\mathcal{G}$ and the height of $\mathcal{G}$ is the number $h$ of levels. A level planar drawing of $\mathcal{G}$ is an upward planar drawing of $G$ where the y-coordinate of each vertex $v$ is $\gamma(v)$. We use $L_{i}$ to denote the horizontal line with y-coordinate $i$. Algorithms for computing level planar drawings usually just determine a level planar embedding of a level planar drawing, which for each $i~{}\in\{1,2,\dots,h\}$ lists the left-to-right sequence of vertices and edges intersected by $L_{i}$. Note that this corresponds to an equivalence class of drawings from which an actual drawing is easily derived. The level graph $\mathcal{G}$ is called proper if $\gamma(v)=\gamma(u)+1$ for every edge $(u,v)\in E$.
The problem Level Planarity asks whether a given level graph admits a level planar drawing. It can be solved in linear time [10, 19, 21, 22]; see [15] for a more detailed discussion on this series of papers. It is easy to see that Level Planarity is polynomial time/space equivalent to the variant where $\gamma$ maps to $h$ arbitrary distinct real numbers.
Constrained and ordered level planarity.
In 2017, Brückner and Rutter [7] and Klemz and Rote [24] independently introduced and studied two closely related variants of Level Planarity, defined as follows. A constrained (ordered) level graph $\mathcal{G}=(G,\gamma,(\prec_{i})_{1\leq i\leq h})$ is a triplet corresponding to a level graph $(G,\gamma)$ of height $h$ equipped with a family containing, for each $1\leq i\leq h$, a partial (total) order on the vertices in $V_{i}$. A constrained (ordered) level planar drawing of $\mathcal{G}$ is a level planar drawing of $(G,\gamma)$ where, for each $1\leq i\leq h$, the left-to-right order of the vertices in $V_{i}$ corresponds to a linear extension of $\prec_{i}$ (is $\prec_{i}$). For a pair of vertices $u,v\in V_{i}$ with $u\prec_{i}v$, we refer to $u\prec_{i}v$ as a constraint on $u$ and $v$.
The problem Constrained Level Planarity (CLP) / Ordered Level Planarity (OLP) asks whether a given constrained / ordered level graph admits a constrained / ordered level planar drawing, in which case the input is called a constrained / ordered level planar graph. The special case where the height of all instances is restricted to a given value $h$ is called $h$-level CLP / $h$-level OLP. In CLP, each partial order $\prec_{i}$ is assumed to be given in form of a directed acyclic graph including all of its transitive edges. In OLP, each total order $\prec_{i}$ is encoded by equipping each vertex of level $V_{i}$ with an integer that is equal to its rank in the order $\prec_{i}$. Note that OLP is polynomial time/space equivalent to the variant of Level Planarity where each vertex is equipped with a prescribed x-coordinate, implying that the only challenge is to draw the edges.
Klemz and Rote [24] showed that OLP (and, thus, CLP) is NP-hard even when restricted to the case where the underlying undirected graph of $G$ is a disjoint union of paths. Note that such a graph has bounded pathwidth (and treewidth), maximum degree, and feedback vertex set number, ruling out efficient parameterized algorithms with respect to these classical parameters. Additionally, the instances produced by their reduction have a levelwidth of only two. Independently, Brückner and Rutter [7] provided a very different reduction for showing that CLP is NP-hard (in fact, their reduction shows the NP-hardness of Partial Level Planarity, which can be seen as a generalization of OLP and a special case of CLP; see below). The instances constructed by their reduction are connected and have bounded maximum degree. They also present a polynomial time algorithm for CLP for the case where the graph has a single sink, which they later sped up [8].
Other related work.
The problem Partial Level Planarity (PLP) (introduced and studied by Brückner and Rutter [7]), asks whether a given level planar drawing of a subgraph $H$ of the input graph $G$ can be extended to a level planar drawing of $G$. This can be seen as a generalization of OLP and, in the proper case, as a specialization of CLP. Several other problems related to the construction of level planar drawings have been studied, including problems with other kinds of ordering constraints (e.g., Clustered Level Planarity [14, 1, 24] and T-Level Planarity [28, 1, 24]), problems with a more geometric touch (see, e.g., [20, 23]), and variants of Level Planarity seeking drawings on surfaces different from the plane (see, e.g., [3, 2, 4]).
Contribution and organization.
As stated above, Klemz and Rote [24] showed that OLP (and, thus, CLP) is NP-hard even when restricted to instances of levelwidth two. In this paper, we focus on the other “dimension”: we study CLP and OLP parameterized by height.
We show that OLP parameterized by the height is complete with respect to the complexity class XNLP, which was first studied by Elberfeld, Stockhusen, and Tantau [12] (under the name $\mathrm{N}[f\ \mathrm{poly},f\ \mathrm{log}]$) and recently made more prominent by Bodlaender, Groenland, Nederlof, and Swennenhuis [6]. It contains all parameterized problems that can be solved nondeterministically in time $f(k)\cdot n^{\mathcal{O}(1)}$ and space $f(k)\cdot\log n$ (where $f$ is a computable function, $n$ the input size, and $k$ the parameter). Elberfeld et al. and Bodlaender et al. study properties of (problems in) this class and provide several problems that are XNLP-complete. In particular, if a problem is XNLP-complete, it lies in XP, but is $W[t]$-hard for every $t$ [6].
Theorem 1.1.
Ordered Level Planarity parameterized by the height of the input graph is XNLP-complete (and, thus, it lies in XP, but is $W[t]$-hard for every $t$). XNLP-hardness holds even when restricted to the case where the input graph is connected. Moreover, there is a constructive XP-time algorithm for Ordered Level Planarity (w.r.t. the height).
Parameterizing OLP by height captures the “linear” nature of the solution – this is reminiscent of recent results by Bodlaender, Groenland, Jacob, Jaffke, and Lima [5] who established XNLP-completeness for several problems parameterized by linear width measures (e.g., Capacitated Dominating Set by pathwidth and Max Cut by linear cliquewidth). However, to the best of our knowledge, this is the first graph drawing (or computational geometry) problem shown to be XNLP-complete. The algorithms are described in Section 2, whereas the hardness is shown in Section 3. In contrast to the fact that OLP parameterized by the height lies in XP, it is not difficult to see that the socket/plug gadget described by Brückner and Rutter [7] can be utilized in the context of a reduction from 3-Partition to show that (PLP and, thus) CLP remains NP-hard even when restricted to instances of constant height. In fact, the unpublished full version of [7] features such a construction with a height of 7 [26]. Here, we present a reduction that is tailor-made for CLP, showing that it is NP-hard even when restricted to instances of height 4. We complement this result by showing that the (surprisingly challenging) case of instances with height at most three can be solved in polynomial time.
Theorem 1.2.
Constrained Level Planarity is NP-hard even when restricted to height 4, but instances of height at most 3 can be solved constructively in polynomial time.
We show the hardness in Appendix C, describe the algorithm in Section 4, and state open problems in Section 5. Proofs of statements marked with a (clickable) $\star$ are in the appendix.
Notation and conventions.
Given an integer $k>0$, we use $[k]$ as shorthand for $\{1,2,\dots,k\}$. Given a directed or undirected graph $G$, let $V(G)$ denote the vertex set of $G$, and let $E(G)$ denote the edge set of $G$. Recall that level graphs are directed with each edge $(u,v)$ pointing upwards, i.e., $\gamma(u)<\gamma(v)$. As a shorthand and when the direction is not important, we use both $uv$ and $vu$ to refer to a directed edge $(u,v)$.
2 An XP / XNLP Algorithm for Ordered Level Planarity
In this section, we show that Ordered Level Planarity is in XNLP (and thus in XP) when parameterized by the height $h$ of the input graph. Moreover, we show how to construct an ordered level planar drawing (if it exists) in XP-time. The main idea of our approach is to continuously sweep the plane with an unbounded y-monotone curve $s$ from left to right in a monotone fashion such that for each edge $(u,v)$, there is a point in time where $u$ and $v$ are consecutive vertices along $s$. When this happens, the edge can be drawn without introducing any crossings due to the fact that $s$ moves monotonically. To discretize this idea and turn it into an algorithm, we instead determine a sequence of unbounded y-monotone curves $S=(s_{1},s_{2},\dots,s_{z})$ that is sorted from left to right (i.e., no point of $s_{i+1}$ is to the left of $s_{i}$), has a length of $z\in\mathcal{O}(n)$, and contains for every edge $(u,v)$ a curve $s_{i}$ along which $u$ and $v$ are consecutive vertices. Now, given $S$, the desired drawing can be constructed in polynomial time; for an illustration see Fig. 1. Moreover, the sequence $S$ can be obtained in XP-time/space by exhaustively enumerating all possibilities or in XNLP-time/space by nondeterministic guessing. Let us proceed to formalize these ideas.
Gaps and positions.
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph and consider one of its levels $V_{i}$. Let $(v_{1},v_{2},\dots,v_{\lambda_{i}})$ be the linear order of $V_{i}$ corresponding to $\prec_{i}$. In an ordered level planar drawing of $\mathcal{G}$, the vertices $V_{i}$ divide the line $L_{i}$ into a sequence of open line-segments and rays, which we call the gaps of $L_{i}$. A position on $L_{i}$ is a gap of $L_{i}$ or a vertex of $V_{i}$. Each position on $L_{i}$ is encoded by an index in $P_{i}=\{0\}\cup[2\lambda_{i}]$: the index 00 represents the gap that precedes $v_{1}$; an odd index $p$ represents the vertex $v_{\lceil p/2\rceil}$; and an even index $p\neq 0$ represents the gap that succeeds $v_{\lceil p/2\rceil}$.
Separations.
A separation for $\mathcal{G}$ is an element of $P_{1}\times P_{2}\times\dots\times P_{h}$. Intuitively, a separation $s=(p_{1},p_{2},\dots,p_{h})$ represents the equivalence class of unbounded y-monotone curves that intersect line $L_{i}$ in position $p_{i}$ for each $1\leq i\leq h$; see Fig. 1. We say that a vertex $v\in V$ is on $s$ if $p_{\gamma(v)}$ represents $v$. Moreover, we say that $v$ is to the left (right) of $s$ if the index corresponding to the position of $v$ is strictly smaller (larger) than $p_{\gamma(v)}$. Consider two vertices $u,v\in V$ that are on $s$ and where $\gamma(u)<\gamma(v)$. We say that $u$ and $v$ are consecutive on $s$ if all of the indices $p_{\gamma(u)+1},p_{\gamma(u)+2},\dots,p_{\gamma(v)-1}$ represent gaps. In this case, $v$ is the successor of $u$ on $s$ and $u$ is the predecessor of $v$ on $s$. We say that $s$ uses an edge $e=(u,v)\in E$ if $u$ and $v$ are consecutive along $s$.
Sweeping sequences.
A sweeping sequence for $\mathcal{G}$ is a sequence $S=(s_{1},s_{2},\dots,s_{z})$ of separations for $\mathcal{G}$ such that for each $j\in[z-1]$, we have $s_{j}\leq s_{j+1}$, componentwise. We say that a sweeping sequences uses an edge $e\in E$, if it contains a separation that uses $e$.
[…]
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph and let $S=(s_{1},s_{2},\dots,s_{z})$ be a sweeping sequence for $\mathcal{G}$ that uses every edge of $E$. Given $\mathcal{G}$ and $S$, an ordered level planar drawing of $\mathcal{G}$ can be constructed in $\mathcal{O}(zh+n^{2})$ time, where $n$ is the number of vertices.
Let $S=(s_{1},s_{2},\dots,s_{z})$ be a sweeping sequence for our ordered level graph $\mathcal{G}$. We say that $S$ is nice if for each $i\in[z-1]$, the separation $s_{i+1}$ is obtained from $s_{i}$ by incrementing exactly one component by $1$. Moreover, we say that $S$ is exhaustive if it is nice and $s_{1}=(0,0,\dots,0)$ and $s_{z}=(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1)$; see Fig. 1.
[…]
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph. Then $\mathcal{G}$ admits an ordered level planar drawing if and only if there is an exhaustive sweeping sequence for $G$ that uses every edge in $E$.
Proof 2.3 (Proof (sketch)).
The “if”-direction follows from Lemma 2.1. For the “only if”-direction, we show that the drawing can be internally triangulated by inserting y-monotone edges (allowing the use of parallel edges). For each y-monotone path $P$ that spans all levels, there is a nice sweeping sequence from $s_{1}=(0,\dots,0)$ to the separation corresponding to $P$ that uses all edges left of $P$. This is shown by arguing inductively on the number of triangles left of $P$.
We remark that an exhaustive sweeping sequence using all edges corresponds directly to a particularly well-structured path decomposition. Thus, Lemma 2.2 implies that every (ordered) level planar drawing of height $h$ represents a graph of pathwidth at most $h-1$; a statement that was independently proven in [11]. However, the path decompositions constructed in the proof of [11, Lemma 1] do not exhibit the same properties that are inherent to exhaustive sweeping sequences and on which our algorithms heavily rely. In particular, these path decompositions may contain bags with multiple vertices of a given level (unless the drawing is proper). Moreover, the existence of a path decomposition of width at most $h-1$ for an ordered level graph $\mathcal{G}$ of height $h$ does not characterize the fact that $\mathcal{G}$ is ordered level planar (recall that, in fact, OLP is NP-hard even when restricted to instances of pathwidth $1$ [24]).
Figure 1: An exhaustive sweeping sequence using all edges of the depicted graph and the corresponding (cf. Lemma 2.1) ordered level planar drawing, as well as the corresponding sequence of true dynamic programming table entries $T[s,U]$ (cf. Lemma 2.4). Note that in any ordered level planar drawing of the graph, exactly one of the edges $(2,4),(1,3)$ is located to the left of the path $(1,2,3,4)$, while the other is located to the right. Similarly, in any exhaustive sweeping sequence containing separation $s_{5}$, exactly one of the edges $(2,4),(1,3)$ is used by a separation preceding $s_{5}$, while the other is used by a separation succeeding $s_{5}$. Hence, when iteratively building an exhaustive sweeping sequence, it is key to remember which edges between vertices of the current separation have already been used – this is exactly the purpose of the sets $U$. E.g., the fact that $(2,4)\in U_{5}$ corresponds to $(2,4)$ being used before $s_{5}$ (in $s_{4}$), from which one can infer how to proceed.
[…]
There is an algorithm that determines whether a given ordered level graph admits an exhaustive sweeping sequence using all of its edges. It can be implemented deterministically using $\mathcal{O}^{*}(2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1))\subseteq\mathcal{O}^{*}(2^{\binom{h}{2}}(2\lambda+1)^{h})\subseteq\mathcal{O}^{*}(2^{\binom{h}{2}}(2n+1)^{h})$ time and space, or nondeterministically using polynomial time and $\mathcal{O}(h^{2}+h\log I)$ space, where $\lambda$ and $h$ denote the width and height of the input graph, respectively, $n$ denotes the number of vertices, $\lambda_{j}$ denotes the width of level $j\in[h]$, and $I$ denotes the input size. Further, the deterministic version can report the sequence (if it exists).
Proof 2.5 (Proof (sketch)).
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph and let $s=(p_{1},p_{2},\dots,p_{h})$ be a separation for $\mathcal{G}$. Further, let $U\subseteq E$ be a subset of the edges that are joining pairs of (not necessarily consecutive) vertices on $s$. We define $T[s,U]=\texttt{true}$ if there exists a nice sweeping sequence for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s$, and uses all edges in $U$, as well as all edges in $E$ incident to at least one vertex to the left of $s$. Otherwise, $T[s,U]=\texttt{false}$. Additionally, we allow $U=\bot$, in which case $T[s,U]=\texttt{false}$. Fig. 1 illustrates several true table entries $T[s_{i},U_{i}]$ along with the corresponding sweeping sequences $s_{1},s_{2},\dots,s_{i}$. Our goal is to determine $T[(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1),\emptyset]$, which is true if and only if there exists an exhaustive sweeping sequence for $\mathcal{G}$ using all edges in $E$.
We will determine the entries $T[s,U]$ by means of a dynamic programming recurrence. For the base case, we simply set $T[(0,0,\dots,0),\emptyset]=\texttt{true}$. Now assume that $s\neq(0,0,\dots,0)$. For each index $1\leq j\leq h$ where $p_{j}\geq 1$, we define a separation
$s_{j}^{\prime}=(p_{1},p_{2},\dots,p_{j-1},p_{j}-1,p_{j+1},\dots,p_{h}).$
Further, we define an edge set $U_{j}^{\prime}\subseteq E$ as follows:
  • If $p_{j}$ represents a vertex $v$, then
    • $U_{j}^{\prime}=\bot$ if $U$ contains edges joining $v$ with vertices on $s$ that are not its predecessor or successor on $s$;
    • otherwise $U_{j}^{\prime}$ is created from $U$ by removing the (up to two) edges incident to $v$.
  • If $p_{j}$ represents a gap and, thus, $p_{j-1}$ represents a vertex $v$, then
    • $U_{j}^{\prime}=\bot$ if $v$ is adjacent to a vertex to the right of $s$;
    • otherwise $U_{j}^{\prime}$ is created from $U$ by
        removing the edge between the predecessor and successor of $v$ along $s_{j}^{\prime}$ (if it exists and is contained in $U$); and
        adding all edges in $E$ that join $v$ with some vertex on $s_{j}^{\prime}$.
For illustrations, refer to Fig. 1 (where for each $i\in[9]$ and $s=s_{i}$ and $U=U_{i}$, we have $s_{j}^{\prime}=s_{i-1}$ and $U_{j}^{\prime}=U_{i-1}$ for some $j\in[4]$). The following technical claim describes our recurrence relation.
[…]
$T[s,U]=\texttt{true}$ if and only if there exists an index $1\leq j\leq h$ where $p_{j}\geq 1$ and such that $T[s_{j}^{\prime},U_{j}^{\prime}]=\texttt{true}$.
Our table $T$ has $2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1)$ entries. In view of 1, we can compute the value of a table entry $T[s,U]$ in polynomial time (disregarding the time spent for the recursive calls) by simply constructing the tuples $(s_{1}^{\prime},U_{1}^{\prime}),\dots,(s_{h}^{\prime},U_{h}^{\prime})$ and then setting $T[s,U]=\bigvee_{j\in[h]}T[s_{j}^{\prime},U_{j}^{\prime}]$ (only taking into account the defined quantities). Hence, by employing memoization and the usual back-linked strategy, we obtain the claimed deterministic algorithm. For the nondeterministic algorithm, we perform $2n$ steps. In each step, we nondeterministically guess the next separation $s$ and its set of prescribed edges $U$ and check whether for some $j\in[h]$, the previous separation $s^{\prime}$ is equal to $s_{j}^{\prime}$ and the previous set of prescribed edges $U^{\prime}$ is equal to $U_{j}^{\prime}$ (starting with $s^{\prime}=(0,0,\dots,0)$ and $U^{\prime}=\emptyset$). At each point in time, we only need to keep two separations, two edge sets, and $\mathcal{O}(1)$ pointers in memory, which can be done using $\mathcal{O}(h^{2}+h\log I)$ space.
As a corollary of Lemmas 2.1, 2.2 and 2.4, we obtain the algorithmic statements in Theorem 1.1.
3 XNLP-Hardness of Ordered Level Planarity
[…]
Basic building blocks of our reduction.
Our construction of $\mathcal{G}$ is heavily based on two gadgets that we call plugs and sockets (a very basic version of these gadgets was already used earlier, in an NP-hardness proof for CLP [7]; here we introduce generalized versions). We define both in terms of the list of levels their vertices occupy. Their $\prec_{i}$ orderings are according to the indices of their vertices. The vertices are deliberately given in an unintuitive order to allow for the ordering in $\prec_{i}$ by indices and to make the (degenerate) cases behave nicely later. Let $\ell_{1},\ell_{2},\ell_{3},\ell_{4}\in[h]$ such that $\ell_{1}<\ell_{2}<\ell_{3}<\ell_{4}$. A (non-degenerate) $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-plug, see Fig. 2, contains vertices $u_{1}$, $u_{2}$, $u_{3}$, $u_{4}$, $u_{5}$, and $u_{6}$, where $\gamma(u_{5})=\ell_{1}$, $\gamma(u_{3})=\gamma(u_{6})=\ell_{2}$, $\gamma(u_{2})=\gamma(u_{4})=\ell_{3}$, and $\gamma(u_{1})=\ell_{4}$. It contains the edges $u_{1}u_{2}$, $u_{2}u_{3}$, $u_{3}u_{4}$, $u_{4}u_{6}$, and $u_{6}u_{5}$; i.e., a plug is a path that traverses its four levels in the order $\ell_{4},\ell_{3},\ell_{2},\ell_{3},\ell_{2},\ell_{1}$. Similarly, an $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-socket, see Fig. 2, consists of vertices $v_{1},v_{2},\dots,v_{10}$ such that $\gamma(v_{3})=\ell_{1}$, vertices $v_{2},v_{4},v_{5},v_{9}$ occupy level $\ell_{2}$, vertices $v_{1},v_{6},v_{7},v_{10}$ occupy level $\ell_{3}$, and $\gamma(v_{8})=\ell_{4}$. It contains the edges $v_{1}v_{2}$, $v_{2}v_{3}$, $v_{3}v_{5}$, $v_{5}v_{7}$, $v_{4}v_{6}$, $v_{6}v_{8}$, $v_{8}v_{10}$, and $v_{10}v_{9}$. Observe that a socket consists of two disconnected paths whose vertices interleave on levels $\ell_{2}$ and $\ell_{3}$. Let vertices $v_{1}$ and $v_{9}$ of plugs and vertices $u_{1}$ and $u_{5}$ of sockets be connecting vertices. Connecting vertices of sockets will be identified with other vertices of the construction and connecting vertices of some plugs may be connected to other plugs via additional edges.
{subfigure}
[…]
Figure 2: Plugs and sockets; (a) a $(1,3,5,7)$-plug, (b) a $(2,3,5,6)$-socket, (c) a degenerate $(3,3,5,5)$-plug, (d) a degenerate $(2,4,4,6)$-socket, (e) a $(1,3,5,7)$-plug that is linked to a degenerate $(2,4,4,6)$-socket. Connecting vertices are filled in black. Note how in degenerate gadgets ((c) and (d)), the edges and vertices of the repeated levels are contracted.
Now we lift the strict inequality restriction on the levels of our gadgets. For plugs we only require $\ell_{1}\leq\ell_{2}<\ell_{3}\leq\ell_{4}$ and for sockets we require $\ell_{1}<\ell_{2}\leq\ell_{3}<\ell_{4}$ and we call a plug or a socket degenerate if it has at least one pair of repeated levels, i.e., for some $i\in[3]$, $\ell_{i}=\ell_{i+1}$. We create the degenerate plugs and sockets by contracting the edges between vertices of the repeated levels while keeping the vertex with the lower index; see Fig. 2 and 2.
Let $P$ be a plug, and let $\mathrm{com}(P)$ be the connected component of $P$ in $G$. A plug fits into a socket if the gadgets could be “weaved” as illustrated in Fig. 2. Formally, we say an $(a_{1},a_{2},a_{3},a_{4})$-plug $P$ fits into a $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $\min_{v\in\mathrm{com}(P)}\{\gamma(v)\}\leq s_{1}$, $\max_{v\in\mathrm{com}(P)}\{\gamma(v)\}\geq s_{4}$, and $s_{1}<a_{2}<s_{2}\leq s_{3}<a_{3}<s_{4}$. In an ordered level planar drawing of $\mathcal{G}$, we say a plug $P$ links to an $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $P$ fits into $S$, and $P$ is drawn between the connecting vertices of $S$ (that is, the edges of $P$ traversing level $s_{3}$ are to the right of $v_{1}$ and the edges of $P$ traversing level $s_{2}$ are to the left of $v_{9}$); see Fig. 2.
A defining feature of our constructions is a division into vertical strips, which is accommodated by the following notion. Given a set $L$ of levels, a wall of $L$ is a path that starts in a vertex on the bottommost level in $L$, goes through a vertex on each intermediate level of $L$, and ends at a vertex on the topmost level of $L$. Note that each gadget type essentially has a unique drawing in the sense that it corresponds to an ordered level planar graph for which all of its ordered level planar drawings have the same level planar embedding. This and further gadget properties are detailed in Appendix B. We use Lemma 3.1 to design specific plugs that can or cannot link to specific sockets in the same ordered level planar drawing.
[…]
Consider an ordered level graph $\mathcal{G}$ that contains an $(a_{1},a_{2},a_{3},a_{4})$-plug $A$, a $(b_{1},b_{2},b_{3},b_{4})$-plug $B$, and an $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ that occupy three disjoint sets of levels. There is an ordered level planar drawing of the subgraph of $\mathcal{G}$ spanned by $A$, $B$, and $S$ where $A$ and $B$ link to $S$ if and only if both $A$ and $B$ fit into $S$ and $a_{2}<b_{2}$ and $a_{3}<b_{3}$, or, both vice versa, $a_{2}>b_{2}$ and $a_{3}>b_{3}$.
Proof 3.2 (Proof (sketch)).
To distinguish the vertices of the plugs $A$ and $B$, we use the name of the plug as superscript, so, e.g., we refer to $u_{3}^{A}$ and $u_{3}^{B}$. If the conditions are satisfied, we can draw both plugs as depicted in Fig. 3. Towards a contradiction, assuming the conditions are not satisfied, we observe that the drawing must locally “nest” plug $B$ into plug $A$ while drawing them around the edge $v_{4}v_{6}$ of $S$, at the same time $B$ “nests” into $A$ while going around the edge $v_{7}v_{5}$ of $S$; see Fig. 3. The first nesting implies that the edge $u^{B}_{3}u^{B}_{4}$ is to the left of $u^{A}_{3}u^{A}_{4}$ and the second nesting implies the opposite order – a contradiction.
{subfigure}
[…]
Figure 3: Sketches for the proof of Lemma 3.1. The edges of the socket $S$ are bold.
$W[1]$-hardness.
Multicolored Independent Set (MCIS) is a well-known $W[1]$-hard problem [13, 25]. In this problem, we are given a graph $H$, an integer $k$ (the parameter), and a $k$-partition $C_{1},C_{2},\dots,C_{k}$ of $V(H)$, and the task is to decide whether $H$ has an independent set $X\subseteq V(H)$ that contains, for every $j\in[k]$, exactly one vertex of $C_{j}$. We may assume without loss of generality that in $H$, there is no edge whose endpoints have the same color $C_{j}$.
[…]
Ordered Level Planarity is $W[1]$-hard with respect to the height.
Proof 3.4 (Proof (sketch)).
We give an overview of a parameterized reduction from MCIS to OLP; refer to Fig. 4 for an illustration. The ordered level graph constructed in this reduction admits a grid-cell structure where the vertical strips are separated by walls and the horizontal strips represent related groups of levels. More precisely, we allocate a constant number of consecutive levels per color of the MCIS instance, to which we refer as a color band. Between each pair of neighboring color bands, we have a collision band. We group several consecutive vertical strips to obtain blocks. We call the leftmost and rightmost block choice blocks and the other blocks edge blocks. Edge blocks correspond bijectively to edges of the MCIS instance.
In the grid cells of the color bands, we insert sockets and we add as many plugs as we have sockets there. No two plugs can link to the same socket, so all available sockets are occupied by one plug. Each plug is a separate connected component, which allows that plug to freely “choose” a socket within its color band where it fits. However, we have different types of plugs and sockets – the first type of plug, called high plug, fits only into the sockets of the choice blocks. In a drawing of our ordered level graph, the number of high plugs that link to the left (and not to the right) choice block within a color band forces the second type of plugs, called color plugs, to be shifted further to the right. In the MCIS instance, a shift by $i$ corresponds to selecting the $i$-th vertex of the respective color to be in the solution.
To obtain an independent set, we need to prevent that the shifts select both endpoints of an edge. To that end, we model an edge $uv$ as follows. The color plugs of one color band all occupy the same set of levels so they form a sequence specified in the ordered level graph. Because of this, we can identify, for a given shift, what socket a particular color plug links to. Consider the edge block of $uv$ and the color plugs that necessarily link to its sockets (a block has circa twice as many sockets than the maximal shift). For an illustration of the construction, assume that we have a shift in the color band of $u$ that selects $u$ and a shift in the color band of $v$ that selects $v$. We select a specific color plug of the color band of $u$ and one in the color band of $v$ to be extended so that they “reach” towards each other and end up in the same cell of a collision band. To this cell, we add a collision socket. If at most one extended plug ends up in a collision socket, then it can be drawn planarly. However, if both extended plugs end up in the same collision socket, then this cannot be drawn without crossings. Hence, there is an ordered level planar drawing of our ordered level graph if and only if there is a solution (i.e., an independent set) to the given MCIS instance.
Figure 4: Example of our parameterized reduction from Multicolored Independent Set (MCIS) to Ordered Level Planarity. On the left side, there is an instance of MCIS with $k=5$ colors. On the right side, there is the schematized grid structure of the ordered level graph constructed from the MCIS instance. The orange disks in the edge blocks represent the places where the collision sockets are placed. Here, the solution $\{v_{2},v_{3},v_{6},v_{7},v_{8}\}$ for the MCIS instances is found.
XNLP-hardness.
We now extend our construction for $W[1]$-hardness to obtain XNLP-hardness as well. We design a parameterized tractable log-space reduction [6, Section V.B.] – a parameterized reduction that runs in $\mathcal{O}(g(k)\cdot n^{c})$ time and uses only $\mathcal{O}(f(k)\cdot\log n)$ space for internal computation, where $g$ and $f$ are computable functions, $k$ is the parameter, and $n$ is the input size. It suffices to design the reduction such that it returns a single bit of the output on demand. Observe that with such a reduction, we can retrieve the entire output by requesting one bit at a time. We reduce from Chained Multicolored Independent Set (CMCIS) – an XNLP-complete problem defined by Bodlaender et al. [6], which is a sequential version of MCIS. As in MCIS, we are given a $k$-colored graph $H$ with color classes $C_{1},\dots,C_{k}$ and, additionally, there is an $r$-partition $V_{1},\dots,V_{r}$ of $V(H)$ such that for every $vw\in E(H)$, if $v\in V_{i}$ and $w\in V_{j}$, then $|i-j|\leq 1$. The task is to select an independent set $X\subseteq V(G)$ such that, for each $i\in[r]$ and for each color $j\in[k]$, $|X\cap V_{i}\cap C_{j}|=1$.
[…]
Ordered Level Planarity is XNLP-hard with respect to the height.
Proof 3.6 (Proof (sketch)).
The main idea in our reduction from CMCIS to OLP is to use $r$ instances of the $W[1]$-hardness construction from Theorem 3.3. We have one instance for each subgraph induced by $V_{i}$ (for $i\in[r]$) where additionally the edges connecting $V_{i}$ with $V_{i-1}$ and $V_{i+1}$ (if existent) are represented by edge blocks shared by the corresponding two instances; see Fig. 5. To this end, we arrange all instances in a zigzag pattern in the top and the bottom half of the combined construction. Now to make it a parameterized tractable log-space reduction, which outputs bit by bit, we essentially traverse the grid structure with a vertical sweep-line from left to right (i.e., column by column) and save as the status only the $\mathcal{O}(k)$ information belonging to the current column and the $\mathcal{O}(\log n)$ information telling us in which column we are currently globally. We describe this in more detail in the appendix.
Figure 5: Comparison of the $W[1]$-hardness construction (left) and the XNLP-hardness construction (right) with $r=3$ and $k=5$. The thick walls terminating the choice blocks have vertices on all levels. Vertical bars and orange disks between color bands represent the extended plugs and the collision sockets, respectively, which are used for the collision mechanism.
The construction in the proof of Theorem 3.5 can be altered to make the graph connected as we show at the end of Appendix B. The main idea is to add a new vertex $q$ on top of the construction and connect all connected components to it. All sockets are already connected to walls and from walls it suffices to create edges to $q$ from their topmost vertex. For the plugs, we duplicate every column into a left and a right part such that each left part behaves as before, while each right part hosts paths that connect the “old” plugs to $q$. These paths are made of a sequence of a new kind of plugs. This comes at the expense of using more levels, namely $\Theta(k^{2})$ instead of $\Theta(k)$. This yields the XNLP-hardness statement in Theorem 1.1.
4 Tractability of 3-Level Constrained Level Planarity
In this section, we sketch our polynomial-time algorithm for 3-level CLP (see Theorem 1.2). For simplicity, we assume that the given level graph $\mathcal{G}$ is constrained level planar and show how to compute a constrained level planar drawing of $\mathcal{G}$. We ask the reader to consult the figures to which we refer in order to get an intuitive understanding of the notions that we cannot define formally due to space constraints. For details and proofs, see Appendix E.
Without loss of generality, we make the following assumptions: (i) $\mathcal{G}$ is proper; otherwise we subdivide each edge that crosses level 2. (ii) $\mathcal{G}$ has no isolated vertices; they can easily be inserted in a postprocessing step. (iii) The component–constraint graph $H$ is strongly connected. This directed graph has a node for each connected component (for short, component) of $\mathcal{G}$ and an arc for each pair $(C,C^{\prime})$ of components such that there is a constraint from a vertex in $C$ to a vertex in $C^{\prime}$.
We often refer to level 1 as bottom level, to level 2 as middle level, and to level 3 as top level. Similarly, we call the pair of bottom and middle level lower band and the pair of middle and top level upper band. Our algorithm successively adds new constraints to the given constrained level graph $\mathcal{G}$ that do not violate planarity. We deduce these new constraints from the structure of $(G,\gamma)$ and from the current set of constraints. In the end, this yields a total order of the vertices for each of the three levels and, hence, a constrained level planar drawing of $\mathcal{G}$. In the very beginning and whenever we add new constraints, we exhaustively add the following implicit constraints:
  • transitivity: $\forall a,b,c\in V(G),\forall i\in[3]\colon a\prec_{i}b\land b\prec_{i}c\Rightarrow a\prec_{i}c$
  • planarity: $\forall ab,cd\in E(G)\text{ with }i:=\gamma(a)=\gamma(c)\text{ and }j:=\gamma(b)=\gamma(d)\colon\\ a\prec_{i}c\Rightarrow(b\prec_{j}d\lor b=d)$
The former ensures that the orderings $(\prec_{i})_{i\in[3]}$ remain transitive while the latter can be added without violating realizability, as they need to be respected in every constrained level planar drawing. The propagation of these constraints is quite useful as it can dictate the relative positions of vertices that are initially unrelated, see Fig. 6.
{subfigure}
[…]

### v2

@@
-Constrained and Ordered Level Planarity Parameterized by the Number of Levels
+
 Abstract
@@
 We complement this result by showing that CLP can be solved in polynomial time for instances of height at most 3.
-keywords:
-Parameterized Complexity, Graph Drawing, XNLP, XP, W[t]-hard, Level Planarity, Planar Poset Diagram, Computational Geometry
 1 Introduction
@@
 The instances constructed by their reduction are connected and have bounded maximum degree.
-They also present a polynomial time algorithm for CLP for the case where the graph has a single sink, which they later sped up [8].
+They also present a polynomial time algorithm for CLP for the case where the graph has a single sink, which they later [9] sped up for the special case where the single-sink graph is also biconnected.
+Very recently, Klemz and Sieper [28] showed that CLP (and, thus, OLP) is FPT if parameterized by the vertex cover number of the input graph.
+They also observed that the reduction by Klemz and Rote (Brückner and Rutter) can easily be modified to show that OLP (CLP) is NP-hard even if restricting to (proper) instances of bounded treedepth.
 Other related work.
@@
 Contribution and organization.
-As stated above, Klemz and Rote [24] showed that OLP (and, thus, CLP) is NP-hard even when restricted to instances of levelwidth two.
+As discussed above, the parameterized complexity of OLP and CLP with respect to classical graph parameters (vertex cover number, feedback vertex set number, treedepth, pathwidth, treewidth, maximum degree) has been explored exhaustively [27, 8, 28].
+In terms of more problem specific parameters, Klemz and Rote [27] showed that OLP (and, thus, CLP) is NP-hard even if restricted to instances of levelwidth two.
 In this paper, we focus on the other “dimension”: we study CLP and OLP parameterized by height.
@@
 In particular, if a problem is XNLP-complete, it lies in XP, but is $W[t]$-hard for every $t$ [7].
-Theorem 1.1.
+Theorem 1.
 Ordered Level Planarity parameterized by the height of the input graph is XNLP-complete (and, thus, it lies in XP, but is $W[t]$-hard for every $t$).
@@
 However, to the best of our knowledge, this is the first graph drawing (or computational geometry) problem shown to be XNLP-complete.
-The algorithms are described in Section 2, whereas the hardness is shown in Section 3.
+The algorithms are described in Section 2 (Theorem 3), whereas the hardness is shown in Section 3 (Theorem 5).
 In contrast to the fact that OLP parameterized by the height lies in XP, it is not difficult to see that the socket/plug gadget described by Brückner and Rutter [8] can be utilized in the context of a reduction from 3-Partition to show that (PLP and, thus) CLP remains NP-hard even when restricted to instances of constant height.
 In fact, the unpublished full version of [8] features such a construction with a height of 7 [31].
-Here, we present a reduction that is tailor-made for CLP, showing that it is NP-hard even when restricted to instances of height 4.
-We complement this result by showing that the (surprisingly challenging) case of instances with height at most three can be solved in polynomial time.
-Theorem 1.2.
-Constrained Level Planarity is NP-hard even when restricted to height 4, but instances of height at most 3 can be solved constructively in polynomial time.
-We show the hardness in Appendix C, describe the algorithm in Section 4, and state open problems in Section 5. Proofs of statements marked with a (clickable) $\star$ are in the appendix.
+Here, we present a reduction that is tailor-made for CLP, showing that it is NP-hard even if restricted to instances of height 4.
+We complement this result by showing that the (surprisingly challenging) case of instances with height at most 3 can be solved in polynomial time.
+Theorem 2.
+Constrained Level Planarity is NP-hard even if restricted to height 4, but instances of height at most 3 can be solved constructively in polynomial time.
+We show the hardness in Section 4 (Theorem 6) and present the algorithm in Section 6 (Theorem 8).
+As a warm-up for the rather technical positive result, we show that CLP can be solved in linear time when restricted to instances of height at most 2; see Section 5 (Theorem 7).
+We conclude by stating some open problems in Section 7.
 Notation and conventions.
@@
 2 An XP / XNLP Algorithm for Ordered Level Planarity
-In this section, we show that Ordered Level Planarity is in XNLP (and thus in XP) when parameterized by the height $h$ of the input graph.
+In this section, we show that Ordered Level Planarity is in XNLP (and thus in XP) if parameterized by the height $h$ of the input graph.
 Moreover, we show how to construct an ordered level planar drawing (if it exists) in XP-time.
@@
 To discretize this idea and turn it into an algorithm, we instead determine a sequence of unbounded y-monotone curves $S=(s_{1},s_{2},\dots,s_{z})$ that is sorted from left to right (i.e., no point of $s_{i+1}$ is to the left of $s_{i}$), has a length of $z\in\mathcal{O}(n)$, and contains for every edge $(u,v)$ a curve $s_{i}$ along which $u$ and $v$ are consecutive vertices.
-Now, given $S$, the desired drawing can be constructed in polynomial time; for an illustration see Fig.
-1.
+Now, given $S$, the desired drawing can be constructed in polynomial time; for an illustration see Figure 1.
 Moreover, the sequence $S$ can be obtained in XP-time/space by exhaustively enumerating all possibilities or in XNLP-time/space by nondeterministic guessing.
@@
 A separation for $\mathcal{G}$ is an element of $P_{1}\times P_{2}\times\dots\times P_{h}$.
-Intuitively, a separation $s=(p_{1},p_{2},\dots,p_{h})$ represents the equivalence class of unbounded y-monotone curves that intersect line $L_{i}$ in position $p_{i}$ for each $1\leq i\leq h$; see Fig.
-1.
+Intuitively, a separation $s=(p_{1},p_{2},\dots,p_{h})$ represents the equivalence class of unbounded y-monotone curves that intersect line $L_{i}$ in position $p_{i}$ for each $1\leq i\leq h$; see Figure 1.
 We say that a vertex $v\in V$ is on $s$ if $p_{\gamma(v)}$ represents $v$.
@@
 We say that a sweeping sequences uses an edge $e\in E$, if it contains a separation that uses $e$.
-[…]
+Drawing algorithm.
+The following lemma formalizes the idea of our drawing algorithm.
+Lemma 1.
 Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph and let $S=(s_{1},s_{2},\dots,s_{z})$ be a sweeping sequence for $\mathcal{G}$ that uses every edge of $E$.
 Given $\mathcal{G}$ and $S$, an ordered level planar drawing of $\mathcal{G}$ can be constructed in $\mathcal{O}(zh+n^{2})$ time, where $n$ is the number of vertices.
+Proof.
+Without loss of generality, we assume that $s_{1}=(0,0,\dots,0)$ (otherwise, we simply prepend this separation to $S$).
+Recall that a separation $s_{j}=(p_{1},p_{2},\dots,p_{h})$ corresponds to an equivalence class of unbounded y-monotone curves that intersect line $L_{i}$ in position $p_{i}$ for each $1\leq i\leq h$.
+For each $j\in[z]$, let $c_{j}$ denote a member of the class of curves corresponding to $s_{j}$.
+The fact that $S$ is a sweeping sequence implies that we may assume that for every $j\in[z-1]$, no point of $c_{j+1}$ lies to the left of $c_{j}$ and that the only points shared by $c_{j}$ and $c_{j+1}$ are those that correspond to vertices.
+In particular, the curve $c_{1}$ passes through the left-most gap on each level.
+To prove the existence of the desired drawing, it suffices to show that for each $j\in[z]$ there is an ordered level planar drawing $\Gamma_{j}$ that contains all vertices that are not to the right of $s_{j}$ and all edges that are used by one of the separations $s_{1},s_{2},\dots,s_{j}$ and where the vertices on $s_{j}$ and the edges used by $s_{j}$ are drawn on the curve $c_{j}$ and the remaining vertices and edges are drawn to the left of $c_{j}$.
+Note that, indeed, $\Gamma_{z}$ is the desired drawing of $\mathcal{G}$.
+The (empty) drawing $\Gamma_{1}$ is trivially constructed.
+To construct $\Gamma_{j+1}$ from $\Gamma_{j}$, we simply place all (isolated) vertices that are to the right of $s_{j}$ and to the left of $s_{j+1}$ between $c_{j}$ and $c_{j+1}$ and then draw on $c_{j+1}$ all the new vertices on $s_{j+1}$, as well as the not yet drawn edges that are used by $s_{j+1}$.
+By definition of the curves $c_{1},c_{2},\dots,c_{j+1}$, this process cannot introduce any crossings.
+For illustrations, refer to Figure 1.
+The above inductive proof directly corresponds to an iterative algorithm (note that we do not have to explicitly construct the curves $c_{1},c_{2},\dots,c_{z}$).
+The total time for adding the vertices is $\mathcal{O}(zh+n)$.
+To add the edges, whenever there is a new pair of consecutive vertices on the current separation, we need to check whether these vertices are adjacent, which takes $\mathcal{O}(n)$ time.
+New pairs of consecutive vertices are only created when a vertex appears or disappears from the current separation.
+Moreover, the appearance/disappearance of a vertex can only create $\mathcal{O}(1)$ new pairs of consecutive vertices.
+Combined with the fact that for each vertex there is an interval of separations in $S$ on which it is located, it follows that the total number of adjacency checks we have to perform is $\mathcal{O}(n)$.
+Thus, the total time for the adjacency checks is $\mathcal{O}(n^{2})$.
+Drawing an edge takes time linear in the number of levels it spans, which is at most $h\in\mathcal{O}(n)$.
+Moreover, by planarity (note that the inductive proof shows that $G$ is planar), the number of edges is $\mathcal{O}(n)$.
+Hence, the time to draw all the edges is $\mathcal{O}(n^{2})$, and the total runtime is $\mathcal{O}(zh+n^{2})$, as claimed.
+∎
+In the remainder of this section, we will show that for each ordered level graph that admits an ordered level planar drawing there exists a particularly well-structured sweeping sequence that uses all of its edges.
+Moreover, we will show that the existence of such a sequence can be tested efficiently.
+Combined with Lemma 1, this results in the desired XP / XNLP algorithms.
+Nice and exhaustive sweeping sequences.
 Let $S=(s_{1},s_{2},\dots,s_{z})$ be a sweeping sequence for our ordered level graph $\mathcal{G}$.
 We say that $S$ is nice if for each $i\in[z-1]$, the separation $s_{i+1}$ is obtained from $s_{i}$ by incrementing exactly one component by $1$.
-Moreover, we say that $S$ is exhaustive if it is nice and $s_{1}=(0,0,\dots,0)$ and $s_{z}=(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1)$; see Fig.
-1.
-[…]
+Moreover, we say that $S$ is exhaustive if it is nice and $s_{1}=(0,0,\dots,0)$ and $s_{z}=(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1)$; see Figure 1.
+Lemma 2.
 Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph.
-Then $\mathcal{G}$ admits an ordered level planar drawing if and only if there is an exhaustive sweeping sequence for $G$ that uses every edge in $E$.
-Proof 2.3 (Proof (sketch)).
-The “if”-direction follows from Lemma 2.1.
-For the “only if”-direction, we show that the drawing can be internally triangulated by inserting y-monotone edges (allowing the use of parallel edges).
-For each y-monotone path $P$ that spans all levels, there is a nice sweeping sequence from $s_{1}=(0,\dots,0)$ to the separation corresponding to $P$ that uses all edges left of $P$.
-This is shown by arguing inductively on the number of triangles left of $P$.
+Then $\mathcal{G}$ admits an ordered level planar drawing if and only if there is an exhaustive sweeping sequence for $\mathcal{G}$ that uses every edge in $E$.
+Proof.
+The “if”-direction follows immediately from Lemma 1.
+For the “only if”-direction, suppose that $\mathcal{G}$ admits an ordered level planar drawing $\Gamma$.
+We augment $\mathcal{G}$ and $\Gamma$ as follows: first, we insert a new bottommost and topmost level with a single vertex $a$ and $b$, respectively.
+Second, we insert a maximal set of y-monotone crossing-free edges (i.e., the resulting drawing is supposed to be crossing-free).
+We explicitly allow the introduction of parallel edges, but we require that for each pair of parallel edges $e,e^{\prime}$ there is a vertex in the interior of the simple closed curve formed by $e,e^{\prime}$.
+Let $\mathcal{G}^{\prime}$ and $\Gamma^{\prime}$ denote the resulting graph and drawing, respectively, which is an internal triangulation:
+Claim 1.
+The outer face of $\Gamma^{\prime}$ is bounded by two parallel $(a,b)$-edges and every internal face of $\Gamma^{\prime}$ is a triangle (i.e., its boundary corresponds to a cycle of length three).
+Proof.
+The statement regarding the outer face obviously holds.
+Now consider an internal face $f$.
+Towards a contradiction, assume that $f$ is not a triangle.
+Let $v$ be a bottommost vertex of $f$, which is incident to (at least) two edges of $f$.
+Let $e_{1}=(v,v_{1}),e_{2}=(v,v_{2})$ be two edges that are consecutive in the cyclic order of edges incident to $v$ and consecutive along the boundary of $f$.
+Without loss of generality, we may assume that $\gamma(v_{2})\leq\gamma(v_{1})$.
+Moreover, without loss of generality, we may assume that $e_{2}$ is to the right of $e_{1}$.
+We distinguish two main cases.
+Case 1.
+There is a vertex $u$ with $\gamma(v)<\gamma(u)\leq\gamma(v_{2})$ that is located to the right of $e_{1}$ and to the left of $e_{2}$.
+We can add an edge from $v$ to the bottommost vertex with these properties in a y-monotone crossing-free fashion in $f$.
+Moreover, this edge cannot be parallel to some edge of $\Gamma^{\prime}$; a contradiction to the definition of $\Gamma^{\prime}$.
+Case 2.  There is no vertex $u$ with $\gamma(v)<\gamma(u)\leq\gamma(v_{2})$ that is located to the right of $e_{1}$ and to the left of $e_{2}$. We distinguish two subcases.
+Case 2.1.$\gamma(v_{2})<\gamma(v_{1})$.
+In this case, we can add an edge from $v_{2}$ to $v_{1}$ in a y-monotone crossing-free fashion in $f$ such that, even if this edge is parallel to some other edge, the interior of the region bounded by these edges contains at least one vertex (recall that $f$ is not a triangle by assumption).
+So again, we obtain a contradiction to the definition of $\Gamma^{\prime}$.
+Case 2.2.$\gamma(v_{2})=\gamma(v_{1})$.
+In this case, we draw a y-monotone curve $c$ in $f$ that starts at $v$ and extends upwards until it crosses some edge $e_{3}=(u,w)$ with $e_{3}\notin\{e_{1},e_{2}\}$.
+The existence of $e_{3}$ follows from the existence of the two edges of the outer face.
+By applying the assumption of Case 2 to $u$, it follows that the upper endpoint $w$ of $e_{3}$ is neither $v_{1}$ nor $v_{2}$.
+Hence, we can add an edge from $v$ to $w$ in a y-monotone crossing-free fashion in $f$ such that, even if this edge is parallel to some other edge, the interior of the region bounded by these edges contains at least one vertex ($v_{1}$ or $v_{2}$).
+So again, we obtain a contradiction to the definition of $\Gamma^{\prime}$.
+∎
+Let $e_{\ell}$ and $e_{r}$ denote the left and right edge of the outer face of $\Gamma^{\prime}$.
+To prove the lemma, it suffices to show that there is a nice sweeping sequence for $\mathcal{G}^{\prime}$ that uses all of its edges and starts with separation $s_{1}=(1,0,0,\dots,1)$ and ends with separation $s_{z}=(1,|P_{1}|-1,|P_{2}|-1,\dots,|P_{z}|-1,1)$, which use $e_{\ell}$ and $e_{r}$, respectively.
+To this end, we show:
+Claim 2.
+Let $P$ be a directed (y-monotone) $ab$-path in $\Gamma^{\prime}$ and let $s$ be the (unique) separation that corresponds to $P$.
+Then there is a nice sweeping sequence from $s_{1}$ to $s$ that uses all edges of $\Gamma^{\prime}$ that are not located to the right of $P$.
+Proof.
+The proof is by induction on the number $i$ of triangles to the left of $P$. If $i=0$ (that is, $P=(e_{\ell})$), the statement obviously holds. So assume $i>0$.
+Suppose there is an edge $(u,v)$ on $P$ that is incident to a triangle $uvw$, where $w$ is to the left of $P$ and $\gamma(u)<\gamma(w)<\gamma(v)$; we call this Configuration (a).
+In this case, we can replace $(u,v)$ with $(u,w)$ and $(w,v)$ to obtain a directed $ab$-path $P^{\prime}$ from $P$ with $i-1$ triangles to its left and the claim follows by induction.
+Next, suppose that there is a triangle $uvw$ to the left of $P$ whose edges $(u,w)$ and $(w,v)$ belong to $P$; we call this Configuration (b).
+We can replace $(u,w)$ and $(w,v)$ with $(u,v)$ to obtain a directed $ab$-path $P^{\prime}$ from $P$ with $i-1$ triangles to its left and the claim follows by induction.
+We claim that one of these two configurations always exists.
+Towards a contradiction, assume otherwise.
+Let $e=(u,v)$ be an edge on $P$ that is incident to a triangle $\triangle=uvw$, where $w$ is to the left of $P$.
+Since Configuration (a) does not exist, we may assume without loss of generality that $(\gamma(u)<)\gamma(v)<\gamma(w)$.
+Further, we may assume without loss of generality that $e$ is the topmost edge on $P$ with this property.
+Since Configuration (b) does not exist, the edge $(v,w)$ of $\triangle$ does not belong to $P$.
+Let $x$ denote the successor of $v$ along $P$ and let $vxy$ denote the triangle to the left of $(v,x)$ on $P$.
+Since the edge $(v,w)$ of $\triangle$ is crossing-free and Configuration (b) does not exist, it follows that $(\gamma(v)<)\gamma(x)<\gamma(y)$; a contradiction to the choice of $e$.
+∎
+The desired sequence from $s_{1}$ to $s_{z}$ is obtained by Claim 2, choosing $P=(e_{r})$. ∎
 We remark that an exhaustive sweeping sequence using all edges corresponds directly to a particularly well-structured path decomposition.
-Thus, Lemma 2.2 implies that every (ordered) level planar drawing of height $h$ represents a graph of pathwidth at most $h-1$; a statement that was independently proven in [11].
-However, the path decompositions constructed in the proof of [11, Lemma 1] do not exhibit the same properties that are inherent to exhaustive sweeping sequences and on which our algorithms heavily rely.
+Thus, Lemma 2 implies that every (ordered) level planar drawing of height $h$ represents a graph of pathwidth at most $h-1$; a statement that was independently proven in [12].
+However, the path decompositions constructed in the proof of [12, Lemma 1] do not exhibit the same properties that are inherent to exhaustive sweeping sequences and on which our algorithms heavily rely.
 In particular, these path decompositions may contain bags with multiple vertices of a given level (unless the drawing is proper).
 Moreover, the existence of a path decomposition of width at most $h-1$ for an ordered level graph $\mathcal{G}$ of height $h$ does not characterize the fact that $\mathcal{G}$ is ordered level planar (recall that, in fact, OLP is NP-hard even when restricted to instances of pathwidth $1$ [27]).
-Figure 1: An exhaustive sweeping sequence using all edges of the depicted graph and the corresponding (cf.
-Lemma 2.1) ordered level planar drawing, as well as the corresponding sequence of true dynamic programming table entries $T[s,U]$ (cf.
-Lemma 2.4).
-Note that in any ordered level planar drawing of the graph, exactly one of the edges $(2,4),(1,3)$ is located to the left of the path $(1,2,3,4)$, while the other is located to the right.
-Similarly, in any exhaustive sweeping sequence containing separation $s_{5}$, exactly one of the edges $(2,4),(1,3)$ is used by a separation preceding $s_{5}$, while the other is used by a separation succeeding $s_{5}$.
-Hence, when iteratively building an exhaustive sweeping sequence, it is key to remember which edges between vertices of the current separation have already been used – this is exactly the purpose of the sets $U$.
-E.g., the fact that $(2,4)\in U_{5}$ corresponds to $(2,4)$ being used before $s_{5}$ (in $s_{4}$), from which one can infer how to proceed.
-[…]
+Computing suitable sweeping sequences.
+In view of Lemma 2, recognizing ordered level planar graphs is equivalent to testing for the existence of suitable sweeping sequences. We now show that this can be done efficiently.
+Lemma 3.
 There is an algorithm that determines whether a given ordered level graph admits an exhaustive sweeping sequence using all of its edges.
@@
 Further, the deterministic version can report the sequence (if it exists).
-Proof 2.5 (Proof (sketch)).
-Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph and let $s=(p_{1},p_{2},\dots,p_{h})$ be a separation for $\mathcal{G}$.
-Further, let $U\subseteq E$ be a subset of the edges that are joining pairs of (not necessarily consecutive) vertices on $s$.
+Proof.
+Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph, let $s=(p_{1},p_{2},\dots,p_{h})$ be a separation for $\mathcal{G}$, and let $U\subseteq E$ be a subset of the edges that are joining pairs of (not necessarily consecutive) vertices on $s$.
 We define $T[s,U]=\texttt{true}$ if there exists a nice sweeping sequence for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s$, and uses all edges in $U$, as well as all edges in $E$ incident to at least one vertex to the left of $s$.
@@
 Additionally, we allow $U=\bot$, in which case $T[s,U]=\texttt{false}$.
-Fig.
-1 illustrates several true table entries $T[s_{i},U_{i}]$ along with the corresponding sweeping sequences $s_{1},s_{2},\dots,s_{i}$.
+Figure 1 illustrates several true table entries $T[s_{i},U_{i}]$ along with the corresponding sweeping sequences $s_{1},s_{2},\dots,s_{i}$.
 Our goal is to determine $T[(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1),\emptyset]$, which is true if and only if there exists an exhaustive sweeping sequence for $\mathcal{G}$ using all edges in $E$.
+Recurrence relation.
 We will determine the entries $T[s,U]$ by means of a dynamic programming recurrence.
@@
         adding all edges in $E$ that join $v$ with some vertex on $s_{j}^{\prime}$.
-For illustrations, refer to Fig.
-1 (where for each $i\in[9]$ and $s=s_{i}$ and $U=U_{i}$, we have $s_{j}^{\prime}=s_{i-1}$ and $U_{j}^{\prime}=U_{i-1}$ for some $j\in[4]$).
+Figure 1: An exhaustive sweeping sequence using all edges of the depicted graph and the corresponding (cf.
+Lemma 1) ordered level planar drawing, as well as the corresponding sequence of true dynamic programming table entries $T[s,U]$ (cf.
+Lemma 3).
+Note that in any ordered level planar drawing of the graph, exactly one of the edges $(2,4),(1,3)$ is located to the left of the path $(1,2,3,4)$, while the other is located to the right.
+Similarly, in any exhaustive sweeping sequence containing separation $s_{5}$, exactly one of the edges $(2,4),(1,3)$ is used by a separation preceding $s_{5}$, while the other is used by a separation succeeding $s_{5}$.
+Hence, when iteratively building an exhaustive sweeping sequence, it is key to remember which edges between vertices of the current separation have already been used – this is exactly the purpose of the sets $U$.
+E.g., the fact that $(2,4)\in U_{5}$ corresponds to $(2,4)$ being used before $s_{5}$ (in $s_{4}$), from which one can infer how to proceed.
+For illustrations, refer to Figure 1 (where for each $i\in[9]$ and $s=s_{i}$ and $U=U_{i}$, we have $s_{j}^{\prime}=s_{i-1}$ and $U_{j}^{\prime}=U_{i-1}$ for some $j\in[4]$).
 The following technical claim describes our recurrence relation.
-[…]
+Claim 3.
 $T[s,U]=\texttt{true}$ if and only if there exists an index $1\leq j\leq h$ where $p_{j}\geq 1$ and such that $T[s_{j}^{\prime},U_{j}^{\prime}]=\texttt{true}$.
-Our table $T$ has $2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1)$ entries.
-In view of 1, we can compute the value of a table entry $T[s,U]$ in polynomial time (disregarding the time spent for the recursive calls) by simply constructing the tuples $(s_{1}^{\prime},U_{1}^{\prime}),\dots,(s_{h}^{\prime},U_{h}^{\prime})$ and then setting $T[s,U]=\bigvee_{j\in[h]}T[s_{j}^{\prime},U_{j}^{\prime}]$ (only taking into account the defined quantities).
-Hence, by employing memoization and the usual back-linked strategy, we obtain the claimed deterministic algorithm.
-For the nondeterministic algorithm, we perform $2n$ steps.
-In each step, we nondeterministically guess the next separation $s$ and its set of prescribed edges $U$ and check whether for some $j\in[h]$, the previous separation $s^{\prime}$ is equal to $s_{j}^{\prime}$ and the previous set of prescribed edges $U^{\prime}$ is equal to $U_{j}^{\prime}$ (starting with $s^{\prime}=(0,0,\dots,0)$ and $U^{\prime}=\emptyset$).
-At each point in time, we only need to keep two separations, two edge sets, and $\mathcal{O}(1)$ pointers in memory, which can be done using $\mathcal{O}(h^{2}+h\log I)$ space.
-As a corollary of Lemmas 2.1, 2.2 and 2.4, we obtain the algorithmic statements in Theorem 1.1.
+Proof.
+For the “if”-direction, assume that there exists an index $1\leq j\leq h$ where $p_{j}\geq 1$ and such that $T[s_{j}^{\prime},U_{j}^{\prime}]=\texttt{true}$.
+Then there exists a nice sweeping sequence $S^{\prime}=(s_{1},s_{2},\dots,s_{z}=s_{j}^{\prime})$ for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s_{j}^{\prime}$, and uses all edges in $U_{j}^{\prime}$, as well as all edges in $E$ incident to at least one vertex to the left of $s_{j}^{\prime}$.
+Let $S$ denote the nice sweeping sequence obtained by appending $s$ to $S^{\prime}$.
+We will show that $S$ is a certificate for the fact that $T[s,U]$ is indeed true.
+We start by showing that $S$ uses all edges that are incident to at least one vertex to the left of $s$.
+Let $e\in E$ be an edge with an endpoint $u$ to the left of $s$.
+If at least one endpoint of $e$ is also to the left of $s_{j}^{\prime}$, the definition of $S^{\prime}$ implies that ($S^{\prime}$ and, hence) $S$ uses $e$.
+So assume that no endpoint of $e$ is to the left of $s_{j}^{\prime}$.
+It follows that $p_{j}$ represents a gap, $p_{j-1}$ represents the vertex $u$, and the other endpoint of $e$ is on $s_{j}^{\prime}$ (it cannot lie to the right of ($s$ and, hence) $s_{j}^{\prime}$ since $T[s_{j}^{\prime},U_{j}^{\prime}]=\texttt{true}$ and, hence, $U_{j}^{\prime}\neq\bot$).
+Consequently, $e\in U_{j}^{\prime}$ (by definition of $U_{j}^{\prime}$) and, thus, ($S^{\prime}$ and) $S$ uses $e$, as desired.
+It remains to show that $S$ uses all edges in $U$.
+Obviously, ($s$ and) $S$ use all edges in $U$ that join two vertices that are consecutive along $s$.
+So let $e\in U$ be an edge that joins two vertices that are nonconsecutive along $s$.
+When creating $U_{j}^{\prime}$ from $U$, only edges between consecutive vertices on $s$ are removed.
+Hence, $e\in U_{j}^{\prime}$ and, thus, ($S^{\prime}$ and) $S$ uses $e$, as desired.
+Altogether, it follows that $S$ is indeed a certificate for the fact that $T[s,U]$ is true.
+For the “only if”-direction, assume that $T[s,U]=\texttt{true}$.
+Then there exists a nice sweeping sequence $S=(s_{1},s_{2},\dots,s_{z}=s)$ for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s$, and uses all edges in $U$, as well as all edges in $E$ incident to at least one vertex to the left of $s$.
+Since $S$ is nice, there is an index $1\leq j\leq h$ such that
+$s_{z-1}=(p_{1},p_{2},\dots,p_{j-1},p_{j}-1,p_{j+1},\dots,p_{h}),$
+i.e., $s_{z-1}=s_{j}^{\prime}$.
+Let $S^{\prime}=(s_{1},s_{2},\dots,s_{z-1}=s^{\prime})$.
+We will show that $S^{\prime}$ is a certificate for the fact that $T[s_{j}^{\prime},U_{j}^{\prime}]$ is indeed true.
+We begin by showing that $U_{j}^{\prime}\neq\bot$.
+To this end, assume otherwise, i.e., $U_{j}^{\prime}=\bot$.
+We distinguish two cases.
+First, assume that $p_{j}$ represents a vertex $v$.
+This implies that $U$ contains an edge $e$ joining $v$ with a vertex on $s$ that is not its predecessor or successor.
+The edge $e$ is not used by $s$.
+However, it is also not used by $S^{\prime}$ since $v$ is to the right of $s_{j}^{\prime}$; a contradiction to the fact that $S$ uses all edges in $U$.
+Second, assume that $p_{j}$ represents a gap and, thus, $p_{j-1}$ represents a vertex $v$.
+This implies that $v$ has an edge $e$ to a vertex to the right of $s$.
+Thus, the sequence $S$ does not use $e$.
+However, the endpoint $v$ of $e$ is to the left of $s$; a contradiction to the fact that $S$ uses all edges with at least one endpoint to the left of $s$.
+Altogether, this shows that indeed $U_{j}^{\prime}\neq\bot$.
+We now show that $S^{\prime}$ uses all edges that are incident to at least one vertex to the left of $s_{j}^{\prime}$.
+To this end, let $u\in V$ be a vertex to the left of $s_{j}^{\prime}$.
+By definition of $s_{j}^{\prime}$, this vertex is also to the left of $s$.
+By definition of $S$, all edges incident to $u$ are used by $S$.
+However, none of these edges can be used by $s$ (since $u$ is to the left of $s$) and, hence, each of these edges is used by (some separation in) $S^{\prime}$, as desired.
+It remains to show that $S^{\prime}$ uses all edges in $U_{j}^{\prime}$.
+To this end, let $e\in U_{j}^{\prime}$.
+First, assume that $e\in U_{j}^{\prime}\setminus U$.
+The definition of $U_{j}^{\prime}$ implies that $p_{j}$ represents a gap and $p_{j-1}$ represents a vertex $v$ that is incident to $e$.
+By definition of $S$ and the fact that $v$ is to the left of $s$, it follows that $S$ uses $e$.
+In fact, since $s$ cannot use $e$, it follows that (some separation in) $S^{\prime}$ uses $e$, as desired.
+It remains to consider the case $e\notin U_{j}^{\prime}\setminus U$.
+Note that this case assumption implies that both endpoints of $e$ belong to $s$.
+We distinguish two subcases: first assume that the endpoints of $e$ are consecutive vertices along $s$.
+Combining this assumption with the fact that $e\in U_{j}^{\prime}\cap U$ and the definition of $U_{j}^{\prime}$ implies that the endpoints of $e$ are also consecutive vertices along $s_{j}^{\prime}$.
+Hence, ($s^{\prime}$ and) $S^{\prime}$ uses $e$, as desired.
+For the other subcase, assume that the two endpoints of $e$ are nonconsecutive along $s$.
+In this case, the separation $s$ cannot use $e$ and, hence, the definition of $S$ implies that (some separation in) $S^{\prime}$ uses $e$, as desired.
+Altogether, we have established that $S^{\prime}$ is indeed a certificate for the fact that $T[s_{j}^{\prime},U_{j}^{\prime}]$ is true. This concludes the proof of the claim. ∎
+In view of Claim 3, we can now state our algorithms.
+Deterministic construction algorithm.
+Our dynamic programming table $T$ has a total of $2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1)$ entries.
+To compute the value of a table entry $T[s,U]$, we simply construct the tuples $(s_{1}^{\prime},U_{1}^{\prime}),\dots,(s_{h}^{\prime},U_{h}^{\prime})$ and then set $T[s,U]=\bigvee_{j\in[h]}T[s_{j}^{\prime},U_{j}^{\prime}]$ (the set $U_{j}^{\prime}$ is only constructed and $T[s_{j}^{\prime},U_{j}^{\prime}]$ only taken into account when $s_{j}^{\prime}$ contains no negative entry).
+Disregarding the time spend for the recursive calls, the time to process a table entry is clearly polynomial.
+Hence, by employing memoization, we can fill the table using $\mathcal{O}^{*}(2^{\binom{h}{2}}\prod_{j\in[h]}(2\lambda_{j}+1))$ time and space.
+To construct the solution, we can employ the usual back-linking strategy.
+In particular, for each true entry $T[s,U]$, it suffices to store an index $j$ such that $T[s_{j}^{\prime},U_{j}^{\prime}]$ is true, which does not change the asymptotic time/space requirements.
+Figure 1 illustrates a sequence of table entries corresponding to a solution.
+Nondeterministic decision algorithm.
+We perform $2n$ steps.
+In each step, we nondeterministically guess the next separation $s$ and its set of prescribed edges $U$, and check whether for some $j\in[h]$, the previous separation $s^{\prime}$ is equal to $s_{j}^{\prime}$ and the previous set of prescribed edges $U^{\prime}$ is equal to $U_{j}^{\prime}$ (starting with $s^{\prime}=(0,0,\dots,0)$ and $U^{\prime}=\emptyset$).
+At each point in time, we only need to keep two separations and two edge sets in memory.
+A separation can be stored using $\mathcal{O}(h\log I)$ space.
+Each edge set $U$ can be encoded by means of a $h\times h$ Boolean adjacency matrix of size $\mathcal{O}(h^{2})$.
+The verification in each step is easy to do using polynomial time and $\mathcal{O}(\log I)$ additional space (in particular, we only need to store $\mathcal{O}(1)$ pointers / indices).
+∎
+As a corollary of Lemmas 1, 2 and 3, we obtain the algorithmic statements in Theorem 1:
+Theorem 3.
+Ordered Level Planarity parameterized by the height is XNLP. Moreover, there is a constructive XP-time algorithm for Ordered Level Planarity parameterized by the height.
 3 XNLP-Hardness of Ordered Level Planarity
@@
 Let $\ell_{1},\ell_{2},\ell_{3},\ell_{4}\in[h]$ such that $\ell_{1}<\ell_{2}<\ell_{3}<\ell_{4}$.
-A (non-degenerate) $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-plug, see Fig.
-2, contains vertices $u_{1}$, $u_{2}$, $u_{3}$, $u_{4}$, $u_{5}$, and $u_{6}$, where $\gamma(u_{5})=\ell_{1}$, $\gamma(u_{3})=\gamma(u_{6})=\ell_{2}$, $\gamma(u_{2})=\gamma(u_{4})=\ell_{3}$, and $\gamma(u_{1})=\ell_{4}$.
+A (non-degenerate) $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-plug, see Figure 2(a), contains vertices $u_{1}$, $u_{2}$, $u_{3}$, $u_{4}$, $u_{5}$, and $u_{6}$, where $\gamma(u_{5})=\ell_{1}$, $\gamma(u_{3})=\gamma(u_{6})=\ell_{2}$, $\gamma(u_{2})=\gamma(u_{4})=\ell_{3}$, and $\gamma(u_{1})=\ell_{4}$.
 It contains the edges $u_{1}u_{2}$, $u_{2}u_{3}$, $u_{3}u_{4}$, $u_{4}u_{6}$, and $u_{6}u_{5}$; i.e., a plug is a path that traverses its four levels in the order $\ell_{4},\ell_{3},\ell_{2},\ell_{3},\ell_{2},\ell_{1}$.
-Similarly, an $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-socket, see Fig.
-2, consists of vertices $v_{1},v_{2},\dots,v_{10}$ such that $\gamma(v_{3})=\ell_{1}$, vertices $v_{2},v_{4},v_{5},v_{9}$ occupy level $\ell_{2}$, vertices $v_{1},v_{6},v_{7},v_{10}$ occupy level $\ell_{3}$, and $\gamma(v_{8})=\ell_{4}$.
+Similarly, an $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-socket, see Figure 2(b), consists of vertices $v_{1},v_{2},\dots,v_{10}$ such that $\gamma(v_{3})=\ell_{1}$, vertices $v_{2},v_{4},v_{5},v_{9}$ occupy level $\ell_{2}$, vertices $v_{1},v_{6},v_{7},v_{10}$ occupy level $\ell_{3}$, and $\gamma(v_{8})=\ell_{4}$.
 It contains the edges $v_{1}v_{2}$, $v_{2}v_{3}$, $v_{3}v_{5}$, $v_{5}v_{7}$, $v_{4}v_{6}$, $v_{6}v_{8}$, $v_{8}v_{10}$, and $v_{10}v_{9}$.
@@
 Connecting vertices of sockets will be identified with other vertices of the construction and connecting vertices of some plugs may be connected to other plugs via additional edges.
-{subfigure}
-[…]
 Figure 2: Plugs and sockets; (a) a $(1,3,5,7)$-plug, (b) a $(2,3,5,6)$-socket, (c) a degenerate $(3,3,5,5)$-plug, (d) a degenerate $(2,4,4,6)$-socket, (e) a $(1,3,5,7)$-plug that is linked to a degenerate $(2,4,4,6)$-socket.
@@
 For plugs we only require $\ell_{1}\leq\ell_{2}<\ell_{3}\leq\ell_{4}$ and for sockets we require $\ell_{1}<\ell_{2}\leq\ell_{3}<\ell_{4}$ and we call a plug or a socket degenerate if it has at least one pair of repeated levels, i.e., for some $i\in[3]$, $\ell_{i}=\ell_{i+1}$.
-We create the degenerate plugs and sockets by contracting the edges between vertices of the repeated levels while keeping the vertex with the lower index; see Fig.
-2 and 2.
+We create the degenerate plugs and sockets by contracting the edges between vertices of the repeated levels while keeping the vertex with the lower index; see Figures 2(c) and 2(d).
 Let $P$ be a plug, and let $\mathrm{com}(P)$ be the connected component of $P$ in $G$.
-A plug fits into a socket if the gadgets could be “weaved” as illustrated in Fig.
-2.
+A plug fits into a socket if the gadgets could be “weaved” as illustrated in Figure 2(e).
 Formally, we say an $(a_{1},a_{2},a_{3},a_{4})$-plug $P$ fits into a $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $\min_{v\in\mathrm{com}(P)}\{\gamma(v)\}\leq s_{1}$, $\max_{v\in\mathrm{com}(P)}\{\gamma(v)\}\geq s_{4}$, and $s_{1}<a_{2}<s_{2}\leq s_{3}<a_{3}<s_{4}$.
-In an ordered level planar drawing of $\mathcal{G}$, we say a plug $P$ links to an $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $P$ fits into $S$, and $P$ is drawn between the connecting vertices of $S$ (that is, the edges of $P$ traversing level $s_{3}$ are to the right of $v_{1}$ and the edges of $P$ traversing level $s_{2}$ are to the left of $v_{9}$); see Fig.
-2.
+In an ordered level planar drawing of $\mathcal{G}$, we say a plug $P$ links to an $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ when $P$ fits into $S$, and $P$ is drawn between the connecting vertices of $S$ (that is, the edges of $P$ traversing level $s_{3}$ are to the right of $v_{1}$ and the edges of $P$ traversing level $s_{2}$ are to the left of $v_{9}$); see Figure 2(e).
 A defining feature of our constructions is a division into vertical strips, which is accommodated by the following notion.
@@
 Note that each gadget type essentially has a unique drawing in the sense that it corresponds to an ordered level planar graph for which all of its ordered level planar drawings have the same level planar embedding.
-This and further gadget properties are detailed in Appendix B.
-We use Lemma 3.1 to design specific plugs that can or cannot link to specific sockets in the same ordered level planar drawing.
-[…]
+We use Lemma 4 to design specific plugs that can or cannot link to specific sockets in the same ordered level planar drawing.
+Lemma 4.
 Consider an ordered level graph $\mathcal{G}$ that contains an $(a_{1},a_{2},a_{3},a_{4})$-plug $A$, a $(b_{1},b_{2},b_{3},b_{4})$-plug $B$, and an $(s_{1},s_{2},s_{3},s_{4})$-socket $S$ that occupy three disjoint sets of levels.
 There is an ordered level planar drawing of the subgraph of $\mathcal{G}$ spanned by $A$, $B$, and $S$ where $A$ and $B$ link to $S$ if and only if both $A$ and $B$ fit into $S$ and $a_{2}<b_{2}$ and $a_{3}<b_{3}$, or, both vice versa, $a_{2}>b_{2}$ and $a_{3}>b_{3}$.
-Proof 3.2 (Proof (sketch)).
+Proof.
 To distinguish the vertices of the plugs $A$ and $B$, we use the name of the plug as superscript, so, e.g., we refer to $u_{3}^{A}$ and $u_{3}^{B}$.
-If the conditions are satisfied, we can draw both plugs as depicted in Fig.
-3.
-Towards a contradiction, assuming the conditions are not satisfied, we observe that the drawing must locally “nest” plug $B$ into plug $A$ while drawing them around the edge $v_{4}v_{6}$ of $S$, at the same time $B$ “nests” into $A$ while going around the edge $v_{7}v_{5}$ of $S$; see Fig.
-3.
-The first nesting implies that the edge $u^{B}_{3}u^{B}_{4}$ is to the left of $u^{A}_{3}u^{A}_{4}$ and the second nesting implies the opposite order – a contradiction.
-{subfigure}
-[…]
-Figure 3: Sketches for the proof of Lemma 3.1. The edges of the socket $S$ are bold.
+As we have only one socket, we use just $v_{1},\dots,v_{10}$ for the vertices of $S$.
+For the names of the vertices, we assume that none of the involved gadgets is degenerate.
+If some of the gadgets are degenerate, adjusting the vertex names is straight-forward.
+(a) If the order of the layers follows the pattern $a_{2},b_{2},b_{3},a_{3}$, then there is no crossing-free drawing that has both $A$ and $B$ linked to $S$.
+(b) If the order of the layers follows the pattern $a_{2},b_{2},a_{3},b_{3}$, then there is a crossing-free drawing that has both $A$ and $B$ linked to $S$.
+Figure 3: Sketches for the proof of Lemma 4. The edges of the socket $S$ are bold.
+For the “only if”-direction.
+Let an ordered level planar drawing where both $A$ and $B$ link to $S$ be given.
+By the definition of linking, $A$ and $B$ fit into $S$ and, hence, $a_{2},b_{2}<s_{2}\leq s_{3}<a_{3},b_{3}$.
+Suppose for a contradiction that $a_{2}<b_{2}$ and $a_{3}>b_{3}$; see Figure 3(a) for an illustration.
+(The case $a_{2}>b_{2}$ and $a_{3}<b_{3}$ is identical up to renaming $A$ and $B$.) By the definition of linking and due to the two (uncrossed) paths of $S$, the edges $u_{2}^{A}u_{3}^{A}$ and $u_{2}^{B}u_{3}^{B}$ are to the left of $v_{4}$ and $v_{6}$, while the edges $u_{3}^{A}u_{4}^{A}$ and $u_{3}^{B}u_{4}^{B}$ are to the right of $v_{4}$ and $v_{6}$.
+Given $a_{2}<b_{2}$, it follows that this is possible if and only if $u_{2}^{B}$, $u_{3}^{B}$, and $u_{4}^{B}$ lie (horizontally) between the edges $u_{2}^{A}u_{3}^{A}$ and $u_{3}^{A}u_{4}^{A}$.
+Symmetrically, the edges $u_{3}^{A}u_{4}^{A}$ and $u_{3}^{B}u_{4}^{B}$ are to the left of $v_{7}$ and $v_{5}$, while the edges $u_{4}^{A}u_{6}^{A}$ and $u_{4}^{B}u_{6}^{B}$ are to the right of $v_{7}$ and $v_{5}$.
+Given $a_{3}>b_{3}$, it follows that this is possible if and only if $u_{3}^{B}$, $u_{4}^{B}$, and $u_{6}^{B}$ lie between the edges $u_{3}^{A}u_{4}^{A}$ and $u_{4}^{A}u_{6}^{A}$.
+Therefore, $u_{3}^{B}$ and $u_{4}^{B}$ lie at the same time to the left and to the right of $u_{3}^{A}u_{4}^{A}$, which is a contradiction.
+For the “if”-direction.
+Assume that $s_{1}<a_{2}<b_{2}<s_{2}<s_{3}<a_{3}<b_{3}<s_{4}$; see Figure 3(b) for an illustration.
+The case $s_{1}<b_{2}<a_{2}<s_{2}<s_{3}<b_{3}<a_{3}<s_{4}$ is identical up to renaming $A$ and $B$.
+We can get an ordered level planar drawing where both plugs $A$ and $B$ are to the right of $v_{1}$, and $A$ and $B$ are to the left of $v_{9}$ if we place the edges occurring between the involved layers in the following order from left to right: first $v_{1}v_{2}$ and $v_{2}v_{3}$, then $u_{2}^{A}u_{3}^{A}$, then $u_{2}^{B}u_{3}^{B}$, then $v_{4}v_{6}$ and $v_{6}v_{8}$, then $u_{3}^{B}u_{4}^{B}$, then $u_{3}^{A}u_{4}^{A}$, then $v_{3}u_{5}$ and $v_{5}v_{7}$, then $u_{4}^{A}u_{6}^{A}$, then $u_{4}^{B}u_{6}^{B}$, and finally $v_{8}u_{10}$ and $v_{10}v_{9}$.
+∎
+In an (ordered) level planar drawing of $\mathcal{G}$, we define spatial relations between two gadgets in the following way.
+For a gadget $X$, we say a level $\ell$ (or a gadget $Y$) is above $X$ if the level $\ell$ (each level hosting vertices of $Y$) has a larger index than any level hosting a vertex of $X$.
+Symmetrically, we define below for the levels with a smaller index.
+We say a gadget $Y$ is left (resp.
+right) of $X$ if on every level all vertices and edges of $Y$ lie to the left (resp.
+right) of all the vertices of $X$.
+We connect plugs and sockets with other gadgets using their connecting vertices.
+The spatial relation of two gadgets is usually clear, so we shortcut the description of how the connection is formed in the following way.
+Connecting a socket $S$ to a wall $W$ is done by identifying a vertex $x$ of $S$ with the vertex of $W$ on level $\gamma(x)$.
+The $x=v_{1}$ if $W$ is left of $S$ and $x=v_{5}$ if $W$ is right of $S$.
+Connecting a plug $P$ to a gadget $Y$ means adding an edge from a connecting vertex $z$ of $P$ to $Y$.
+We $z=u_{1}$ if $Y$ is above $P$ and $z=u_{5}$ if $Y$ is below $P$.
+The other endpoint of the edge depends on $Y$.
+In our case, $Y$ is either a plug or a single vertex.
+In case $Y$ is a plug its connecting vertex is determined in the same way we did it for $P$.
+Hence, the connection of $P$ to $Y$ results in an edge between $u_{1}$ of $P$ and $u_{5}$ of $Y$, or vice-versa.
+See connecting vertices of plugs and sockets in Figure 2.
+An ordered level planar graph is called rigid if all of its ordered level planar drawings have a common level planar embedding.
+Recall that $\mathcal{G}$ is proper if $\gamma(v)=\gamma(u)+1$ for every edge $(u,v)\in E(G)$.
+Observation 1.
+Every proper ordered level planar graph is rigid.
+Observation 2.
+A $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-plug or a $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$-socket correspond to a proper ordered level graph if and only if $\ell_{i+1}\in\{\ell_{i},\ell_{i}+1\}$ for every $i\in[3]$.
+By inserting a level $i$ into an ordered level graph $\mathcal{G}$, we mean that the number $h$ of levels of $\mathcal{G}$ increases by one and, for each vertex $v$ of $\mathcal{G}$ with $\gamma(v)\geq i$, its level $\gamma(v)$ increases by one.
+Observation 3.
+A rigid ordered level graph remains rigid even if we insert new levels.
+Moreover, a rigid graph $\mathcal{G}$ remains rigid even if we subdivide an edge $uv$ by a new vertex $w$, i.e., if we replace the directed edge $uv$ by the directed edges $uw$ and $wv$, and set the x-coordinate of $w$ according to where $uv$ crossed $L_{\gamma(w)}$ in the unique drawing of $\mathcal{G}$.
+From Observations 1, 2, and 3, it follows that plugs, sockets, and walls are all rigid since we can start with a proper ordered level graph and then insert additional levels.
+We are now done with the auxiliary definitions.
+We first give the full proof of $W[1]$-hardness of OLP.
+Then we extend the construction to show XNLP-hardness of OLP.
+Last in this section, we show that the XNLP-hardness result also works for connected ordered level graphs.
 $W[1]$-hardness.
@@
 We may assume without loss of generality that in $H$, there is no edge whose endpoints have the same color $C_{j}$.
-[…]
+Lemma 5.
 Ordered Level Planarity is $W[1]$-hard with respect to the height.
-Proof 3.4 (Proof (sketch)).
-We give an overview of a parameterized reduction from MCIS to OLP; refer to Fig.
-4 for an illustration.
-The ordered level graph constructed in this reduction admits a grid-cell structure where the vertical strips are separated by walls and the horizontal strips represent related groups of levels.
-More precisely, we allocate a constant number of consecutive levels per color of the MCIS instance, to which we refer as a color band.
-Between each pair of neighboring color bands, we have a collision band.
-We group several consecutive vertical strips to obtain blocks.
-We call the leftmost and rightmost block choice blocks and the other blocks edge blocks.
-Edge blocks correspond bijectively to edges of the MCIS instance.
-In the grid cells of the color bands, we insert sockets and we add as many plugs as we have sockets there.
-No two plugs can link to the same socket, so all available sockets are occupied by one plug.
-Each plug is a separate connected component, which allows that plug to freely “choose” a socket within its color band where it fits.
-However, we have different types of plugs and sockets – the first type of plug, called high plug, fits only into the sockets of the choice blocks.
-In a drawing of our ordered level graph, the number of high plugs that link to the left (and not to the right) choice block within a color band forces the second type of plugs, called color plugs, to be shifted further to the right.
-In the MCIS instance, a shift by $i$ corresponds to selecting the $i$-th vertex of the respective color to be in the solution.
-To obtain an independent set, we need to prevent that the shifts select both endpoints of an edge.
-To that end, we model an edge $uv$ as follows.
-The color plugs of one color band all occupy the same set of levels so they form a sequence specified in the ordered level graph.
-Because of this, we can identify, for a given shift, what socket a particular color plug links to.
-Consider the edge block of $uv$ and the color plugs that necessarily link to its sockets (a block has circa twice as many sockets than the maximal shift).
-For an illustration of the construction, assume that we have a shift in the color band of $u$ that selects $u$ and a shift in the color band of $v$ that selects $v$.
-We select a specific color plug of the color band of $u$ and one in the color band of $v$ to be extended so that they “reach” towards each other and end up in the same cell of a collision band.
-To this cell, we add a collision socket.
-If at most one extended plug ends up in a collision socket, then it can be drawn planarly.
-However, if both extended plugs end up in the same collision socket, then this cannot be drawn without crossings.
-Hence, there is an ordered level planar drawing of our ordered level graph if and only if there is a solution (i.e., an independent set) to the given MCIS instance.
+Proof.
+Let $\mathcal{S}$ be an instance of MCIS, that is, a graph $H$, an integer $k$, and a $k$-partition $C_{1},C_{2},\dots,C_{k}$ of $V(H)$.
+We prove the theorem by constructing an OLP instance $\mathcal{G}$, with $h\in\mathcal{O}(k)$ levels, that has an ordered level planar drawing if and only if $\mathcal{S}$ contains a multicolored independent set of size $k$.
+Let $n=|V(H)|$ and $m=|E(H)|$.
+For a color $j\in[k]$, let $m_{j}=|\{uv\colon c(u)=j\lor c(v)=j\}|$, i.e., $m_{j}$ is the number of edges incident to a vertex with color $j$.
+Note that $\sum_{j=1}^{k}m_{j}=2m$ by double counting since we assume that the colors of the endpoints of every edge are distinct.
+For this proof, we also assume that, without loss of generality, the color classes have the same size $n^{\prime}$, i.e., $n^{\prime}=|C_{j}|=n/k$ for each $j\in[k]$.
 Figure 4: Example of our parameterized reduction from Multicolored Independent Set (MCIS) to Ordered Level Planarity.
@@
 Here, the solution $\{v_{2},v_{3},v_{6},v_{7},v_{8}\}$ for the MCIS instances is found.
+Bands, columns, blocks, and cells.
+We divide the construction vertically and horizontally in the following way, see Figure 4.
+The levels are partitioned by their purpose into five types – rigid, color, high, pass-through, and collision levels.
+We use (horizontal) bands to denote consecutive sets of levels that belong together conceptually; we use two band types called color bands and collision bands.
+We consider any vertices that lie on levels of a band and any edges between such vertices to be part of the band.
+Note that bands contain a mixture of level types.
+We typically use relative numbering to speak about the levels of a band, e.g., the lowest level of a band is numbered $1$.
+The vertical slices of the construction are delimited by the walls.
+We begin the construction with $37\cdot k+3\cdot(k-1)=10k-$ 3 rigid levels.
+We group these rigid levels into $k$ color bands that alternate with $k-1$ collision bands: Each color band contains $7$ rigid levels and each collision band contains $3$ rigid levels.
+Let the color-$j$ band be the $j$-th color band from the bottom and let it be associated with color $j$.
+Similarly, the collision-$j$ band denotes the $j$-th collision band from the bottom, i.e., the collision band between the color-$j$ and the color-$(j+1)$ band.
+We later enlarge the bands by inserting more levels using Observation 3.
+We add $1+2\cdot(n^{\prime}-1)+m\cdot(2n^{\prime}-1)=(m+1)\cdot(2n^{\prime}-1)$ walls that divide each level created so far into $m\cdot(2n^{\prime}-1)+2n^{\prime}-2$ line segments and two rays.
+We order the line segments naturally by $\prec_{\ell}$ for each level $\ell$.
+For $i\in[m\cdot(2n^{\prime}-1)+2n^{\prime}-2]$, we let the $i$-th column be the union of the $i$-th line segments over all levels, vertices placed on those line segments, and edges between those vertices; we will not use the rays.
+The first and the last $(n^{\prime}-1)$ columns on the left/right are called left/right choice block.
+We partition the remaining $m\cdot(2n^{\prime}-1)$ columns into $m$ edge blocks each with $12n^{\prime}-$ 1 consecutive columns.
+Every edge block is associated with one edge of $E(H)$.
+Let the edge-$uv$ block be the edge block associated with the edge $uv\in E(H)$.
+We typically use relative numbering to denote columns in a block; so the leftmost column would be the first one etc.
+We call the intersection of a band and a column a cell.
+So far, the construction contains many empty cells.
+The aim is to fill each color band cells with plugs and sockets of various types to create a “shifting” mechanism that represents a selection of a vertex of that color.
+The collision bands come into play later to prevent a selection of neighboring vertices.
+Placement of sockets.
+Now we place the sockets into the cells.
+Let us call the two walls on the boundary of a cell its left and right wall.
+By adding a socket to a cell we mean adding the socket to $\mathcal{G}$ and identifying its right connecting vertex with the right wall vertex on the same level and doing the same for the left one.
+See Figure 5 to see sockets placed in cells.
+Recall that color bands contain $7$ rigid levels.
+We add a $(2,4,4,6)$-socket, see Figure 2(d), to each cell in all color bands within the choice blocks (note that we use relative level indices).
+We call these choice sockets.
+For every $uv\in E(H)$ we add $(3,4,4,5)$-sockets, called color sockets, to all cells that are in the edge-$uv$ block and in either color-$c(u)$ band or the color-$c(v)$ band.
+To all cells in the color-$i$ bands in the edge-$uv$ block where $i$ is different from $c(u)$ and $c(v)$, we add $(1,2,6,7)$-sockets, called pass-through sockets.
+Recall that collision bands contain three rigid levels.
+We do not add any sockets to the collision band cells that are in the choice blocks.
+In each edge block, we add a single socket as follows.
+For $uv\in E(H)$, let $q_{uv}=\min\{c(u),c(v)\}$; note that $q_{uv}<k$.
+Into the collision-$q_{uv}$ band of edge-$uv$ block, we add to a $n^{\prime}$-th cell a $(1,2,2,3)$-socket, which we call collision socket.
+See construction overview in Figure 4.
+Figure 5: Plugs and sockets drawn on color band levels. Sockets are placed in cells. (a) color plug, (b) high plug, (c) pass-through plug, (d) choice socket, (e) color socket, (f) pass-through socket.
+Refinement.
+Next, we make the gadgets proper by finding edges that span more than one level and subdividing them by vertices on all the intermediate levels, setting the x-coordinates of the new vertices according to the unique drawings of the gadgets.
+By Observation 1 the construction is rigid and by Observation 3 it will remain rigid even if we insert further levels.
+This completes the part of the construction on rigid levels.
+Now we add the other types of levels, which host the plugs.
+To each color band, where we already have $7$ rigid levels (for now denoted by R), we insert $4$ color levels (C), $4$ high levels (H), and $2$ pass-through levels (P) so that the final order from bottom to top is HCRPRHRCRCRHRPRCH; see table 1.
+Then we subdivide the leftmost and the rightmost wall to contain vertices on all levels – forcing everything else to be drawn between them.
+We refer to the levels by their relative order within the band.
+Note that the level insertion changed the indices of rigid levels and of the gadgets they contain.
+level 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17
+type H C R P R H R C R C R H R P R C H
+Table 1: Order of the levels within each color band.
+H stands for high levels hosting the high plugs, C for color levels hosting the color plugs, R for rigid levels hosting the rigid structure, and P for pass-through levels.
+Placement of plugs.
+To each of the $k$ color bands, we now add a set of plugs that will guide the mechanism for selecting vertices in the MCIS instance $\mathcal{S}$.
+None of these plugs are connected to the rest of the graph, i.e., each plug is an isolated connected component of $\mathcal{G}$.
+In the color-$j$ band, we add as many plugs as there are choice and color sockets in total (so we do not include pass-through sockets here): Recall that the color-$j$ band contains in total $(2n^{\prime}-2)+m_{j}\cdot(2n^{\prime}-1)$ choice and color sockets.
+We add $n^{\prime}-1$ of $(1,6,12,17)$-plugs, which we call high plugs, and $n^{\prime}-1+m_{j}\cdot(2n^{\prime}-1)$ of $(2,8,10,16)$-plugs, which we call color plugs.
+See Figure 5a and 5b.
+We place the high plugs from left to right one after another so that their x-coordinates do not overlap, similarly we place the color plugs.
+Shifting mechanism (choosing vertices).
+Observe that color plugs can fit into choice sockets and color sockets while high plugs can fit only into choice sockets.
+Further, neither color plugs nor high plugs can fit into pass-through sockets.
+The leftmost and rightmost walls of $\mathcal{G}$ force the plugs to be drawn into the cells while a high plug and a color plug cannot occupy the same socket due to Lemma 4.
+Hence, in every realization of $\mathcal{G}$ every choice and color socket is linked with exactly one plug.
+When we say that we shift color $j$ by $i$ steps, we mean that a realization of $\mathcal{G}$ links $i$ high plugs in the left choice block and $(n^{\prime}-1)-i$ high plugs in the right choice block of the color-$j$ band.
+The high plugs can be interspersed among color plugs in the choice blocks arbitrarily.
+Shifting color $j$ by $i$ represents selecting the $(i+1)$-th vertex of color $j$; see Figure 6.
+As we have $n^{\prime}-1$ choice sockets in the left choice block, the represented vertex choice is in $[n^{\prime}]$.
+Let $A$ be a color-$j$ plug that links to the $t$-th color socket of an edge block when color $j$ is shifted by 0 steps.
+Recall that the relative order of the color plugs is fixed because they occupy the same set of levels.
+Hence, with shift $i$ plug $A$ links to the $(t+i)$-th color socket of the edge block.
+Note that shift $i$ forces $i$ of the last color-$j$ plugs in the edge block overflow to the next block.
+Figure 6: An example of a color band in our $W[1]$-hardness reduction for color classes of size 4 with high-plugs (h), color-plugs (c), pass-through-plugs (p), and the rigid structure (r).
+Note how two high-plugs within the choice-block force all the color-plugs to be shifted by two columns to the right.
+Further note the special role of the dashed color-plug that extends to the collision-gadget beyond these levels.
+Collisions (ensuring independence).
+Next, we make sure that for each $uv\in E(H)$, the vertices $u$ and $v$ cannot be selected at the same time.
+Intuitively, we achieve this by extending two color plugs within the edge-$uv$ block so that they end up occupying the same collision socket, where at most one can fit – making the shifts that represents selecting $u$ and $v$ impossible to draw.
+Recall that to the edge-$uv$ block we already placed a collision socket into $n^{\prime}$-th cell of collision-$q_{uv}$ band, where $q_{uv}=\min\{c(u),c(v)\}$.
+Now to each collision band (whose three rigid levels are denoted by R), we insert two types of levels – one called above levels (A) and the other called below levels (B).
+The resulting sequence of levels within each collision band is ARABRBARB, also noted in Table 2.
+We add $(1,3,7,7)$-plugs called A-collision plugs and $(4,4,6,9)$-plugs called B-collision plugs.
+We set their x-coordinates to be in the middle cell but due to the choice of levels the collision plugs are not restricted to be drawn in a particular cell.
+Note that with the inserted levels, each collision socket is now a $(2,5,5,8)$-socket.
+level 1 2 3 4 5 6 7 8 9
+type A R A B R B A R B
+Table 2: Order of the levels within each collision band.
+A stands for above levels hosting the A-collision plugs, B for below levels hosting the B-collision plugs, and R for rigid levels hosting the rigid structure.
+For this description let $c(u)<c(v)$, without loss of generality by swapping $u$ and $v$.
+Let $\operatorname{idx}(u)$ and $\operatorname{idx}(v)$ denote the index of the vertices $u$ and $v$ within their color, i.e., vertex $u$ is the $\operatorname{idx}(u)$-th vertex of color $c(u)$ and $v$ the $\operatorname{idx}(v)$-th vertex of color $c(v)$.
+Let $X$ be the $(n^{\prime}-\operatorname{idx}(u)+1)$-th color plug in the color-$c(u)$ band of the edge-$uv$ block and let $Y$ be the $(n^{\prime}-\operatorname{idx}(v)+1)$-th color plug in the color-$c(v)$ band of the edge-$uv$ block.
+Finally, we describe how to connect $X$ and $Y$ to the collision plugs.
+We connect $X$ to the B-collision plug, this is straight-forward as they occupy neighboring bands.
+However, to connect $Y$ with the A-collision plug the connection needs to traverse over many color bands.
+To accommodate that, we insert $(4,4,14,14)$-plugs called pass-through plugs, see Figure 5c, into each color band between the collision-$c(u)$ band and the color-$c(v)$ band.
+We place the pass-through plugs in the middle cell of the edge-$uv$ block.
+We connect the A-collision plug in the collision-$c(u)$ band with the pass-through plug in the color-$(c(u)+1)$ band, then connect pass-through plug in color-$i$ band with the pass-through plug in color-$(i+1)$ band for all $i\in\{c(u)+1,\dots,c(v)-2\}$, and last we connect the pass-through plug in color-$(c(v)-1)$ band with $Y$.
+We refer to the two connected components of $X$ and $Y$ as the (pair of) extended color plugs of $uv$.
+Note that the pass-through plugs fit only into pass-through sockets.
+Claim 4.
+Let $uv\in E(H)$. If $u$ is selected in color $c(u)$ and $v$ is selected in color $c(v)$, the pair of extended color plugs in the edge-$uv$ block lies in the $n^{\prime}$-th column of the block.
+Proof.
+Since color $c(u)$ is shifted by $\operatorname{idx}(u)-1$ steps to the right, the corresponding extended color plug is in column $n^{\prime}+1-\operatorname{idx}(u)+(\operatorname{idx}(u)-1)=n^{\prime}$ of the edge-$uv$ block.
+Similarly, for the corresponding extended color plug on the color-$c(v)$ band shifted by $\operatorname{idx}(v)-1$, we have $n^{\prime}+1-\operatorname{idx}(v)+(\operatorname{idx}(v)-1)=n^{\prime}$.
+∎
+We finish the construction of $\mathcal{G}$ by performing the described way of adding a pair of extended color plugs for each edge of $H$.
+Correctness.
+We started off by adding a rigid structure that divides the ordered level planar drawing into cells.
+Then we inserted extra levels that host various plugs and allow the plugs to flow freely between cells.
+The bounding walls occupy all levels and so restrict every plug to be drawn between them.
+We added sockets to cells that restrict the number of plugs in a cell to one.
+We modelled selection of vertices in MCIS by shifting a set of color plugs within the color band.
+Then we ensured that the selected vertices must be independent by creating collision via extended color plugs.
+We finish the proof by arguing correctness.
+Claim 5.
+For $uv\in E(H)$, there is no ordered level planar drawing of $\mathcal{G}$ where the color-$c(u)$ band is shifted by $\operatorname{idx}(u)-1$ and the color-$c(v)$ band is shifted by $\operatorname{idx}(v)-1$.
+Proof.
+Towards a contradiction, assume that the color plugs in the color-$c(u)$ band are shifted by $\operatorname{idx}(u)-1$ and the color plugs in the color-$c(v)$ band are shifted by $\operatorname{idx}(v)-1$.
+Then by Claim 4 the extended plugs of $uv$ end up in the $n^{\prime}$-th column – the same column that has a collision socket of edge-$uv$ block in the collision-$c(u)$ band.
+The two extended color plugs form connected components to the collision-$c(u)$ band so in an ordered level planar drawing the extended parts must stay in the same column as they cannot cross rigid walls.
+We also see that A-collision plug and B-collision plug cannot both be drawn in the same collision socket due to Lemma 4.
+Hence, such a drawing does not exist.
+∎
+Claim 6.
+There exists an ordered level planar drawing of $\mathcal{G}$ if and only if $\mathcal{S}$ is a yes-instance.
+Proof.
+For the one direction, assume $\mathcal{S}$ is a yes-instance.
+Let $s_{j}$ be the index of a vertex of color $j$ that is chosen in the solution $\mathcal{S}$.
+We draw the rigid part of $\mathcal{G}$ in its unique embedding.
+Then we draw high and color plugs of color-$j$ band shifted by $s_{j}-1$.
+All types of plugs occupy disjoint sets of levels, hence, drawing one type does not influence the possibility of drawing the other.
+In $\mathcal{S}$ we have, for each edge, chosen at most one incident vertex.
+So in each edge-$uv$ block at most one extended plug links to the collision socket of $uv$.
+In the other direction, assume we have an ordered level planar drawing of $\mathcal{G}$.
+Focusing on the rigid part we note it still has only one embedding, which must be used in this drawing as well.
+This naturally divides the final drawing into columns and bands.
+The leftmost and rightmost walls force all the remaining gadgets to be drawn in between them, placing each plug in some column.
+For each color $j$ we can identify the number $s_{j}$ of high plugs in the color-$j$ band left choice block.
+Due to Claim 5, we know that these shifts cannot represent adjacent vertices in $\mathcal{S}$ and, hence, correspond to a solution for $\mathcal{S}$.
+∎
+MCIS is $W[1]$-hard with respect to the solution size $k$.
+Our construction uses $917\cdot k+9\cdot(k-1)=26k-$ 9 levels and can clearly be carried out in polynomial time.
+Hence, Ordered Level Planarity is $W[1]$-hard with respect to the number of levels.
+∎
 XNLP-hardness.
 We now extend our construction for $W[1]$-hardness to obtain XNLP-hardness as well.
-We design a parameterized tractable log-space reduction [6, Section V.B.] – a parameterized reduction that runs in $\mathcal{O}(g(k)\cdot n^{c})$ time and uses only $\mathcal{O}(f(k)\cdot\log n)$ space for internal computation, where $g$ and $f$ are computable functions, $k$ is the parameter, and $n$ is the input size.
+We design a parameterized tractable log-space reduction [7, Section V.B.] – a parameterized reduction that runs in $\mathcal{O}(g(k)\cdot n^{c})$ time and uses only $\mathcal{O}(f(k)\cdot\log n)$ space for internal computation, where $g$ and $f$ are computable functions, $k$ is the parameter, and $n$ is the input size.
 It suffices to design the reduction such that it returns a single bit of the output on demand.
@@
 The task is to select an independent set $X\subseteq V(G)$ such that, for each $i\in[r]$ and for each color $j\in[k]$, $|X\cap V_{i}\cap C_{j}|=1$.
-[…]
+Theorem 4.
 Ordered Level Planarity is XNLP-hard with respect to the height.
-Proof 3.6 (Proof (sketch)).
-The main idea in our reduction from CMCIS to OLP is to use $r$ instances of the $W[1]$-hardness construction from Theorem 3.3.
-We have one instance for each subgraph induced by $V_{i}$ (for $i\in[r]$) where additionally the edges connecting $V_{i}$ with $V_{i-1}$ and $V_{i+1}$ (if existent) are represented by edge blocks shared by the corresponding two instances; see Fig.
-5.
-To this end, we arrange all instances in a zigzag pattern in the top and the bottom half of the combined construction.
-Now to make it a parameterized tractable log-space reduction, which outputs bit by bit, we essentially traverse the grid structure with a vertical sweep-line from left to right (i.e., column by column) and save as the status only the $\mathcal{O}(k)$ information belonging to the current column and the $\mathcal{O}(\log n)$ information telling us in which column we are currently globally.
-We describe this in more detail in the appendix.
+Proof.
+Let $\mathcal{S}$ be an instance of the Chained Multicolored Independent Set (CMCIS).
+We prove this theorem by constructing an OLP instance $\mathcal{G}$, with $h\in\Theta(k)$ levels, that has an ordered level planar drawing if and only if there exists a solution to $\mathcal{S}$.
+Then we argue how to retrieve a single output bit using limited memory.
+The instance we build uses the structures that we have described in detail in the proof of Lemma 5, namely (different types of) plugs, sockets, color/collision bands, the whole mechanism that represents the choice of a vertex in the solution by the shifts of plugs within a color band, and the collision mechanism that makes it impossible to use shifts that would represent choosing neighboring vertices.
+We go through the construction while assuming familiarity with the aforementioned terms.
+For $i\in[r]$, we denote the edges between vertices of $V_{i}$ by $E_{i}$, and we denote the edges that connect a vertex from $V_{i}$ with a vertex from $V_{i+1}$ by $E_{i,i+1}$.
+Moreover, we let $n_{i}=|V_{i}|$, $m_{i}=|E_{i}|$, and $m_{i,i+1}=|E_{i,i+1}|$.
+For the reduction, we may assume that all of the $r$ partitions have the same size $n_{i}$, for each $V_{i}$ there are $n_{i}/k=:\hat{n}$ vertices per color, and each $E_{i}$ contains at least one edge, i.e., $m_{i}\geq 1$.
+If this was not the case, we could simply add vertices (with a specific color $j$) until all partitions have $n_{i}$ vertices and $\hat{n}$ vertices per color.
+The new vertices get an edge to every vertex in its partition except for the vertices of color $j$.
+Hence, they cannot appear in the solution if there are two or more colors.
+The number of vertices being added is in $\mathcal{O}(k\cdot r\cdot\max_{i\in r}n_{i})$.
+Roughly speaking, the whole construction combines multiple times the $W[1]$-hardness constructions – partially overlapping horizontally.
+More precisely, we first create $2k$ color bands that alternate with $12k-$ 1 collision bands.
+As before, each color band starts with $7$ rigid levels and each collision band starts with $3$ rigid levels.
+We divide the color bands into two halves; we call the $k$ color bands that occupy the lower levels bottom half and the other $k$ color bands top half; see Figure 7 for an illustration.
+Then, we add $1+2\cdot(\hat{n}-1)+m\cdot(2\hat{n}-1)$ walls where $m$ is the total number of edges, i.e., $m=\sum_{i=1}^{r}m_{i}+\sum_{i=1}^{r-1}m_{i,i+1}$.
+The walls divide the final drawing into columns, each column in a band forms a cell.
+Several columns form a block, however, we will also use blocks that span only one half.
+In such a case, the construction fills all unused cells with pass-through sockets.
+The first and last $(\hat{n}-1)$ columns form the first and the last choice block.
+The remaining columns are divided into $m$ edge blocks each with $(2\hat{n}-1)$ columns.
+These edge blocks are partitioned into sectors.
+Going from left to right, we call the first $m_{1}$ edge blocks the $(1)$-sector, which represent the edges in $E_{1}$, the next $m_{1,2}$ blocks the $(1,2)$-sector, which represent the edges in $E_{1,2}$, and so on, alternating between $E_{i}$ and $E_{i,i+1}$ until $E_{r}$ is reached.
+For $i\in[r]$, let $V_{i}$ be assigned to the top half if $i$ is odd, and to the bottom half otherwise.
+The colors of the vertices in $V_{i}$ are represented by the $k$ color bands in its assigned half.
+We now add all the remaining parts.
+For each edge block, we add color, pass-through, and collision sockets accordingly as in the proof of Lemma 5.
+For edges in $E_{i,i+1}$, we put the color sockets in one of the top-half color bands and in one of the bottom-half color bands.
+Note that an edge block representing an edge that is incident to a vertex from $V_{i}$ lies in the $(i-1,i)$-, $(i)$-, or $(i,i+1)$-sectors.
+For each $i\in[r]$, we place two choice blocks in the half where $V_{i}$ is assigned to: one in the $\hat{n}-1$ columns just before the $(i-1,i)$-sector and one just after the $(i,i+1)$-sector.
+Last, we insert the non-rigid levels into each color and collision band as before.
+We subdivide the very first and last wall to have vertices on all levels.
+For each $i\in[r]$, we also subdivide the first wall of the left choice block of $V_{i}$ in the half of $V_{i}$ and do the same for the last wall of the right choice block.
+The resulting construction has $r$ times (some instance of) the $W[1]$-hardness construction lying in the top or the bottom half, where the only difference is that some of the extended color plugs are shared between two of these constructions in different halves.
+Note that this does not influence the shifting mechanism; it still works the same way.
+Each edge block internally looks the same, the ones for $E_{i}$ just span fewer levels than those for $E_{i,i+1}$.
+Hence, the edge block arguments from Claims 4 and 5 work the same and we can derive a statement analogous to Claim 6.
+Recall that for a parameterized tractable log-space reduction, we need to specify how to retrieve from the input CMCIS instance in FPT-time a single bit of the output OLP instance while maintaining only $\mathcal{O}(f(k)\cdot\log n)$ space across all requests of a bit.
+We traverse our grid structure column by column.
+To locate the plugs somewhere (for the reduction, we just need to represent the OLP instance not a specific drawing of it), we assume that all high plugs are assigned to the right choice block.
+This gives every color, pass-through, and collision plug its fixed column.
+Therefore, we need to save as our status only the information belonging to the current column, e.g., to which edge (if any) this column belongs etc.
+Every column spans $\mathcal{O}(k)$ levels and every socket has $\mathcal{O}(1)$ vertices and edges.
+Hence, the total number of vertices and edges in one column is $\mathcal{O}(k)$ and what type they are can be computed in polynomial time and $\mathcal{O}(k\log n)$ space where the $\log n$ factor comes from saving the number of the currently considered column, vertex and edge.
+When we need to proceed with the next column, we can find the next edge in polynomial time if needed, which means we do not need to first sort all edges.
+Moreover, if our output encoding expects first all vertices, then all edges, etc., our algorithm might do several iterations through our column structure when considering all requests of single bits together.
+∎
 Figure 7: Comparison of the $W[1]$-hardness construction (left) and the XNLP-hardness construction (right) with $r=3$ and $k=5$.
@@
 Vertical bars and orange disks between color bands represent the extended plugs and the collision sockets, respectively, which are used for the collision mechanism.
-The construction in the proof of Theorem 3.5 can be altered to make the graph connected as we show at the end of Appendix B.
-The main idea is to add a new vertex $q$ on top of the construction and connect all connected components to it.
-All sockets are already connected to walls and from walls it suffices to create edges to $q$ from their topmost vertex.
-For the plugs, we duplicate every column into a left and a right part such that each left part behaves as before, while each right part hosts paths that connect the “old” plugs to $q$.
-These paths are made of a sequence of a new kind of plugs.
-This comes at the expense of using more levels, namely $\Theta(k^{2})$ instead of $\Theta(k)$.
-This yields the XNLP-hardness statement in Theorem 1.1.
+The construction in the proof of Theorem 4 can be altered to make the graph connected as we show now.
+Theorem 5.
+Ordered Level Planarity is XNLP-hard with respect to the height, even when the input graph $G$ is connected.
+Proof.
+Recall that in the proof of Theorem 4 (and Lemma 5), we first create many walls, and then fill in sockets and plugs.
+The only gadgets that are disconnected from the rigid structure are the plugs.
+In short, our goal is to extend all plugs to connect them to a single vertex that is placed at a new topmost level.
+To this end, we add a new tunnel column to every column.
+A tunnel column contains sockets that are not much different from pass-through sockets, though they allow up to $3k$ plugs to fit at the same time.
+These sockets are placed in new bands, which are inserted inside the color bands.
+The new bands have their own levels for the different types of plugs because they need to be drawn mutually independent.
+Therefore, we use roughly $6k$ extra levels per color band, resulting in a total number of levels in $\Theta(k^{2})$.
+We now proceed to show the construction in more detail.
+We begin with a slightly altered construction from the proof of Theorem 4.
+We create the rigid levels and add the walls that form the columns.
+Then we split each column by a wall into two vertical strips – we refer to the right strip as tunnel, the left one we still call column.
+We continue with the construction from the proof of Theorem 4 as if the tunnels did not exist.
+After this construction is complete, we continue as follows.
+Observe that each connected component of the rigid structure contains a wall because every connected component of every socket is attached to a wall.
+We add a new topmost level and create a new vertex $q$ on it.
+We connect $q$ to all topmost vertices of all the walls making the rigid structure connected, see Figure 8d.
+Figure 8: Our modifications to make the ordered level graph in our XNLP-hardness reduction connected.
+Tunnels have a transparent blue background, while the tunnel band has a gray background.
+(a) Simplified color band with a color plug in a color socket; (b) column divided into the left column and a tunnel, showing how a plug continues in the tunnel; (c) a tunnel socket and three tunnel plugs within the tunnel band of a tunnel; and (d) the top level where the tunnel ends by connecting to $q$.
+Into each color band, we insert a tunnel band between the levels $9$ and $10$ (in the relative level numbering of the color bands), see Figure 8a and b.
+A tunnel band consists of $16k+$ 1 levels, which are grouped in a top and a bottom half.
+Each half contains $k$ triplets of levels.
+The $j$-th triplet is made of the rigid-$j$ level, color-$j$ level, and high-$j$ level, as shown in table 3.
+To close off the tunnel band, it contains an extra rigid level on top.
+Table 3: Purposes of the levels of a tunnel band: $\texttt{R}_{j}$ stands for rigid-$j$ level, $\texttt{C}_{j}$ for color-$j$ level, and $\texttt{H}_{j}$ for high-$j$ level.
+Furthermore, R stands for the rigid level on top.
+level $1$ $2$ $3$ $4$ $5$dots $13k-$ 1 $3k$ $13k+$ 1 $23k+$ 2 $\dots$ $6k$ $16k+$ 1
+type $\texttt{R}_{1}$ $\texttt{C}_{1}$ $\texttt{H}_{1}$ $\texttt{R}_{2}$ $\texttt{C}_{2}$dots $\texttt{C}_{k}$ $\texttt{H}_{k}$ $\texttt{R}_{1}$ $\texttt{C}_{1}$dots $\texttt{H}_{k}$ R
+Let a $(1,3k+1,3k+1,6k+1)$-socket with respect to a tunnel band be called tunnel socket.
+Let an $(i,i,i+3k,i+3k)$-plug be called tunnel-$i$ plug for $i\in\{2,\dots,3k\}$.
+By Lemma 4, any number of distinct tunnel plugs can fit into a tunnel socket, moreover, they occupy mutually disjoint sets of levels.
+For a vertex $v$ and $i\in\{2,\dots,3k\}$, let tunnel from $v$ to $q$ through $i$ mean adding a tunnel-$i$ plug to each color band between the color band of $v$ and the topmost vertex $q$, connecting $v$ to the bottommost of these tunnel-$i$ plugs, connecting pairs of tunnel-$i$ plugs in neighboring color bands, and connecting the topmost plug to $q$.
+In case $v$ is in the topmost color band, we just connect $v$ with $q$ with an edge.
+Our aim is to make the construction connected by tunneling from a vertex of every connected component to $q$.
+The remaining connected components are the plugs and extended plugs described in the proofs of Lemmas 5 and 1.
+Similar to the closing argument of Theorem 4, let us assume for the ease of description that all high plugs are placed in the right choice block.
+Let us focus on a single column-tunnel pair along with their plugs to describe the procedure that is performed on each such pair.
+In each color-$j$ band for $j\in[k-1]$ we take the first rigid vertex $u$ above the color-$j$ band in the wall between the column and its tunnel, we duplicate $u$ to create $u^{\prime}$, place $u^{\prime}$ to the right of $u$ and set the edge of the wall incident to $u$ from below to be connected to $u^{\prime}$ instead.
+This operation splits the wall into two parts and the lower part becomes disconnected from $q$; we fix this by tunnelling from $u^{\prime}$ to $q$ through $23j-$ 2.
+In each color-$j$ band for $j\in[k]$ within this column-tunnel pair, if the socket is assigned a high or a color plug, then we take the top vertex $v$ of the plug assigned the socket and tunnel from $v$ to $q$ through $13j-$ 1 if the plug is a color plug, or through $3j$ if it is a high plug.
+Now we argue that our alterations did not impact how the construction works.
+First, each tunnel band was inserted between two adjacent levels of a color band so it is evident that no color or high plugs can fit the tunnel socket, and no tunnel plugs can fit the choice or color sockets.
+We let the connections of the tunnel-$(3i-2)$ plugs, which we use to attach vertices of the walls to $q$, occupy the rigid levels so their positions within the rigid structure is fixed.
+Similarly, we let the connections of the tunnel-$(3i-1)$ plugs occupy levels that are reserved for color plugs, and we let the connections of the tunnel-$(3i)$ plugs occupy levels that are used only for high plugs.
+Hence, the set of feasible drawing does not change by adding tunnels and this connected instance of Ordered Level Planarity has a planar drawing if and only if the CMCIS instance has a solution.
+We argue that to retrieve a single output bit, we can still use the same approach as in the proof of Theorem 4.
+We again traverse our construction with a vertical sweep-line.
+For the argument, we again assume that all high plugs are located in the right choice blocks.
+This fixes not only color, pass-through, and collision plugs, but also the newly added tunnel plugs.
+The cornerstone observation is that even though we altered the construction, we know that in each column-tunnel pair, every level contains a constant number of vertices.
+Every column-tunnel pair spans $\mathcal{O}(k^{2})$ levels that in total contain no more than $\mathcal{O}(k^{2})$ vertices and edges.
+For a global orientation of the column-tunnel, $\mathcal{O}(n)$ bits suffices.
+To get one bit of the output, we can go over the columns-tunnel pairs one by one.
+∎
+Remark 1.
+We now hint at the reason why the bound of $\Theta(k^{2})$ levels might be necessary to make our construction connected.
+At first glance, it may seem feasible to restrict the number of levels in each tunnel socket to two.
+Every edge block contains at most two plugs in each column so why do we need $\Omega(k)$ levels in each tunnel socket?
+The choice blocks do contain many plugs in each column, but those could be spread out into separate choice blocks, each being relevant for only one color.
+The main issue is that even for the lowest numbered color we need to connect its plugs to the topmost vertex $q$.
+To do that we build a path over all other color bands.
+If we did not add extra levels to such bands, then choices for the lowest color would not have been independent of the upper colors.
+4 NP-Hardness of 4-Level Constrained Level Planarity
+We show the hardness of 4-level CLP by reducing from the $3$-Partition problem.
+In the $3$-Partition problem, we are given a multiset $S=\{s_{1},\dots,s_{n}\}$ of $n=3m$ positive integers whose sum is $mB$.
+The task is to decide whether there exist a partition of $S$ into $m$ triplets $S_{1},S_{2},\dots,S_{m}$ such that the sum of the numbers in each triplet is equal to $B$.
+$3$-Partition is known to be strongly NP-hard even when every integer in $S$ is strictly between $B/4$ and $B/2$ [19].
+This is important for us because we represent each number $k$ in $S$ in unary encoding by a $k$-clip, a graph structure that we define next; see Figure 9.
+The bounds on these numbers also play a role in ensuring that each triplet contains exactly three numbers.
+Definition 1 ($k$-Clip).
+A $k$-clip is graph that contains $12k+$ 1 edges that have one endpoint on level $2$ and alternate their second endpoint between levels $3$ and $1$ (starting with level $3$); the clip also contains a central vertex on level $4$ which is connected to all vertices on level $3$.
+The order of the level-$2$ vertices is fixed to a linear order by constraints.
+Figure 9: An example of a $4$-clip that represents number $4$. Red arrows show second-level constraints that fix the drawing of the clip.
+Lemma 6.
+If $\mathcal{G}$ contains two clips $P_{1}$ and $P_{2}$, then, in any $4$-level planar drawing $\Gamma$ of $\mathcal{G}$, the clips $P_{1}$ and $P_{2}$ do not cross, i.e., on every level either all vertices of $P_{1}$ lie before all vertices of $P_{2}$ or vice versa.
+Proof.
+First, consider the level planar drawing $\Gamma$ and let $v_{1}\in P_{1}$ and $v_{2}\in P_{2}$ be the unique level-$4$ vertices of these clips.
+We assume without loss of generality (by swapping names of $P_{1}$ and $P_{2}$) that $v_{1}$ is drawn to the left of $v_{2}$.
+We immediately see that on the levels $2$ and $3$ all vertices of the clip $P_{1}$ must be drawn before all vertices of the clip $P_{2}$.
+For the level-$1$ vertices of a clip it holds that they must be drawn in the same order as their adjacent vertices on level $2$.
+Therefore, the statement also follows for the level-$1$ vertices.
+∎
+For each $j\in[3m]$, the number $s_{j}\in S$ shall have its own $s_{j}$-clip and be forced to be drawn in a mountain chain, shown in Figure 10.
+Each mountain chain represents one of the final $m$ “buckets” of size $B$.
+An assignment of $k$-clips onto these mountain chains represents a partition of $S$ into triplets.
+Definition 2 (Mountain Chain).
+A mountain is a path on $5$ vertices $(v_{1},v_{2},v_{3},v_{4},v_{5})$ where $\gamma(v_{1})=1$, $\gamma(v_{2})=2$, $\gamma(v_{3})=3$, $\gamma(v_{4})=2$, and $\gamma(v_{5})=1$.
+A mountain $k$-chain is created by combining $k$ mountains in a chain by identifying, for each pair of consecutive mountains, $v_{5}$ of the first one with $v_{1}$ of the second one.
+After forming the chain, we add two walls that consist of a path from level $1$ to level $4$.
+The level-$1$ vertices of these walls are identified with the leftmost and rightmost level-$1$ vertices of the mountain chain.
+We can have sequences of mountain chains if we identify the second and the first wall of two mountain chains.
+We observe that each mountain $k$-chain contains $k+1$ valleys, one between each pair of adjacent mountains, and another between the outer mountains and the adjacent walls.
+Figure 10: A drawing of mountains in a sequence of two mountain chains.
+The mountain chains are made up of mountains glued together by their first-level vertices and with a wall from level $1$ to level $4$ on both ends.
+Observe that the valleys are formed between adjacent mountains and between the left-, respectively, rightmost mountain and the neighboring walls.
+Having all basic building blocks at hand, we are ready to prove the hardness of CLP.
+Theorem 6.
+Constrained Level Planarity is NP-hard for four levels.
+Proof.
+We describe a polynomial-time reduction from $3$-Partition.
+We are given integers $m$, $B$ and a multiset of integers $S$ of $n=3m$ positive integers such that their sum is $mB$.
+We create the CLP instance $\mathcal{G}=(G,\gamma,(\prec_{i})_{i\in[4]})$ as follows.
+For each $s_{j}\in S$ where $j\in[3m]$, create an $s_{j}$-clip.
+We also create a sequence of $m$ mountain $B$-chains $C_{1},\dots,C_{m}$.
+Let $w_{0},\dots,w_{m}$ denote (in order) the vertices of the walls on level 4.
+We add the constraint $w_{0}\prec_{4}w_{m}$ and, for each $s_{j}$-clip, we add the constraints $w_{0}\prec_{4}x$ and $x\prec_{4}w_{m}$, where $x$ is the level-4 vertex of the $s_{j}$-clip.
+First, let us prove that, given a Yes-instance of $3$-Partition, the constrained level planar obtained by our reduction admits a level planar drawing on four levels.
+We draw the mountain chain naturally from left to right in the ordering which satisfies $w_{0}\prec_{4}w_{m}$.
+Consider a solution of the $3$-Partition instance.
+For each triplet $(a,b,c)$ in that solution, we take the respective $a$-clip, $b$-clip, and $c$-clip, and draw them all within a single mountain $B$-chain.
+In particular, we take one clip and draw it such that the edges between the levels $1$ and $2$ occupy the spaces below a set of consecutive mountains, and the edges between the levels $2$ and $3$ occupy the valleys around these mountains.
+This way the $a$-clip fills $a$ of the mountains, the $b$-clip fills $b$ of them, and $c$-clip fills $c$ mountains.
+Together, their edges on the levels $1$ and $2$ occupy exactly the $B$ mountains of a mountain $B$-chain as $a+b+c=B$.
+See Figure 11 for a simplified example of such a drawing.
+The $m$ triplets fit into the $m$ mountain $B$-chains as described above.
+This drawing is clearly planar and satisfies all the constraints, concluding the first part of the proof.
+Figure 11: A level planar drawing of clips of size $2,2,1,1,3,1$ into mountain chains of size $5$.
+For clarity, we show the constraints that force clips between first and last wall only for a single clip.
+Second, we prove a series of claims showing that if there is a level planar drawing on four levels, then it must look exactly like we described above.
+Hence, such a planar drawable constrained level graph represents a Yes-instance of $3$-Partition.
+We start with the drawing of the mountain chains.
+Claim 7.
+There is a unique level planar drawing of the mountain $B$-chains $C_{1},\dots,C_{m}$.
+Proof.
+The mountains of the mountain chains together form of a path of length $m\cdot 4B+1$, which are arranged on the levels in order $1,2,3,2,1,\dots,2,1$ and that can be drawn in two ways – left-to-right or right-to-left.
+Then, as the walls go up to level $4$, they can only be drawn on the upper side of that long path, so there is a unique space for each wall to be drawn.
+Finally, due to the $w_{0}\prec_{4}w_{m}$, only one of the two ways to draw the path and the walls remains.
+∎
+Claim 8.
+In any level planar drawing, any mountain contains at most one edge (between levels 1 and 2) of any clip.
+Proof.
+There is a unique way to draw a clip because the constraints give a total order on its level-$2$ vertices.
+Notice that in this drawing every edge $e$ between the level 1 and 2 is next to two edges between levels 2 and 3 of the same clip.
+These edges between levels 2 and 3 cannot be drawn “under” the mountain so one of them ends up to the left and one to the right of the mountain containing $e$.
+Therefore, no two edges between levels 1 and 2 of the same clip can end up “under” the same mountain.
+Assume that two edges between levels 1 and 2 of different clips end up under the same mountain.
+Then their neighboring edges between the levels 2 and 3 would interleave each other, which contradicts contradicting Lemma 6.
+∎
+Claim 9.
+In any level planar drawing, every $s_{j}$-clip ($j\in[3m]$) occupies mountains of a single mountain $B$-chain and is drawn in such a way that each of its edges between levels 1 and 2 occupies one mountain.
+Proof.
+The unique drawing of the sequence of mountain $B$-chains and the constraints $w_{0}\prec_{4}x$ and $x\prec_{4}w_{m}$ for the level-$4$ vertex of each $s_{j}$-clip imply that the clips must to be drawn in the same space as the mountains.
+The walls split the drawing into a right-of-the-wall part and a left-of-the-wall part, so each clip can be drawn onto at most a single mountain $B$-chain.
+The only space to the draw level-$1$ vertices is “under” the mountains, hence, the edges between levels 1 and 2 of each clip occupy mountains.
+Claim 8 showed that each mountain contains at most one such edge.
+∎
+By design, the number of edges between levels 1 and 2 within the clips is equal to the number of mountains within the mountain $B$-chains.
+In the claims, we proved that each mountain holds exactly one such edge and that each clip falls into a single mountain $B$-chain.
+As the clips represent positive numbers that are strictly between $B/2$ and $B/4$ it is clear that each mountain $B$-chain contains exactly $3$ clips, which directly corresponds to a solution of the corresponding $3$-Partition instance.
+∎
+5 A Linear-Time Algorithm for 2-Level Constrained Level Planarity
+In this section, we present an optimal linear (in the input size) time algorithm for 2-level CLP, which serves as a warm-up for our much more involved approach for 3-level CLP (discussed in Section 6).
+We first discuss how to handle isolated vertices and then how to handle connected input graphs.
+Finally, we observe that connected components of a disconnected input graph can be handled individually, which allows us to state our main algorithm.
+Handling isolated vertices.
+Observe that a 1-level CLP instance $\mathcal{G}=(G,\gamma,(\prec_{i})_{i\in[1]})$ contains no edges, so that all drawings of $\mathcal{G}$ are crossing-free by default.
+Hence, finding a constrained level drawing of $\mathcal{G}$ is equivalent to finding a linear extension of the partial order $\prec_{1}$, which can be done in linear time by performing a topological sorting of the directed graph corresponding to $\prec_{1}$ (recall that a topological sorting of a directed (multi)graph can be performed in time that is linear in its number of vertices and edges; here the number of edges corresponds to the number of constraints).
+Following this idea, any constrained level graph can be transformed into an equivalent simpler instance by removing its isolated vertices:
+Lemma 7 ([28, Lemma 4]).
+Let $\mathcal{G}=(G,\gamma,(\prec_{i})_{i})$ be a constrained level graph, let $G^{\prime}$ be the subgraph of $G$ induced by the non-isolated vertices $V^{\prime}$, and let $\gamma^{\prime}$ and $(\prec^{\prime}_{i})_{i}$ be the restrictions of $\gamma$ and $(\prec_{i})_{i}$ to $V^{\prime}$, respectively.
+There is an algorithm that, given $\mathcal{G}$ and a constrained level planar drawing $\Gamma^{\prime}$ of $\mathcal{G}^{\prime}=(G^{\prime},\gamma^{\prime},(\prec^{\prime}_{i})_{i})$, constructs a constrained level planar drawing of $\mathcal{G}$ in linear [^2] The authors of [28] did not analyze the polynomial factors in the runtimes of their algorithms and just stated that the runtime of [28, Lemma 4]] is polynomial.
+However, it is easy to see that their approach only requires linear time.
+(in the size of the input) time.
+To give a more complete picture, we include the proof idea:
+Proof of Lemma 7.
+For each level $V_{\ell}$, the drawing $\Gamma^{\prime}$ induces a linear order $<_{\ell}^{\prime}$ on $V_{\ell}^{\prime}$ that extends $\prec^{\prime}_{\ell}$.
+Klemz and Sieper [28] showed that the order $<_{\ell}=(<_{\ell}^{\prime}\cup\prec_{\ell})$ is acyclic and that, hence, the desired drawing can be obtained by computing a linear extension of $<_{\ell}$ for each level $V_{\ell}$.
+This can be done by performing a topological sorting of the directed graph corresponding to $\prec_{\ell}$ augmented with the directed edges corresponding to the transitive reduction (i.e., not including transitive edges) of $<_{\ell}^{\prime}$.
+Note that the transitive reduction of $<_{\ell}^{\prime}$ contains only $\mathcal{O}(V_{\ell})$ edges, which are given as part of the encoding of $\Gamma^{\prime}$.
+Thus, the total runtime for computing the topological sortings on all levels is indeed linear in the size of the input ($\mathcal{G}$ and $\Gamma^{\prime}$).
+∎
+Figure 12: A 2-level drawing of a caterpillar. The vertices of its spine are drawn as squares.
+Handling connected graphs.
+A caterpillar is a tree in which all vertices are within distance at most one of a central path (in other words, removing all degree-1 vertices results in a graph that is a path), which is called spine.
+For a caterpillar $G$ with at least three vertices, let $p(G)$ be the subgraph of $G$ induced by $\{v\in V(G)\mid\mathrm{deg}(v)\geq 2\}$, i.e., $p(G)$ is the spine of $G$.
+A drawing of a caterpillar on two levels can be seen in Figure 12.
+Caterpillars are precisely the connected graphs that can be embedded on two levels [21] and we can construct constrained level planar drawings of them efficiently:
+Lemma 8.
+Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{i\in[2]})$ be a constrained level graph where $G$ is a connected.
+There is an algorithm that, given $\mathcal{G}$, either creates a constrained level planar drawing of $\mathcal{G}$ or determines that such a drawing does not exist in linear (in the size of the input) time.
+Proof.
+If $G$ is not a caterpillar (which can be tested in linear time), $\mathcal{G}$ is not (constrained) level planar.
+So assume that $G$ is a caterpillar.
+We may assume that $G$ has at least three vertices, otherwise a constrained level planar drawing is trivial to construct in constant time.
+There are two possible options to draw $p(G)=(v_{1},v_{2},\dots,v_{k})$ on two levels in a level planar fashion: as a zigzag path alternating between the two levels with either $v_{1}v_{2}$ as the left-most edge and $v_{k-1}v_{k}$ as the right most edge or vice versa.
+Suppose we want to test whether a constrained level planar drawing of the former form exists (testing whether a drawing of the latter form exists can be done symmetrically).
+We will do so by performing two topological sortings, one in a directed graph obtained by augmenting the graph $H_{1}$ corresponding to $\prec_{1}$, the other in a directed graph obtained by augmenting the graph $H_{2}$ corresponding to $\prec_{2}$.
+The desired drawing (if it exists) is then obtained by simply placing the vertices of each level in the order given by its topological sorting.
+Without loss of generality, we may assume that $v_{1}\in V_{1}$.
+To ensure that $p(G)$ is drawn crossing-free in the desired zigzag fashion, we add the edges $(v_{1},v_{3}),(v_{3},v_{5}),(v_{5},v_{7}),\dots$ to $H_{1}$ and we add the edges $(v_{2},v_{4}),(v_{4},v_{6}),(v_{6},v_{8}),\dots$ to $H_{2}$.
+To ensure that the remaining edges also have to be drawn in a crossing-free fashion, we observe that for each $i\in\{2,3,\dots,k-1\}$, every neighbor $u$ of the vertex $v_{i}$ has to be drawn between the vertices $v_{i-1}$ and $v_{i+1}$ on the level that does not contain $v_{i}$.
+Similarly, every neighbor $u$ of $v_{1}$ has to be drawn on $L_{2}$ and to the left of $v_{2}$ (if it exists) and every neighbor $u$ of $v_{k}$ has to be drawn on the level that does not contain $v_{k}$ and to the right of $v_{k-1}$ (if it exists).
+It is easy to see that these conditions are not only necessary, but also sufficient to guarantee a crossing free level drawing.
+Moreover, they can be enforced by adding at most two constraints per neighbor $u$ of a spine vertex.
+So overall, it suffices to augment $H_{1}$ and $H_{2}$ with at most $2n$ edges, where $n$ is the number of vertices in $G$, to ensure that the drawing corresponding to the two topological orderings (if they exist) is crossing-free and realizes $p(G)$ in the desired fashion.
+Moreover, by definition of $H_{1}$ and $H_{2}$, the drawing also satisfies all constraints, i.e., it is constrained level planar.
+Conversely, if at least one of our two augmented directed graphs $H_{1}^{\prime},H_{2}^{\prime}$ admits no topological ordering, then there is no constrained level planar drawing of $\mathcal{G}$.
+Given that the sizes of $H_{1}^{\prime}$ and $H_{2}^{\prime}$ are both linear in the size of $\mathcal{G}$, the two topological sortings can be obtained in the desired runtime. ∎
+Main algorithm.
+The fact that each edge of a 2-level planar drawing spans all (two) levels implies that the connected components can be handled individually (Lemma 9), which leads to our linear-time algorithm for 2-level CLP (Theorem 7).
+Lemma 9.
+Let $\mathcal{G}=(G,\gamma)$ be a 2-level graph, and let $A$ and $B$ be two distinct connected components of $G$, both containing at least one edge.
+Then for each level drawing of $\mathcal{G}$ there exists an unbounded y-monotone curve that separates $A$ from $B$ (i.e., $A$ lies to the left of $B$ on every level or $B$ lies to the left of $A$ on every level).
+Proof.
+Let $\Gamma$ be a level planar drawing of $\mathcal{G}$, and let $e=vu\in E(A)$.
+Then $e$ splits the horizontal strip bounded by $L_{1}$ and $L_{2}$ into a left and a right part, and all paths in $G$ from the left to the right part have to intersect $u$ or $v$.
+Given that $A$ and $B$ are disjoint, it follows that $B$ must lie entirely to the left or to the right of $e$.
+A symmetric statement applies to each edge $e^{\prime}\in E(B)$.
+In combination, this yields the claim.
+∎
+Theorem 7.
+There exists an algorithm that, given a constrained 2-level graph $\mathcal{G}=(G,\gamma,$ $(\prec_{i})_{i\in[2]})$, either computes a constrained level planar drawing of $\mathcal{G}$ or determines that such a drawing does not exist in linear (in the size of the input) time.
+Proof.
+In view of Lemma 7, we may assume that $G$ contains no isolated vertices, as they can be handled in a linear time postprocessing step.
+We begin by computing the connected components, denoted by $C^{1},C^{2},\dots,C^{k}$, of $G$, which is easily done in linear time via BFS.
+Then we create a constrained level graph $\mathcal{G}^{j}=(C^{j},\gamma^{j},(\prec_{i}^{j})_{i\in[2]})$ for each connected component $C^{j}$, where $\gamma^{j}$ and $(\prec_{i}^{j})i$ are the restrictions of $\gamma$ and $(\prec_{i})_{i}$ to the vertices of $C^{j}$, respectively.
+This can be done in linear time by sweeping the set of constraints once.
+While doing so, we also create (in linear time) a directed multigraph $H$ with a node for each connected component $C^{j}$ and with a directed edge from a component $C^{v}$ to a component $C^{u}$ for each constraint $v\prec_{i}u$ with $i\in[2],v\in C^{v},u\in C^{u}$.
+In view of Lemma 9, a constrained level planar drawing of $\mathcal{G}$ exists if and only if (a) $H$ is acyclic and (b) each of the graphs $\mathcal{G}^{j}$ admits a constrained level planar drawing.
+Condition (a) can be tested by applying a topological sorting algorithm on $H$ in linear time.
+Condition (b) can be tested by applying Lemma 8 for each $\mathcal{G}^{j}$.
+Given that the graphs $\mathcal{G}^{j}$ partition $\mathcal{G}$, this takes linear time in total.
+Finally, to construct the desired drawing (if it exists), we simply arrange the drawings of the graphs $\mathcal{G}^{j}$ in the order given by the topological sorting of $H$.
+∎
 6 Tractability of 3-Level Constrained Level Planarity
-In this section, we sketch our polynomial-time algorithm for 3-level CLP (see Theorem 1.2).
-For simplicity, we assume that the given level graph $\mathcal{G}$ is constrained level planar and show how to compute a constrained level planar drawing of $\mathcal{G}$.
-We ask the reader to consult the figures to which we refer in order to get an intuitive understanding of the notions that we cannot define formally due to space constraints.
-For details and proofs, see Appendix E.
-Without loss of generality, we make the following assumptions: (i) $\mathcal{G}$ is proper; otherwise we subdivide each edge that crosses level 2.
-(ii) $\mathcal{G}$ has no isolated vertices; they can easily be inserted in a postprocessing step.
-(iii) The component–constraint graph $H$ is strongly connected.
-This directed graph has a node for each connected component (for short, component) of $\mathcal{G}$ and an arc for each pair $(C,C^{\prime})$ of components such that there is a constraint from a vertex in $C$ to a vertex in $C^{\prime}$.
-We often refer to level 1 as bottom level, to level 2 as middle level, and to level 3 as top level.
-Similarly, we call the pair of bottom and middle level lower band and the pair of middle and top level upper band.
-Our algorithm successively adds new constraints to the given constrained level graph $\mathcal{G}$ that do not violate planarity.
-We deduce these new constraints from the structure of $(G,\gamma)$ and from the current set of constraints.
-In the end, this yields a total order of the vertices for each of the three levels and, hence, a constrained level planar drawing of $\mathcal{G}$.
+In this section we show that, given a constrained level graph $\mathcal{G}=(G,\gamma,(\prec_{i})_{i\in[3]})$ of height 3, we can decide in polynomial time whether $\mathcal{G}$ admits a constrained level planar drawing.
+We assume that $\mathcal{G}$ is level planar (which can be tested in linear time [22, 25]) and proper (otherwise we subdivide edges that connect levels 1 and 3).
+Besides referring to levels 1, 2, and 3 by their indices, we also call them bottom level, middle level, and top level, respectively.
+Furthermore, we call the pair of bottom and middle level lower band and the pair of middle and top level upper band.
+Throughout this section, we successively add new constraints to the given constrained level graph $\mathcal{G}$ that we deduce from the structure of $G$, $\gamma$, and the current set of constraints.
+In the end, this yields a total order of the vertices for each of the three levels that corresponds to a constrained level planar drawing of $\mathcal{G}$, or we come to the conclusion that $\mathcal{G}$ does not admit such a drawing.
 In the very beginning and whenever we add new constraints, we exhaustively add the following implicit constraints:
-  • transitivity: $\forall a,b,c\in V(G),\forall i\in[3]\colon a\prec_{i}b\land b\prec_{i}c\Rightarrow a\prec_{i}c$
-  • planarity: $\forall ab,cd\in E(G)\text{ with }i:=\gamma(a)=\gamma(c)\text{ and }j:=\gamma(b)=\gamma(d)\colon\\ a\prec_{i}c\Rightarrow(b\prec_{j}d\lor b=d)$
-The former ensures that the orderings $(\prec_{i})_{i\in[3]}$ remain transitive while the latter can be added without violating realizability, as they need to be respected in every constrained level planar drawing.
-The propagation of these constraints is quite useful as it can dictate the relative positions of vertices that are initially unrelated, see Fig.
-6.
-{subfigure}
-[…]
+  • Transitivity: For every triplet $(a,b,c)$ of vertices on the same level, if there is a constraint from $a$ to $b$ and from $b$ to $c$, then there is also a constraint from $a$ to $c$. Formally,
+    $\forall a,b,c\in V(G)$ with $i:=\gamma(a)=\gamma(b)=\gamma(c)\colon(a\prec_{i}b)\land(b\prec_{i}c)\Rightarrow(a\prec_{i}c)$.
+  • Planarity: For every pair $(ab,cd)$ of edges such that $a$ and $c$, and $b$ and $d$ lie on the same level, if there is a constraint from $a$ to $c$, then there is also a constraint from $b$ to $d$ or $b=d$.
+Formally, $\forall ab,cd\in E(G)\text{ with }i:=\gamma(a)=\gamma(c)\text{ and }j:=\gamma(b)=\gamma(d)\colon$ $(a\prec_{i}c)\Rightarrow(b\prec_{j}d)\lor(b=d)$.
+The transitivity constraints ensure that the orderings $(\prec_{i})_{i\in[3]}$ remain transitive while the planarity constraints can be added without violating realizability, as they need to be respected in every constrained level planar drawing.
+The propagation of these constraints is quite useful as it can dictate the relative positions of vertices that are initially unrelated, see Figure 13.
+Note that we ignore the runtime of adding implicit constraints throughout this section and get back to it in Theorem 8.
+In this section, when we state asymptotic running times such as “linear” or “quadratic”, we refer to the number of vertices of $G$.
+Note that, due to the constraints, the size of the input can already be quadratic (in the number of vertices).
+(a) Initial situation with the given constraint $b_{1}\prec_{1}b_{2}$ but without implicit constraints.
+(b) Constraint $b_{1}\prec_{1}b_{2}$ and edges $b_{1}m_{1}$ and $b_{2}m_{2}$ yield $m_{1}\prec_{2}m_{2}$; edges $m_{1}t_{1}$ and $m_{2}t_{2}$ then yield $t_{1}\prec_{3}t_{2}$.
+(c) Constraint $t_{1}\prec_{3}t_{2}$ and edges $t_{1}m_{1}$ and $t_{2}m_{3}$ yield $m_{1}\prec_{2}m_{3}$. Hence, $m_{3}$ cannot be placed at $p_{1}$.
+Figure 13: Constrained level graph with only one constraint ($b_{1}\prec_{1}b_{2}$).
+Among the three possible positions $p_{1}$, $p_{2}$, and $p_{3}$ to place $m_{3}$ relative to $m_{1}$ and $m_{2}$, position $p_{1}$ is excluded due to the implicit constraints ensuring planarity.
+Some simplifying assumptions.
+For the remainder of this section, we assume that, for each level, the current set of constraints among the vertices on this level is acyclic.
+Otherwise, we reject the instance.
+Due to Lemma 7, we can assume, without loss of generality, that $\mathcal{G}$ contains no isolated vertices.
+We now make some general observations for constrained level graphs that we will later apply to the concrete constrained level graph $\mathcal{G}$ that we are given as input.
+To this end, let $\mathcal{J}=(J,\gamma,(\prec_{i})_{i\in[h]})$ be a constrained level graph of height $h$.
+We define the component–constraint graph $H_{\mathcal{J}}$ of $\mathcal{J}$ as follows.
+The nodes of $H_{\mathcal{J}}$ are the connected components of $J$, and there is an arc from a component $C$ to a component $D$ if there are vertices $u$ in $C$ and $v$ in $D$ with $u\prec_{i}v$ for some level $i$; see Figure 14 (the terms “hook chain”, “hook anchor”, and “hook piece” will be explained later).
+We define, for a component $C$ of $J$ (that is, a node of $H_{\mathcal{J}}$), the graph $\mathcal{J}_{C}$ as the subgraph of $\mathcal{J}$ induced by the vertices of $C$.
+Similarly, we define, for a set $\mathcal{C}$ of components of $J$ (that is, a set of nodes of $H_{\mathcal{J}}$), the subgraph $\mathcal{J}_{\mathcal{C}}$ $=\bigcup_{C\in\mathcal{C}}\mathcal{J}_{C}$ of $\mathcal{J}$.
+Lemma 10.
+A constrained level graph $\mathcal{J}$ admits a constrained level planar drawing if and only if, for every strongly connected component $\mathcal{C}$ of the component–constraint graph $H_{\mathcal{J}}$, the corresponding subgraph $\mathcal{J}_{\mathcal{C}}$ admits a constrained level planar drawing.
+Proof.
+Obviously, a constrained level planar drawing of $\mathcal{J}$ contains, for each strongly connected component $\mathcal{C}$ of $H_{\mathcal{J}}$, a constrained level planar drawing of $\mathcal{J}_{\mathcal{C}}$.
+Now, given, for each strongly connected component $\mathcal{C}$ of $H_{\mathcal{J}}$, a constrained level planar drawing of $\mathcal{J}_{\mathcal{C}}$, we construct a constrained level planar drawing of $\mathcal{J}$.
+Let $H_{\mathcal{J}}^{\prime}$ be the directed graph obtained from $H_{\mathcal{J}}$ by contracting each strongly connected component to a single node.
+Clearly, $H_{\mathcal{J}}^{\prime}$ is acyclic.
+We sort $H_{\mathcal{J}}^{\prime}$ topologically and place the individual drawings of the strongly connected components according to that order from left to right.
+∎
+Hence, from now on we assume, without loss of generality, that the component–constraint graph $H_{\mathcal{J}}$ of any level graph $\mathcal{J}$ is strongly connected.
+(Otherwise, we compute the component–constraint graph $H_{\mathcal{J}}$ of $\mathcal{J}$ in quadratic time and treat each strongly connected component of $H_{\mathcal{J}}$ individually.) In particular, this assumption holds for $\mathcal{G}$.
+Since we assume that $\mathcal{G}$ also does not contain any isolated vertices and is proper, every connected component contains at least one vertex on the middle level.
+For brevity, we call connected components just components and if we speak of components of a constrained level graph $\mathcal{J}=(J,\gamma,(\prec_{i})_{i\in[3]})$, we mean the components of $J$.
+Next, we investigate some properties that are specific to constrained level planar drawings on three levels.
+Provided $\mathcal{G}$ is constrained level planar, these properties will lead to further assumptions on the structure of $\mathcal{G}$.
+On the interaction of components.
+(a) constrained level planar drawing of $\mathcal{J}$: components $C_{1},\dots,C_{7}$ in gray, hook anchors in red, and hook pieces of $C_{3}$ in dark gray
+(b) the corresponding component–constraint graph $H_{\mathcal{J}}$
+Figure 14: Finding the hook chain $\langle C_{1},C_{2},C_{3},C_{7}\rangle$ (corresponding arcs marked in orange).
+Let $\Gamma$ be a constrained level planar drawing of a proper constrained level graph $\mathcal{J}=(J,\gamma,(\prec_{i})_{i\in[3]})$ of height 3.
+Let $\sigma$ be the order of the vertices on the middle level of $\Gamma$ from left to right.
+For a subgraph $X$ of $J$, let the span of $X$ in $\Gamma$, $\operatorname{span}_{\Gamma}(X)$, be the smallest interval of $\sigma$ that contains all middle-level vertices of $X$.
+A component $C$ encloses a component $C^{\prime}$ if $\operatorname{span}_{\Gamma}(C^{\prime})\subsetneq\operatorname{span}_{\Gamma}(C)$; we then call $C^{\prime}$ an enclosed component.
+For example, in Figure 14, $C_{3}$ encloses $C_{4}$, $C_{5}$, and $C_{6}$.
+Recall that a caterpillar is a tree in which all vertices are within distance at most one of a central path, which we call its spine; see Figure 12 for an example.
+Observation 4.
+An enclosed component occupies exactly two levels and is a caterpillar.
+Proof.
+An enclosed component $C$ occupies at least two levels because we assume that the given constrained level graph has no isolated vertices.
+Further, $C$ cannot occupy three levels because otherwise, the component that encloses $C$ could not have middle-level vertices to the left and to the right of the middle level vertices of $C$.
+Caterpillars are precisely the connected graphs that can be embedded into two levels [21].
+∎
+Let $C$ and $C^{\prime}$ be two distinct components of $\mathcal{J}$, and let $\Gamma$ be a constrained level planar drawing of $\mathcal{J}$.
+We say that $C^{\prime}$ hooks into $C$ from the right (in $\Gamma$) if there are vertices $u\neq v$ of $C$ and vertices $u^{\prime}\neq v^{\prime}$ of $C^{\prime}$ such that $u$, $u^{\prime}$, $v$, $v^{\prime}$ occur in this order on the middle level, and $u$ and $v^{\prime}$ are the leftmost and rightmost middle-level vertices of $C\cup C^{\prime}$, respectively.
+For example, in Figure 14, $C_{7}$ hooks into $C_{3}$ from the right, but $C_{3}$ does not hook into $C_{7}$ from the right.
+For a pair of components $(C,C^{\prime})$ such that $C^{\prime}$ hooks into $C$ from the right, we introduce names for some of the vertices of $C$; the same names apply symmetrically to the vertices of $C^{\prime}$.
+We call the vertices in $C\cap\operatorname{span}_{\Gamma}(C^{\prime})$ hook vertices of $C$ towards $C^{\prime}$.
+For example, in Figure 14, $u$ is a hook vertex of $C_{3}$ towards $C_{2}$, and $v$ and $w$ are hook vertices of $C_{3}$ towards $C_{7}$.
+Let the hook piece of $C$ towards $C^{\prime}$ be the subgraph of $J$ induced by the closed neighborhood of the hook vertices of $C$ towards $C^{\prime}$.
+For example, in Figure 14(a), the hook pieces of $C_{3}$ are marked in dark gray.
+Observation 5.
+Let $\mathcal{J}$ be a proper constrained level graph of height 3, and let $\Gamma$ be a constrained level planar drawing of $\mathcal{J}$.
+If $\mathcal{J}$ has two components $C$ and $C^{\prime}$ such that one hooks into the other from the right in $\Gamma$, then the following statements hold.
+  1. 1.
+    The hook piece $p$ of $C$ towards $C^{\prime}$ is a caterpillar on two levels.
+  2. 2.
+    The hook piece $p$ of $C$ towards $C^{\prime}$ contains exactly one vertex with a neighbor outside of $p$.
+  3. 3.
+    All neighbors of the hook vertices of $C$ towards $C^{\prime}$ lie on the same level (see, e.g., the neighbors of $v$ and $w$ in Figure 14(a)).
+  4. 4.
+    If all neighbors of the hook vertices of $C$ towards $C^{\prime}$ lie on the top level, then all neighbors of the hook vertices of $C^{\prime}$ towards $C$ lie on the bottom level, and vice versa (see, e.g., the neighbors of $v$ and $w$ versus the neighbor of $u^{\prime}$ in Figure 14(a)).
+Proof.
+  1. 1.
+    Consider the drawing $\Gamma^{\prime}$ where we have removed $C$ except for $p$.
+By definition, $C^{\prime}$ encloses $p$ in $\Gamma^{\prime}$.
+Due to Observation 4, $p$ is a caterpillar on two levels.
+  2. 2.
+    By definition, the middle-level vertices of $p$ and the middle-level vertices of $C\setminus p$ form two disjoint intervals. Hence, their closed neighborshoods intersect in at most one vertex.
+  3. 3.
+    Since $p$ is a caterpillar on two levels, all neighbors of the hook vertices of $p$ lie on the same level.
+  4. 4.
+    If the neighbors of the hook vertices of $C$ towards $C^{\prime}$ and the neighbors of the hook vertices of $C^{\prime}$ towards $C$ would lie on the same level, their incident edges would cross.
+The vertex in the hook piece of $C$ that has neighbors outside the hook piece is unique due to item 2 of Observation 5; we call it the hook anchor of $C$ towards $C^{\prime}$.
+In Figure 14, the vertices that are hook anchors are marked by red squares.
+Let a hook chain be a maximal sequence $\langle C_{1},\dots,C_{k}\rangle$ of $k\geq 2$ components such that, for every $i\in[k-1]$, $C_{i+1}$ hooks into $C_{i}$ from the right.
+If a component $C$ neither hooks into another component from the right nor another component hooks into $C$ from the right and $C$ is not enclosed by another component, then we call $\langle C\rangle$ a degenerate hook chain; see the red component in Figure 17.
+Lemma 11.
+Let $\mathcal{J}$ be a proper constrained level graph of height 3, and let $\Gamma$ be a constrained level planar drawing of $\mathcal{J}$.
+If $C$, $C^{\prime}$, and $C^{\prime\prime}$ are three distinct components of $\mathcal{J}$, then $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C^{\prime})\cap\operatorname{span}_{\Gamma}(C^{\prime\prime})=\emptyset$.
+Proof.
+Suppose for a contradiction that there are three distinct components $C$, $C^{\prime}$, and $C^{\prime\prime}$ with $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C^{\prime})\cap\operatorname{span}_{\Gamma}(C^{\prime\prime})\neq\emptyset$.
+Without loss of generality, we can assume that $C^{\prime}$ and $C^{\prime\prime}$ have at least two vertices on the middle level each.
+Let $v$ be a vertex of $C$ that lies in $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C^{\prime})\cap\operatorname{span}_{\Gamma}(C^{\prime\prime})$.
+We can also assume that $v$ has a neighbor $\bar{v}$ on the top level.
+Let $u^{\prime}$ be the middle-level vertex of $C^{\prime}$ to the left of $v$ that is closest to $v$.
+Let $w^{\prime}$ be the middle-level vertex of $C^{\prime}$ to the right of $v$ that is closest to $v$.
+Such vertices must exist since $v\in\operatorname{span}_{\Gamma}(C^{\prime})$.
+Let $u^{\prime\prime}$ and $w^{\prime\prime}$ be vertices of $C^{\prime\prime}$ defined in the same way as $u^{\prime}$ and $w^{\prime}$.
+Due to the edge $v\bar{v}$, all paths from $u^{\prime}$ to $w^{\prime}$ must contain a vertex $v^{\prime}$ on the bottom level whose predecessor lies to the left of $v$, while its successor lies to the right of $v$.
+Symmetrically, all paths from $u^{\prime\prime}$ to $w^{\prime\prime}$ must contain a vertex $v^{\prime\prime}$ with the same properties as $v^{\prime}$.
+Then, however, one of the edges incident to $v^{\prime}$ must cross one of the edges incident to $v^{\prime\prime}$.
+This yields the desired contradiction.
+∎
+We next prove some properties of hook chains in constrained level planar drawings on three levels.
+Lemma 12.
+In a constrained level planar drawing $\Gamma$ of a proper constrained level graph $\mathcal{J}$ of height 3, every component is either part of a hook chain or enclosed by a component that is part of a hook chain.
+Proof.
+Observe that Lemma 11 implies that no enclosed component can enclose another component.
+Now, we prove that if there is a component $C$ that is not part of a hook chain, then $C$ is an enclosed component.
+Because $\langle C\rangle$ is not a degenerate hook chain and because of the strong connectivity of $H_{\mathcal{J}}$, there is a component $C^{\prime}$ such that $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C^{\prime})\neq\emptyset$ and $\operatorname{span}_{\Gamma}(C)\nsupseteq\operatorname{span}_{\Gamma}(C^{\prime})$.
+If $\operatorname{span}_{\Gamma}(C)\subseteq\operatorname{span}_{\Gamma}(C^{\prime})$, then $C$ is an enclosed component.
+Otherwise $C$ and $C^{\prime}$ overlap.
+Without loss of generality, there is a vertex $v$ in $C$ on the middle level that lies to the left of $\operatorname{span}_{\Gamma}(C^{\prime})$.
+There is no constraint from a vertex of $C^{\prime}$ to a vertex of $C$ because otherwise $C^{\prime}$ would hook into $C$ from the right.
+However, because of the strong connectivity of $H_{\mathcal{J}}$, there is, for some $\ell>2$, a path $\langle C^{\prime}=C_{1},C_{2},\dots,C_{\ell}=C\rangle$ in $H_{\mathcal{J}}$ such that, for every $i\in[\ell-1]$, it holds that $\operatorname{span}_{\Gamma}(C_{i})\cap\operatorname{span}_{\Gamma}(C_{i+1})\neq\emptyset$.
+Since $\operatorname{span}_{\Gamma}(C^{\prime})\cap\operatorname{span}_{\Gamma}(C_{2})$ is to the right of $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C^{\prime})$, $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C_{\ell-1})$ is to the left of $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C^{\prime})$, and, for every $i\in\{2,\dots,\ell-1\}$, $\operatorname{span}_{\Gamma}(C)\cap\operatorname{span}_{\Gamma}(C^{\prime})\cap\operatorname{span}_{\Gamma}(C_{i})=\emptyset$ (due to Lemma 11), this is not possible.
+Hence, every component is part of a hook chain or is enclosed by a component of a hook chain.
+∎
+Lemma 13.
+A constrained level planar drawing $\Gamma$ of a proper constrained level graph $\mathcal{J}$ of height 3 contains exactly one (possibly degenerate) hook chain.
+Every cycle in the component–constraint graph $H_{\mathcal{J}}$ contains at least one node that corresponds to a component that is part of the hook chain.
+Proof.
+First, we show that there is exactly one hook chain.
+There cannot be two different hook chains $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ because $\operatorname{span}_{\Gamma}(\mathcal{C}_{1})\cap\operatorname{span}_{\Gamma}(\mathcal{C}_{2})=\emptyset$, but then, if say $\mathcal{C}_{1}$ is to the left of $\mathcal{C}_{2}$, there cannot be a directed path from a component in $\mathcal{C}_{2}$ to a component in $\mathcal{C}_{1}$ in $H_{\mathcal{J}}$, which contradicts the strong connectivity of $H_{\mathcal{J}}$.
+Since there needs to be at least one component $C$ whose span is not contained in the span of another component, $C$ is part of a hook chain or $\langle C\rangle$ is a degenerate hook chain.
+Finally, suppose that there is a cycle in $H_{\mathcal{J}}$ containing no node representing a component of the hook chain.
+Then, two consecutive components $C_{i}$ and $C_{i+1}$ of the cycle may not hook.
+Hence, $C_{i+1}$ must be enclosed by $C_{i}$ (or vice versa).
+However, by the same argument $C_{i+1}$ must enclose a component $C_{i+2}$ (if it exists), which is, again, not possible by Lemma 11.
+If the cycle contains only two components, $C_{1}$ may enclose $C_{2}$ (or vice versa), but this would mean that $\langle C_{1}\rangle$ (or $\langle C_{2}\rangle$) is a degenerate hook chain.
+∎
+Lemma 14.
+Let $\Gamma_{1}$ and $\Gamma_{2}$ be two constrained level planar drawings of a proper constrained level graph $\mathcal{J}$ of height 3 whose hook chains are $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively.
+If $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ start in the same component $C$ and end in the same component $C^{\prime}$, then $\mathcal{C}_{1}=\mathcal{C}_{2}$.
+If $\mathcal{J}$ admits a constrained level planar drawing with a hook chain starting in a component $C$ and ending in a component $C^{\prime}$, we can determine this hook chain in linear time given $C$, $C^{\prime}$, and $H_{\mathcal{J}}$.
+Proof.
+In a hook chain, for each pair of consecutive components, the later component has a constraint to the earlier component.
+Therefore, a hook chain from $C$ to $C^{\prime}$ in $\Gamma_{1}$ (resp.
+$\Gamma_{2}$) corresponds to a simple directed path $\pi_{1}$ (resp.
+$\pi_{2}$) from $C^{\prime}$ to $C$ in $H_{\mathcal{J}}$.
+If $\pi_{1}=\pi_{2}$, then clearly $\mathcal{C}_{1}=\mathcal{C}_{2}$.
+Now suppose for a contradiction that $\pi_{1}\neq\pi_{2}$.
+Traverse $\pi_{1}$ and $\pi_{2}$ simultaneously until they differ for the first time.
+Let $\hat{C}$ be the last component that lies on both $\pi_{1}$ and $\pi_{2}$, and let $\hat{C}_{1}$ and $\hat{C}_{2}$ be its distinct successors on $\pi_{1}$ and $\pi_{2}$, respectively.
+In $\Gamma_{1}$, $\hat{C}$ hooks into $\hat{C}_{1}$ from the right, and hence $\operatorname{span}_{\Gamma_{1}}(\hat{C})\cap\operatorname{span}_{\Gamma_{1}}(\hat{C}_{1})\neq\emptyset$.
+According to Lemma 11, the vertices of $\operatorname{span}_{\Gamma_{1}}(\hat{C})\cap\operatorname{span}_{\Gamma_{1}}(\hat{C}_{1})$ (i.e., the middle-level vertices of the hook pieces of $\hat{C}$ and $\hat{C}_{1}$) divide $\Gamma_{1}$ into a left and a right part.
+Hence, every component distinct from $\hat{C}$ and $\hat{C}_{1}$ lies either completely in the left or completely in the right part.
+Since there is a constraint from a vertex of $\hat{C}$ to a vertex of $\hat{C}_{2}$, $\hat{C}_{2}$ lies in the right part.
+Repeating this argument, all components of $\pi_{2}$ succeeding $\hat{C}$ lie in the right part, which, in particular, includes $C$.
+However, due to Lemmas 13 and 12, the first vertex on the middle level in $\Gamma_{1}$ belongs to $C$, which means that $C$ lies in the left part – a contradiction.
+We can find the unique simple path in $H_{\mathcal{J}}$ from $C^{\prime}$ to $C$ in linear time using a breadth-first search starting in $C^{\prime}$. ∎
+Due to Lemma 14, it suffices to check, for a quadratic number of pairs of components $(C,C^{\prime})$, whether $\mathcal{G}$ admits a constrained level planar drawing with a hook chain starting in $C$ and ending in $C^{\prime}$.
+We simply check each pairwise combination of components (including degenerate hook chains where $C=C^{\prime}$).
+Therefore, from now on, we assume that we know the hook chain of $\mathcal{G}$.
+Lemma 15.
+Given a proper constrained level graph $\mathcal{J}=(J,\gamma,(\prec_{i})_{i\in[3]})$ of height 3 and a sequence of components of $\mathcal{J}$ that appears as the hook chain in a level planar drawing of $\mathcal{J}$, we can find in linear time four sets of vertices such that there is a constrained level planar drawing of $\mathcal{J}$ whose set of hook anchors is one of these sets.
+Proof.
+For simplicity, we just use “hook chain” when we speak of the input sequence of components that appears as the hook chain in at least one level planar drawing.
+Note that the statement of the lemma is trivially true for degenerate hook chains.
+Otherwise, we consider every pair $(C,C^{\prime})$ of consecutive components in the hook chain and compute a hook anchor of $C$ towards $C^{\prime}$.
+The hook anchor of $C^{\prime}$ towards $C$ can be computed symmetrically.
+Let $V_{C\leftarrow C^{\prime}}$ be the set of middle-level vertices of $C$ that are involved in constraints from $C^{\prime}$, and let $V_{C^{\prime}\to C}$ be the set of middle-level vertices of $C^{\prime}$ that are involved in constraints to $C$.
+For example, in Figure 15, $V_{C\leftarrow C^{\prime}}=\{v,w^{\prime}\}$.
+Observe that in any constrained level planar drawing of $\mathcal{J}$, the vertices in $V_{C\leftarrow C^{\prime}}$ are hook vertices of $C$ towards $C^{\prime}$, and the vertices in $V_{C^{\prime}\to C}$ are hook vertices of $C^{\prime}$ towards $C$.
+Therefore, the neighbors of $V_{C\leftarrow C^{\prime}}$ all lie on the top level and the neighbors of $V_{C^{\prime}\to C}$ all lie on the bottom level (or vice versa).
+We assume, without loss of generality, that the neighbors of $V_{C\leftarrow C^{\prime}}$ all lie on the top level.
+If $C$ has a predecessor $\overline{C}$ in the hook chain, then there exists a vertex $u\in C$ with a constraint to some vertex in $\overline{C}$.
+Consider the BFS-tree $T$ (of $J$) rooted in $u$.
+Let $a$ be the vertex farthest from $u$ in $T$ such that the subtree rooted in $a$ contains $V_{C\leftarrow C^{\prime}}$.
+We claim that there is a constrained level planar drawing of $\mathcal{J}$ where $a$ is the hook anchor of $C$ towards $C^{\prime}$.
+Note that a hook anchor is a cut vertex of $C$ separating $u$ and $V_{C\leftarrow C^{\prime}}$.
+Moreover, if there is a constrained level planar drawing $\Gamma$ of $\mathcal{J}$ where $C^{\prime}$ hooks into $C$ from the right and the hook anchor of $C$ towards $C^{\prime}$ is not adjacent to a vertex of $V_{C\leftarrow C^{\prime}}$, then there is another constrained level planar drawing $\Gamma^{\prime}$ of $\mathcal{J}$, which is identical to $\Gamma$ except that the hook vertices of $C$ towards $C^{\prime}$ that are to the left of all vertices from $V_{C\leftarrow C^{\prime}}$ are “shifted” to the left of the leftmost middle-level vertex of $C^{\prime}$.
+For example, in Figure 15, $u_{1}$, $u_{2}$, and $u_{3}$ are shifted to the left of $u^{\prime}$.
+Then, the “shifted” vertices are no longer hook vertices and the hook anchor of $C$ towards $C^{\prime}$ is adjacent to a vertex of $V_{C\leftarrow C^{\prime}}$.
+The correctness follows from the fact that the vertices that are not in the subtree of $a$ in $T$ do not have incoming constraints from $C^{\prime}$.
+Note that, due to Lemma 11, there cannot be another component between $C$ and $C^{\prime}$.
+(a) Constrained level planar drawing $\Gamma$ with hook anchor $a_{1}$ in $C$ towards $C^{\prime}$, which is not adjacent to a vertex in $V_{C\leftarrow C^{\prime}}$.
+(b) Constrained level planar drawing $\Gamma^{\prime}$ with hook anchor $a_{2}$ in $C$ towards $C^{\prime}$, which is adjacent to vertex $v\in V_{C\leftarrow C^{\prime}}$.
+Figure 15: By “shifting” hook vertices of $C$ to the left, we obtain, from a constrained level planar drawing $\Gamma$, another constrained level planar drawing $\Gamma^{\prime}$ where the hook anchor of $C$ towards $C^{\prime}$ is adjacent to a vertex with a constraint from a vertex of $C^{\prime}$ (here, $v$).
+Hook pieces are indicated by dark gray background color, hook anchors by red squares, and constraints by blue arrows.
+Now, we consider the case that $C$ is the first component of the hook chain.
+If $C$ has a vertex $u$ on the bottom level, we determine the hook anchor of $C$ towards $C^{\prime}$ using a BFS-tree rooted in $u$ as above.
+Otherwise, $C$ is contained in the upper band and is a caterpillar.
+Since $C$ hooks, $C$ has at least three vertices and a nonempty spine.
+We take each of the at most two endpoints of the spine of $C$ as $u$ and execute the rest of the algorithm once for each of them in order to find the hook anchor of $C$ towards $C^{\prime}$.
+In the symmetric case, when $C^{\prime}$ is the last component and we want to find the hook anchor towards $C$, we do the same as above.
+This results in at most four sets of vertices such that there exists a constrained level planar drawing that has precisely one of these sets as hook anchors.
+Regarding the running time, we execute a (linear-time) BFS at most four times per component of the hook chain.
+Together, the components have size linear in the size of $J$, Hence, the overall running time is linear.
+∎
+Due to Lemma 15, we can find, for $\mathcal{G}$ and a given hook chain $\mathcal{C}$, four sets of vertices one of which must be the set of hook anchors if $\mathcal{G}$ admits a constrained level planar drawing with hook chain $\mathcal{C}$.
+We simply check each of the four possibilities.
+Therefore, from now on, we assume that we know the hook chain and its set of hook anchors.
+Further note that every component contains at least one middle level vertex and that both the number of components and the number of vertices on the middle level are bounded by $O(n)$.
+Hence, within $O(n^{2})$ time, instead of guessing $C$ and $C^{\prime}$ directly, we can guess two middle level vertices $s$ and $t$ and choose the components such that $s\in C$ and $t\in C^{\prime}$.
+We will later use $s$ and $t$ as the first and last vertex on the middle level.
+Drawing a constrained level graph with given hook chain, hook anchors, $s$ and $t$.
+It remains to find a constrained level planar drawing of $\mathcal{G}$, given a (possibly degenerate) hook chain, a set of hook anchors as well as the first and last vertices of the middle level $s$ and $t$.
+In several steps, we modify $\mathcal{G}$ (we call the resulting constrained level graphs $\mathcal{G}=\mathcal{G}_{0},\mathcal{G}_{1},\dots,G_{10}$) such that, in step $i$, $\mathcal{G}_{i}$ is constrained level planar if and only if $\mathcal{G}_{i-1}$ is constrained level planar.
+This culminates in a proper constrained level graph $\mathcal{G}_{10}$ whose constraints yield a total order on each level such that we can easily check whether $\mathcal{G}_{10}$ is constrained level planar.
+In the affirmative case, we know that $\mathcal{G}$ is constrained level planar, too, and we can find a constrained level planar drawing of $\mathcal{G}$ in $\mathcal{O}(n^{5})$ time, where $n$ is the number of vertices of $\mathcal{G}$; see Theorem 8.
+We can assume that every intermediate graph is level planar (when ignoring the constraints) and that its set of constraints is acyclic because these properties can be tested in at most quadratic time and if one does not hold, we can abort.
+We start the construction of $\mathcal{G}_{1}$ with a copy of $\mathcal{G}$.
+Given the hook chain $\langle C_{1},\dots,C_{k}\rangle$ together with a set of hook anchors $\{a_{1},a_{2}^{\prime},a_{2},a_{3}^{\prime},\dots,a_{k-1},a_{k}^{\prime}\}$ determined according to Lemma 15, we add edges to $\mathcal{G}_{1}$ in order to connect all components of the hook chain, as follows.
+For each $i\in[k-1]$, let $w_{i}$ be the vertex on the spine of the hook piece of $C_{i+1}$ towards $C_{i}$ that is farthest from $a^{\prime}_{i+1}$.
+Let $w^{\prime}_{i+1}$ be defined symmetrically within $C_{i}$ and with respect to $a_{i}$.
+For example, in Figure 16, on the spine of the hook piece of $C_{7}$ towards $C_{6}$, $w_{6}$ is farthest from $a^{\prime}_{7}$, and, on the spine of the hook piece of $C_{6}$ towards $C_{7}$, $w^{\prime}_{7}$ is farthest from $a_{6}$.
+Note that $w_{i}$ and $w^{\prime}_{i+1}$ can lie on any level of the band where its hook pieces lies.
+We now add the edges $a_{i}w_{i}$ and $a^{\prime}_{i+1}w^{\prime}_{i+1}$.
+We subdivide every edge that has one endpoint on the upper and one endpoint on the lower level with a new vertex on the middle level in order to make $\mathcal{G}_{1}$ proper.
+See Figure 16 for an example with $i=6$.
+We call the resulting component of $\mathcal{G}_{1}$ the main component.
+Figure 16: Constructing the main component: original components in gray, hooking constraints in blue, hook anchors (according to Lemma 15) in red, and edges in $E(\mathcal{G}_{1})\setminus E(\mathcal{G})$ in green.
+Lemma 16.
+The constrained level graph $\mathcal{G}_{1}$ is constrained level planar if and only if $\mathcal{G}$ is.
+Furthermore, given $\mathcal{G}$, a hook chain, and a set of hook anchors that appear in some constrained level planar drawing of $\mathcal{G}$, we can construct $\mathcal{G}_{1}$ in linear time.
+Proof.
+If $\mathcal{G}$ is not constrained level planar, then certainly $\mathcal{G}_{1}$ is not constrained level planer either (because it has the same set of constraints, but additional edges).
+On the other hand, if $\mathcal{G}$ is constrained level planar, then $\mathcal{G}$ admits a constrained level planar drawing $\Gamma$ with the hook chain $\langle C_{1},\dots,C_{k}\rangle$.
+If the hook chain is degenerate ($k=1$) it holds that $\mathcal{G}=\mathcal{G}_{1}$ and we are done, so we assume $k\geq 2$.
+We now have the hook anchors $\{a_{1},a^{\prime}_{2},a_{2},a^{\prime}_{3},\dots,a_{k-1},a^{\prime}_{k}\}$.
+Consider, for any $i\in[k-1]$, the vertices $a_{i}$ and $w_{i}$ (the argument is symmetric for $a^{\prime}_{i}$ and $w^{\prime}_{i}$ with $i\in\{2,\dots,k\}$).
+The hook piece of $C_{i+1}$ towards $C_{i}$ is a caterpillar whose spine is oriented away (i.e., to the left) of its hook anchor $a^{\prime}_{i+1}$ in $\Gamma$.
+Since $a_{i}$ is a hook anchor, the leftmost vertex of $C_{i+1}$ on the middle level is visible from $a_{i}$.
+This leftmost vertex is either the last vertex $w_{i}$ of the spine of the hook piece of $C_{i+1}$ towards $C_{i}$, or adjacent to it.
+We need to show that we can include a $y$-monotone curve between $a_{i}$ and $w_{i}$ in $\Gamma$.
+This is obvious if $w_{i}$ is on the middle level.
+If $w_{i}$ is not on the middle level, then we can draw the curve connecting $a_{i}$ with a point in the left vicinity of the leftmost vertex of $C_{i+1}$ on the middle level, and follow its edge towards $w_{i}$.
+Thus, $w_{i}$ can be reached by $a_{i}$ in $\Gamma$ and, by adding the (possibly subdivided) edge $a_{i}w_{i}$ to $\Gamma$, $\Gamma$ can be extended to a constrained level planar drawing of $\mathcal{G}_{1}$.
+We can find each last vertex of the spine in linear time and the size of the spines together is at most linear in the size of $\mathcal{G}$.
+Inserting the new edges can again be done in linear time.
+Hence, we can construct $\mathcal{G}_{1}$ in linear time.
+∎
+Figure 17: Drawing of the main component of a level graph with special vertices $s=m_{1}$ and $t=m_{30}$: backbone edges are thick and red, enclosed components gray, pieces orange (edges connecting fingers to their anchors dashed), and edges incident to leaves purple.
+Next we fix the orientation of the main component.
+To this end, we use the “guessed” vertices $s$ and $t$ such that (i) the vertices along every simple $s$$t$ path are ordered from left to right and (ii) the vertices that do not lie on any simple $s$$t$ path induce components with a “simple” structure.
+Accordingly, let the backbone spanned by $s,t$ (or, for short, the backbone if $s$ and $t$ are clear from the context) of the main component be the subgraph of $\mathcal{G}_{1}$ that corresponds to the union of all simple $s$$t$ paths.
+In Figure 17, the edges of the backbone spanned by $s=m_{1}$ and $t=m_{30}$ are drawn thick and red.
+Let a separator vertex be a middle-level vertex on the backbone that has a backbone neighbor on both the top and the bottom level.
+Further let an anchor be a vertex on the backbone that is incident to vertices that are not on the backbone.
+Consider the connected components of $\mathcal{G}_{1}$ after removing the backbone.
+Each such component is incident to a unique anchor in $\mathcal{G}_{1}$, otherwise there would be a simple $s$$t$ path through two of its anchors.
+In this case the path would be part of the backbone.
+Let a piece be such a component including its anchor and all edges connecting the component with the anchor.
+If there is a piece that occupies all three levels ignoring its anchor, we reject the current guess of $s$ and $t$.
+Lemma 17.
+If $\mathcal{G}_{1}$ admits a constrained level planar drawing, then there is a choice of $s$ and $t$ such that each piece ignoring its anchor lies on at most two levels, and that $s$ and $t$ are the leftmost and rightmost vertex of $\mathcal{G}_{1}$ on the middle level, respectively.
+(Hence, ignoring its anchor, every piece is a caterpillar.)
+Proof.
+Let $\Gamma$ be a constrained level planar drawing of $\mathcal{G}_{1}$, let $s$ be the leftmost and $t$ be the rightmost middle-level vertex.
+Since the hook chain is degenerate, $s$ and $t$ both lie in the main component and hence there is a simple path $\pi$ from $s$ to $t$.
+The path $\pi$, which is part of the backbone spanned by $s$ and $t$, already partitions the main component into connected components of height at most 2.
+∎
+Following Lemma 17, we define $\mathcal{G}_{2}$ to be the constrained level graph based on $\mathcal{G}_{1}$ with the additional constraints $s\prec_{2}t$ as well as $s\prec_{2}v$ and $v\prec_{2}t$ for every vertex $v$ on the middle level.
+(Recall that further implicit constraints are added automatically.)
+Lemma 18.
+If $\mathcal{G}_{2}$ admits a constrained level planar drawing, then, for every simple $s$$t$ path $\pi$ and for every pair $(u,v)$ of vertices of $\pi$ such that $u$ precedes $v$ in $\pi$ and $\gamma(u)=\gamma(v)=i$, $\mathcal{G}_{2}$ contains the constraint $u\prec_{i}v$.
+Proof.
+Fix any simple $s$$t$ path $\pi$.
+First, consider the case that $\pi$ lies on only one band, say, the lower band.
+Then, $\pi$ consists of an odd number $l$ of vertices.
+Pick $k$ such that $l=2k-1$.
+We name the vertices along $\pi$ $s=m_{1},b_{1},m_{2},b_{2},m_{3},\dots,b_{k-1},m_{k}=t$, where $m_{1},m_{2},\dots,m_{k}$ are the middle-level vertices of $\pi$ in order and $b_{1},b_{2},\dots,b_{k-1}$ are the bottom-level vertices of $\pi$ in order.
+By definition, $\mathcal{G}_{2}$ contains the constraints $m_{1}\prec_{2}m_{2}$, $m_{1}\prec_{2}m_{3}$, $\dots$, $m_{1}\prec_{2}m_{k}$.
+These constraints imply (due to the implicit planarity constraints) together with the edge $m_{1}b_{1}$ on the one hand and the edges $m_{2}b_{2},m_{3}b_{3},\dots,m_{k-1}b_{k-1}$ on the other hand that $\mathcal{G}_{2}$ contains also the constraints $b_{1}\prec_{1}b_{2}$, $b_{1}\prec_{1}b_{3}$, $\dots$, $b_{1}\prec_{1}b_{k-1}$.
+In turn, these constraints imply together with the edge $b_{1}m_{2}$ on the one hand and the edges $b_{2}m_{3},b_{3}m_{4},\dots,b_{k-1}m_{k}$ on the other hand that $\mathcal{G}_{2}$ contains also the constraints $m_{2}\prec_{2}m_{3}$, $m_{2}\prec_{2}m_{4}$, $\dots$,… , $m_{2}\prec_{2}m_{k}$.
+Repeating this argument shows that, for every pair $(u,v)$ of vertices on $\pi$ such that $u$ precedes $v$ in $\pi$ and $\gamma(u)=\gamma(v)=i$, $\mathcal{G}_{2}$ contains the constraint $u\prec_{i}v$.
+Second, consider the case that $\pi$ lies on all three levels.
+Let $r$ be the first (middle-level) vertex of $\pi$ whose predecessor and successor on $\pi$ lie on distinct levels.
+Since we have added the constraint $s\prec_{2}r$, we repeat the first argument of this proof to show the statement for the $s$$r$ sub-path of $\pi$ (where $r$ takes the role of $t$).
+If $r$ is the only (middle-level) vertex of $\pi$ whose predecessor and successor on $\pi$ lie on distinct levels, then we apply the first argument of this proof to show the statement for the $r$$t$ sub-path of $\pi$.
+Otherwise, we recursively apply the second argument of this proof to show the statement for the $r$$t$ sub-path of $\pi$ (where $r$ takes the role of $s$).
+∎
+Note that, in $\mathcal{G}_{2}$, the order $\prec_{2}$ restricted to the backbone vertices is not necessarily total; for example, in Figure 17, the order of $m_{8}$ and $m_{10}$ as well as the order of $m_{17}$ and $m_{18}$ is not fixed.
+For each such set of permutable middle-level vertices on the backbone individually, pick an arbitrary topological order (with respect to the constraints) and fix it by adding new constraints.
+We call the resulting constrained level graph $\mathcal{G}_{3}$.
+Lemma 19.
+The constrained level graph $\mathcal{G}_{3}$ is constrained level planar if and only if $\mathcal{G}_{2}$ is.
+Furthermore, $\mathcal{G}_{3}$ can be constructed from $\mathcal{G}_{2}$ in quadratic time.
+Proof.
+If there is a constrained level planar drawing of $\mathcal{G}_{3}$, we also have a constrained level planar drawing of $\mathcal{G}_{2}$ because $\mathcal{G}_{3}$ has the same underlying level graph as $\mathcal{G}_{2}$ and a superset of the constraints.
+Suppose for a contradiction that $\mathcal{G}_{2}$ admits a constrained level planar drawing $\Gamma$, but $\mathcal{G}_{3}$ does not.
+Consider each set of vertices whose order has been fixed by adding constraints in $\mathcal{G}_{3}$.
+Since $\mathcal{G}_{3}$ is not constrained level planar, the involved vertices cannot be permuted in $\Gamma$ to match the order prescribed in $\mathcal{G}_{3}$.
+This can only happen if there is a constraint in $\mathcal{G}_{2}$ contradicting the order prescribed in $\mathcal{G}_{3}$.
+Then, however, this does not correspond to a topological order.
+To detect each such set of vertices, construct the auxiliary graph that has a node for every middle-level vertex on the backbone and an edge for each pair of vertices that are not ordered by a constraint.
+This graph can be constructed in quadratic time and each connected component corresponds to such a set of vertices.
+Finding topological orders and inserting the new constraints can be done in overall quadratic time.
+∎
+From Lemmas 19 and 18 the following observation follows.
+Observation 6.
+If $\mathcal{G}_{3}$ admits a constrained level planar drawing, then the orders on all levels restricted to the backbone vertices are total.
+Now let a finger be a piece (including its anchor) that lies on all three levels and hence has an anchor on the top or bottom level.
+Further let a hand be the union of fingers incident to the same anchor.
+For example, in Figure 17, the two yellow components incident to $b_{1}$ are fingers that together form a hand.
+Let $\alpha$ be the anchor of a hand, let $N_{\alpha}$ be the set of neighbors of $\alpha$ within the hand, and let $L_{\alpha}$ ($R_{\alpha}$) be the set of backbone neighbors of $\alpha$ that lie before (after) $\alpha$ on some simple $s$$t$ path.
+For example, consider the hand with anchor $\alpha=b_{1}$ in Figure 17.
+We have $L_{\alpha}=\{m_{1},m_{2}\}$, $N_{\alpha}=\{m_{3},m_{4},m_{6}\}$, and $R_{\alpha}=\{m_{7}\}$.
+Note that $L_{\alpha}$ and $R_{\alpha}$ are nonempty because $\alpha$ cannot be $s$ or $t$.
+We construct the constrained level graph $\mathcal{G}_{4}$ from $\mathcal{G}_{3}$ by iterating over each hand of $\mathcal{G}_{3}$ and performing the following operations: (i) Let $\alpha$ be the anchor of the hand, (ii) add constraints from all vertices in $L_{\alpha}$ to all vertices in $N_{\alpha}$, (iii) add constraints from all vertices in $N_{\alpha}$ to all vertices in $R_{\alpha}$, and (iv) remove all edges between $\alpha$ and $N_{\alpha}$.
+Note that fingers in $\mathcal{G}_{3}$ are represented by enclosed components in $\mathcal{G}_{4}$.
+Due to the implicit constraints, two such enclosed components that originate from different hands are ordered according to their anchors.
+Further note that the component–constraint graph $H_{\mathcal{G}_{4}}$ of $\mathcal{G}_{4}$ remains strongly connected since every finger in $\mathcal{G}_{3}$ that becomes a component in $\mathcal{G}_{4}$ has both incoming and outgoing constraints with respect to the main component.
+Lemma 20.
+The constrained level graph $\mathcal{G}_{4}$ is constrained level planar if and only if $\mathcal{G}_{3}$ is.
+Furthermore, $\mathcal{G}_{4}$ can be constructed from $\mathcal{G}_{3}$ in quadratic time.
+Proof.
+If there is a constrained level planar drawing $\Gamma_{3}$ of $\mathcal{G}_{3}$, we can obtain a constrained level planar drawing of $\mathcal{G}_{4}$ by removing, for each hand with anchor $\alpha$, the edges between $\alpha$ and $N_{\alpha}$.
+Note that the additional constraints in $\mathcal{G}_{4}$ are satisfied because otherwise edges of the fingers in $\mathcal{G}_{3}$ would cross backbone edges in $\Gamma_{3}$.
+If there is a constrained level planar drawing $\Gamma_{4}$ of $\mathcal{G}_{4}$, we can obtain a constrained level planar drawing of $\mathcal{G}_{3}$ by adding, for each hand with anchor $\alpha$, the edges between $\alpha$ and $N_{\alpha}$.
+These edges do not cross any other edges because the additional constraints imply that the vertices of $N_{\alpha}$ lie between the vertices of $L_{\alpha}$ and $R_{\alpha}$.
+Since the edges between $L_{\alpha}\cup R_{\alpha}$ and $\alpha$ have no crossings, the same holds for the new edges.
+Since we know, for each vertex, whether it lies on the backbone, we can determine all anchors and all fingers in linear time. Trivially, the number of constraints that we add is at most quadratic. ∎
+Now each remaining piece lies within one of the two bands (including its unique anchor); hence it is a caterpillar.
+In the simplest case, this is just a leaf.
+Next, we will get rid of such leaves.
+Let $\mathcal{G}_{5}$ be the constrained level graph obtained from $\mathcal{G}_{4}$ by removing all leaves attached to backbone vertices except for possibly $s$ and $t$.
+Recall that if $\mathcal{G}_{4}$ is not level planar (when ignoring the constraints) or its constraints contain a cycle, we abort here before constructing $\mathcal{G}_{5}$.
+Lemma 21.
+The constrained level graph $\mathcal{G}_{5}$ is constrained level planar if and only if $\mathcal{G}_{4}$ is. Furthermore, $\mathcal{G}_{5}$ can be constructed from $\mathcal{G}_{4}$ in linear time.
+Proof.
+If $\mathcal{G}_{4}$ is constrained level planar, then certainly $\mathcal{G}_{5}$ is also constrained level planar (because we have only removed vertices and edges).
+Suppose for a contradiction that $\mathcal{G}_{5}$ is constrained level planar, but $\mathcal{G}_{4}$ is not.
+We remove the leaves of $\mathcal{G}_{4}$ iteratively.
+Let $l$ be the leaf whose removal makes the resulting constrained level graph $\mathcal{G}_{4}^{\prime}$ constrained level planar.
+Let $V^{\rightarrow}$ be the set of vertices having a constraint to $l$, and let $V^{\leftarrow}$ be the set of vertices having a constraint from $l$.
+Furthermore, let $a$ be the unique neighbor of $l$.
+Observe that in every constrained level planar drawing of $\mathcal{G}_{4}^{\prime}$, all vertices from $V^{\rightarrow}$ precede all vertices from $V^{\leftarrow}$ because of the transitivity constraints due to $l$ in $\mathcal{G}_{4}$.
+Fix any constrained level planar drawing $\Gamma$ of $\mathcal{G}_{4}^{\prime}$.
+In $\Gamma$, place $l$ on level $\gamma(l)$ immediately to the right of the rightmost vertex of $V^{\rightarrow}$ or as the first vertex on level $\gamma(l)$ if $V^{\rightarrow}=\emptyset$.
+We call the resulting drawing $\Gamma^{\prime}$.
+Note that $\Gamma^{\prime}$ fulfills all constraints and the only crossings of $\Gamma^{\prime}$ involve the edge $al$.
+Let $X$ be the set of edges that cross the edge $al$.
+Note that $X$ is not empty, otherwise $\mathcal{G}_{4}^{\prime}+al$ would be constraint level planar.
+Suppose that there is an edge $cd\in X$ such that $\gamma(a)=\gamma(c)=:i$, $\gamma(l)=\gamma(d)=:j$, and there is a constraint between $a$ and $c$ or between $l$ and $d$ in $\mathcal{G}_{4}$.
+If $\mathcal{G}_{4}$ contains the constraint $l\prec_{j}d$, then, due to the planarity constraints, $\mathcal{G}_{4}$ contains also the constraint $a\prec_{i}c$, and vice versa.
+As $d\in V^{\leftarrow}$, $c$ needs to be to the left of $a$ in $\Gamma^{\prime}$ to cross $al$; see Figure 18(a).
+This, however, contradicts the constraint $a\prec_{i}c$.
+A symmetric argument holds if there are the constraints $d\prec_{j}l$ and $c\prec_{i}a$.
+Hence, no endpoint of an edge in $X$ is involved in a constraint with $a$ or $l$.
+(a) If a constraint between $a$ and $c$ or between $l$ and $d$ exists, the other exists as well and is violated.
+(b) Edges do not cross from the left (when going from $l$ to $a$).
+(c) The edge $al$ crosses two enclosed components that have no constraints with $a$ or $l$.
+(d) The vertices of the enclosed components to the right of $a$ and $l$ can be moved to the free space directly to the left of $a$ and $l$.
+Figure 18: Configurations appearing in the proof of Lemma 21, where we obtain $\mathcal{G}_{5}$ from $\mathcal{G}_{4}$ by removing leaves attached to the backbone.
+We consider the leaf $l$ and its unique anchor $a$ and assume that the edge $cd$ crosses $al$.
+Next, suppose that there is an edge $cd\in X$ such that $a$ precedes $c$ on level $i=\gamma(a)=\gamma(c)$ and $d$ precedes $l$ on level $j=\gamma(l)=\gamma(d)$ in $\Gamma^{\prime}$; see Figure 18(b).
+Consider the rightmost vertex $b$ of $V^{\rightarrow}$.
+Since all crossings in $\Gamma^{\prime}$ involve the edge $al$, vertex $b$ is not incident to any edge crossing $cd$.
+Moreover, $b$ is not incident to any edge crossing $al$ because $b$ is involved in a constraint with $l$.
+This, however, would force $b$ to be an isolated vertex, which $\mathcal{G}_{4}$ does not contain.
+Hence, all edges of $X$ cross $al$ in the other direction (that is, for every edge $cd$ in $X$, $c$ precedes $a$ and $l$ precedes $d$).
+This means that $a$ has no neighbor on level $j$ that is to the left of $l$ and, hence, we can place vertices with neighbors on level $j$ directly to the left of $a$, which we will exploit in the next step.
+Similarly, $l$ has no neighbor of level $i$ that is to the left or right of $a$, simply because $l$ is a leaf.
+Now, let $cd$ be the leftmost edge of $X$, again with $\gamma(a)=\gamma(c)=:i$ and $\gamma(l)=\gamma(d)=:j$.
+Note that $c$ cannot be a vertex of the backbone because then it would have a constraint with $a$ due to Lemma 19.
+Hence, $c$ (and $cd$) is part on an enclosed component or of a piece (potentially being another leaf).
+This enclosed component or piece lies on the band formed by levels $i$ and $j$.
+In either case, no vertex of the enclosed component or piece is involved in a constraint with $a$ or $l$.
+If $c$ lies in an enclosed component $C$, we can move all vertices of $C$ on level $i$ that are to the right of $a$ directly to the left of $a$, and we can move all vertices of $C$ on level $j$ that are to the right of $l$ directly to the left of $l$.
+For an illustration of this step, see Figures 18(c) and 18(d).
+If $c$ lies in a piece $p$, we distinguish two subcases depending on where $p$ is attached to the backbone with respect to $al$.
+If $p$ is attached to the backbone on the left side of $al$ (or the anchor of $p$ is $a$), we can move all vertices of $p$ on level $i$ that are to the right of $a$ directly to the left of $a$, and we can move all vertices of $p$ on level $j$ that are to the right of $l$ directly to the left of $l$.
+Otherwise, if $p$ is attached to the backbone on the right side of $al$, we can move all vertices of $p$ on level $i$ that are to the left of $a$ directly to the right of $a$ (in this case, $a$ cannot have a neighbor on level $j$ that is to the right of $l$), and we can move all vertices of $p$ on level $j$ that are to the left of $l$ directly to the right of $l$.
+If not all endpoints of $X$ lie in the enclosed component or piece that we have treated above, we let $cd$ be the leftmost untreated edge in $X$ and repeat the above treatment.
+In the end, after treating all edges in $X$, we obtain a constrained level planar drawing of $\mathcal{G}_{4}^{\prime}+al$; a contradiction.
+Summing up, we conclude that, if $\mathcal{G}_{4}$ is not constrained level planar, then neither $\mathcal{G}_{5}$ is constrained level planar.
+Regarding the running time, we can check, for each neighbor of a backbone vertex, in constant time whether it is a leaf and, if so, remove it. ∎
+Now, due to Lemma 21, each remaining piece (including its unique anchor) is a caterpillar on one band that is not a single edge. This leads to the following simple observation.
+Observation 7.
+In $\mathcal{G}_{5}$, there is no set of three pieces that share an anchor.
+Note that the backbone subdivides the bands into regions above and below the backbone.
+In the next step, we construct a constrained level graph $\mathcal{G}_{6}$ by adding further constraints to $\mathcal{G}_{5}$ that assign pieces to these regions.
+To this end, we make the following definitions.
+Recall that a separator vertex is a middle-level vertex on the backbone that has a backbone neighbor on both the top and the bottom level.
+Let a gap be a pair $(a,c)$ of separator vertices such that $a\prec_{2}c$ and there is no separator vertex $b$ with $a\prec_{2}b\prec_{2}c$.
+If, for every simple $s$$t$ path $\pi$ through $a$ and $c$, the subpath from $a$ to $c$ goes through the lower band, then we call $(a,c)$ an upper gap.
+Symmetrically, if it goes through the upper band, then we call $(a,c)$ a lower gap.
+Otherwise, that is, if there is a path through $a$ and $c$ whose subpath goes through the lower band and there is a path through $a$ and $c$ whose subpath goes through the upper band, then we call $(a,c)$ a closed gap (for instance see the gap between $m_{16}$ and $m_{19}$ in Figure 17).
+Closed gaps can enclose only leaves and isolated vertices.
+Hence, they are empty in $\mathcal{G}_{5}$.
+Consider the sequence of upper and lower gaps.
+(If this sequence is empty, simply omit the following construction and continue with $\mathcal{G}_{6}:=\mathcal{G}_{5}$.) In this sequence, we call a maximal sequence of consecutive upper (lower) gaps an upper (lower) gap group.
+We sort these gap groups from left to right according to the order of their separator vertices and name them $W_{1},\dots,W_{l}$ for some $l\geq 1$ (see Figure 19(a) for an example).
+Furthermore, we have two unbounded gap groups $W_{0}$ and $W_{l+1}$ representing the space to the left of the first separator vertex and to the right of the last separator vertex, respectively.
+Now we construct a 2-SAT formula whose solutions represent assignments of the pieces to gap groups.
+Note that some pieces can only be placed into one gap group; hence we ignore them in this step.
+These are all pieces whose anchor lies on the middle level (e.g., the pieces with anchors $m_{14}$ and $m_{28}$ in Figure 17) and whose anchor is incident to a closed gap (e.g., the piece with anchor $t_{8}$ in Figure 17).
+Two pieces with the same anchor need to be assigned to different gap groups.
+(For example, in Figure 19(a), $p_{5}$ and $p_{6}$ must lie in different gap groups.) Three or more pieces with the same anchor do not exist due to Observation 7.
+The assignment of pieces to gap groups depends also on the placement of the enclosed components of $\mathcal{G}_{5}$ (including previous fingers); see Figure 19(a).
+To model these dependencies, we define a directed auxiliary graph $H_{\mathrm{gap}}$ as follows.
+The nodes of $H_{\mathrm{gap}}$ are the pieces (that we do not ignore), the enclosed components (which both span only one band), and the gap groups.
+We now define the arc set of $H_{\mathrm{gap}}$ in order to express that certain pieces and components must be placed to the left (or right) of certain gap groups or other pieces and components.
+For an enclosed component $C$, $H_{\mathrm{gap}}$ has an arc from the gap group node of $W_{i}$ to $C$ if $W_{i+1}$ is the first gap group into which $C$ can be placed due to its band and its constraints from backbone vertices.
+Symmetrically, $H_{\mathrm{gap}}$ has an arc from $C$ to the gap group vertex of $W_{j}$ if $W_{j-1}$ is the last gap group into which $C$ can be placed due to its band and its constraints to backbone vertices.
+For an enclosed component or a piece $x$ and an enclosed component or a piece $y$ on distinct bands, $H_{\mathrm{gap}}$ has an arc from $x$ to $y$ if there is a constraint from a vertex of $x$ to a vertex of $y$ (see Figure 19(b) for an example).
+Next we describe our 2-SAT formula $\Phi$.
+For each piece $p$, $\Phi$ has a variable $x_{p}$ whose truth value (true/false) represents the position of $p$ (left/right) with respect to its anchor.
+For two pieces $p$ and $q$ with the same anchor, we add the condition $x_{p}\neq x_{q}$, which can be expressed by the clauses $(x_{p}\lor x_{q})$ and ($(\lnot x_{p}\lor\lnot x_{q})$ (see $p_{5}$ and $p_{6}$ in Figure 19(a)).
+We now use $H_{\mathrm{gap}}$ to construct further clauses of $\Phi$.
+Let $g_{\mathrm{left}}(p)$ and $g_{\mathrm{right}}(p)$ be the indices of the left and right gap group, respectively, where a piece $p$ can be placed.
+Note that these are the indices of the gap groups to the left and to the right of the anchor of $p$, respectively.
+(a) Clauses of $\Phi$: $p_{1}$ is fixed to the left, $p_{5}$ and $p_{6}$ must be different, $p_{2}$ cannot be right while $p_{3}$ is left, $p_{4}$ and $p_{6}$ force each other to the left and right respectively.
+(Main component in red, pieces in orange, enclosed components in black and constraints in blue.)
+(b) The graph $H_{\mathrm{gap}}$ with two types of constraints. The blue constraints are directly induced by constraints in $G$ while the violet constraints are induced by planarity.
+(c) A solution of the 2-SAT instance allows for a consistent labeling of the nodes in $H_{\mathrm{gap}}$.
+The depicted situation of two inconsistently labeled paths $p\leadsto q$ and $p^{\prime}\leadsto q^{\prime}$ cannot occur.
+Figure 19: A constrained level planar graph $\mathcal{G}$ and the corresponding graph $H_{\mathrm{gap}}$.
+For every pair of a gap group $W_{i}$ ($i\in\{0,\dots,l+1\}$) and a piece $p$, we determine the length $L$ of a longest directed path from $W_{i}$ to $p$ in $H_{\mathrm{gap}}$ if such a path exists.
+This can be done in linear time since $H_{\mathrm{gap}}$ is acyclic (if $\mathcal{G}$ is constrained level planar).
+If a path exists, each node on the path needs to be assigned to a different gap group (in increasing order) since the arcs of $H_{\mathrm{gap}}$ connect objects (i.e., pieces, enclosed components, or gap groups) on distinct bands.
+Hence, if $L>g_{\mathrm{right}}(p)-i$, then $\mathcal{G}_{5}$ is not constrained level planar.
+If $L=g_{\mathrm{right}}(p)-i$, then $p$ needs to lie on the right side of its anchor.
+Thus, we set $x_{p}$ to false.
+Otherwise, if $L<g_{\mathrm{right}}(p)-i$, then the path does not restrict the placement of $p$.
+Symmetrically, if there is a path from $p$ to $W_{i}$, we let $L$ be the length of a longest such path.
+If $L>i-g_{\mathrm{left}}(p)$, then $\mathcal{G}_{5}$ is not constrained level planar.
+Else, if $L=i-g_{\mathrm{left}}(p)$, then $p$ needs to lie on the left side of its anchor.
+Thus, we set $x_{p}$ to true (see $p_{1}$ in Figure 19).
+Otherwise, if $L<i-g_{\mathrm{left}}(p)$, then the path does not restrict the placement of $p$.
+For every pair $(p,q)$ of pieces, we determine the length $L$ of a longest directed path from $p$ to $q$ if it exists.
+If $L>g_{\mathrm{right}}(q)-g_{\mathrm{left}}(p)$, then $\mathcal{G}_{5}$ is not constrained level planar.
+If $L=g_{\mathrm{right}}(q)-g_{\mathrm{left}}(p)$, then $p$ needs to be placed on the left side of its anchor while $q$ needs to be placed on the right side of it anchor (see $p_{4}$ and $p_{6}$ in Figure 19).
+Hence, we set $x_{p}$ to true and $x_{q}$ to false.
+If $L=g_{\mathrm{right}}(q)-g_{\mathrm{left}}(p)-1$, then $p$ cannot be placed on the right side of its anchor if $q$ is placed on the left side of its anchor (see $p_{2}$ and $p_{3}$ in Figure 19).
+Hence, we add the clause $(p\lor\lnot q)$ to $\Phi$.
+Otherwise, the placements of $p$ and $q$ do not depend on each other.
+If $\Phi$ admits a feasible truth assignment, we assign the pieces to gap groups accordingly.
+We do this by adding constraints between the corresponding separator vertices and the middle-level vertices of the pieces.
+The resulting constrained level graph is $\mathcal{G}_{6}$.
+Lemma 22.
+The constrained level graph $\mathcal{G}_{6}$ is constrained level planar if and only if $\mathcal{G}_{5}$ is.
+Furthermore, $\mathcal{G}_{6}$ can be constructed from $\mathcal{G}_{5}$ in quadratic time.
+Proof.
+If $\mathcal{G}_{5}$ is not constrained level planar, then certainly $\mathcal{G}_{6}$ is not constrained level planar either (because it is the same constrained level graph but with additional constraints).
+Otherwise assume that $\mathcal{G}_{5}$ is constrained level planar but $\mathcal{G}_{6}$ is not.
+This means that we have assigned the pieces to gap groups such that, for this assignment, $\mathcal{G}_{5}$ does not admit a constrained level planar drawing.
+Recall that our assignment takes all constraints to/from the backbone into account and that arbitrarily many enclosed components in one band can be assigned to the same gap (group).
+The only remaining case is that chains of enclosed components or pieces on alternating bands (which are both nodes of $H_{\mathrm{gap}}$) cannot be accommodated.
+It suffices to consider objects on alternating bands (as they appear in $H_{\mathrm{gap}}$) because two consecutive objects on the same band fit into the same gap group, and the predecessors of the first object are also predecessors of the second object due to transitivity constraints (symmetrically for successors).
+We claim, however, that we can specify a labeling $\ell$ of the nodes of $H_{\mathrm{gap}}$ to indices of gap groups such that all pieces and enclosed components can be accommodated while observing all constraints.
+First label the nodes corresponding to gap groups with their number and label the nodes corresponding to pieces with the gap group determined by the 2-SAT solution.
+Consider the longest directed paths between two labeled nodes $p$ and $q$ with $\ell(p)\leq\ell(q)$ in $H_{\mathrm{gap}}$.
+Note that $\ell(q)-\ell(p)$ cannot be less than the length of a path between $p$ and $q$ because then we would have violated a constraint in the 2-SAT formula or there was no solution for $\mathcal{G}_{5}$ already.
+We next show a strategy how to iteratively label the nodes on such paths.
+We repeatedly determine the smallest $k\in\mathbb{N}_{0}$ such that there is a path that has labeled end nodes $p$ and $q$ (which can also be enclosed components) but unlabeled inner nodes and has length $\ell(q)-\ell(p)-k$ (i.e., a path with minimum “slack” for its labels).
+If such a path exists, we label its inner nodes with $\ell(p)+1,\dots,\ell(q)-k-1$.
+Note that this process terminates with a consistent labeling of pieces and enclosed components.
+Suppose for a contradiction that, in some step, there was an inner node $z$ on a path between $p$ and $q$ such that the label $\ell(z)$ became inconsistent with another path between $p^{\prime}$ and $q^{\prime}$ (where possibly $p=p^{\prime}$ or $q=q^{\prime}$); see Figure 19(c).
+Without loss of generality, assume that $\ell(z)-\ell(p^{\prime})$ is less than the length of the path between $p^{\prime}$ and $z$.
+As the path between $z$ and $q$ has length $\ell(q)-\ell(z)-k$, the path from $p^{\prime}$ via $z$ to $q$ has length at most $\ell(q)-\ell(p^{\prime})-k-1$ and we would have labeled this path first.
+Hence, $\mathcal{G}_{6}$ admits a constrained-level planar drawing, too.
+Concerning planarity note that pieces and enclosed components are assigned to gap groups and hence, by implicit constraints, they need to lie within a single gap each.
+Since gaps do not contain any edges in their corresponding band, crossings with the backbone are impossible.
+If two pieces are forced to have crossing edges they may not have the same anchor because this would violate the 2-SAT formula $\Phi$.
+Otherwise, and in the cases of one piece and one enclosed components or two enclosed component, any pair of edges that is forced to cross implies a constraint cycle in $\mathcal{G}_{5}$.
+Regarding the running time, note that we can determine and index all gaps and gap groups in linear time.
+Then, for each enclosed component and each piece $X$, we can find its gap group interval in linear time by taking into account the constraints between $X$ and the backbone.
+Overall, this takes quadratic time.
+To construct $H_{\mathrm{gap}}$, it remains to add arcs between non-gap-group nodes.
+For each pair of such nodes on different bands, we check whether there is a constraint between them.
+Again, this can be accomplished in quadratic time.
+The 2-SAT formula has at most a linear number of variables (corresponding to the pieces) and, thus, a quadratic number of clauses.
+To set up these clauses, we determine, for each piece, the longest paths in $H_{\mathrm{gap}}$ to all other pieces.
+Overall, this takes quadratic time.
+Finding a satisfying truth assignment of the 2-SAT formula takes time linear in its length [14], hence, quadratic time in our case.
+∎
+Note that the pieces that we have assigned to gap groups in the previous step have also been assigned to gaps and the orientations of their spines [^3] Recall that the spine of a caterpillar is the subgraph induced by the set of non-leaves.
+have been fixed – both due to implicit constraints.
+Now consider the other pieces.
+Recall that, due to Observation 7, no three pieces share an anchor.
+If two pieces share an anchor and their placement to the left or right of this anchor is not restricted by constraints, then we arbitrarily assign them to distinct sides.
+This assignment is fixed by adding constraints from or to their anchor.
+Again, this fixes the orientation of their spines.
+For each remaining piece that is not restricted to a single gap due to constraints, we arbitrarily assign it to one of its two possible gaps.
+For each remaining piece where the orientation of its spine is not yet fixed due to constraints, we arbitrarily fix the orientation of its spine – unless two pieces are incident to the same anchor and need to lie in the same gap.
+In this case we fix their orientation such that one extends to the left and the other extends to the right; see the two pieces with anchor $m_{14}$ in Figure 17.
+We call the resulting constrained level graph $\mathcal{G}_{7}$.
+Lemma 23.
+The constrained level graph $\mathcal{G}_{7}$ is constrained level planar if and only if $\mathcal{G}_{6}$ is.
+Furthermore, $\mathcal{G}_{7}$ can be constructed from $\mathcal{G}_{6}$ in quadratic time.
+Proof.
+If $\mathcal{G}_{6}$ is not constrained level planar, then certainly $\mathcal{G}_{7}$ is not constrained level planar either (because it is the same constrained level graph but with additional constraints).
+Otherwise assume that $\mathcal{G}_{6}$ is constrained level planar but $\mathcal{G}_{7}$ is not.
+First suppose that by assigning a piece $p$ to a specific gap $g$, the resulting constrained level planar graph does not admit a constrained level planar drawing any more.
+Note that then, in any constrained level planar drawing of $\mathcal{G}_{6}$, $p$ is assigned to the the gap on the other side of its anchor.
+If $p$ cannot be drawn inside of $g$ without crossing edges of the main component, then there would have been an implicit constrained forcing $p$ to be drawn in the other gap.
+Further, observe that $p$ can be drawn arbitrarily close to a backbone edge incident to $p$’s anchor because otherwise there would be a chain of constraints from $p$ to a backbone vertex to $p$’s anchor (or a neighbor of its anchor) or the same chain in the other direction.
+Then, however, there would have been no choice for the gap of $p$.
+If a vertex not in $p$ needs to be drawn between the vertices of $p$ due to constraints, then the orientation of $p$ is already fixed in $\mathcal{G}_{6}$.
+Otherwise, $p$ can be drawn arbitrary small, and then its orientation does not affect the remaining drawing.
+It remains to argue about the running time.
+We can find all pieces in linear time and check whether they are fixed by considering the constraints of the involved vertices.
+This can be done in overall quadratic time.
+Fixing the gaps and orientations can again be done in quadratic time.
+∎
+Now let $H_{\mathcal{G}_{7}}^{\prime}$ be the component–constraint graph of $\mathcal{G}_{7}$ without the main component.
+For both the upper and the lower band, there is a sequence of gaps where the enclosed components can be placed.
+Let the gap interval of an enclosed component $C$ be the subsequence of gaps within its band where $C$ can be placed according to its constraints.
+We go through the enclosed components in some topological order of $H_{G_{7}}^{\prime}$.
+We assign each component $C$ (by adding constraints) into the leftmost gap of its gap interval such that $C$ respects the constraints from components placed before.
+If a component cannot be assigned to a gap in its gap interval due to these constraints, $\mathcal{G}_{7}$ is not constrained level planar.
+We call the resulting constrained level graph $\mathcal{G}_{8}$.
+Lemma 24.
+The constrained level graph $\mathcal{G}_{8}$ is constrained level planar if and only if $\mathcal{G}_{7}$ is.
+Furthermore, $\mathcal{G}_{8}$ can be constructed from $\mathcal{G}_{7}$ in quadratic time.
+Proof.
+If $\mathcal{G}_{7}$ is not constrained level planar, then $\mathcal{G}_{8}$ is certainly not constrained level planar since the underlying graphs are identical, but the set of constraints of $\mathcal{G}_{8}$ is a superset of that of $\mathcal{G}_{7}$.
+Therefore, we can assume that $\mathcal{G}_{7}$ is constrained level planar.
+Let $\Gamma_{7}$ be a constrained level planar drawing of $\mathcal{G}_{7}$.
+Note that the backbones of $\mathcal{G}_{7}$ and $\mathcal{G}_{8}$ and hence their sets of gaps are identical.
+Let $\sigma$ be the topological order of $H_{\mathcal{G}_{7}}^{\prime}$ which we used to assign the enclosed components to gaps when defining $\mathcal{G}_{8}$.
+Let $C$ be the first enclosed component in $\sigma$ such that $C$ is assigned to a gap $g_{8}$ in $\mathcal{G}_{8}$, whereas $C$ lies in a different gap $g_{7}$ in $\Gamma_{7}$.
+By construction, the gap $g_{8}$ is to the left of the gap $g_{7}$.
+We can construct a new constrained level graph $\mathcal{G}_{7}^{(1)}$ from $\mathcal{G}_{7}$ by fixing $C$ to the gap $g_{8}$ with constraints.
+We claim that the graph $\mathcal{G}_{7}^{(1)}$ is constrained level planar.
+Note that the only difference between $\mathcal{G}_{7}^{(1)}$ and $\mathcal{G}_{7}$ is that $\mathcal{G}_{7}^{(1)}$ fixes $C$ to a specific gap $g_{8}$.
+By the choice of $g_{8}$, all constraints of type $a\prec_{i}c$ that involve a vertex $a\not\in C$ and a vertex $c\in C$ are satisfied in $\mathcal{G}_{7}^{(1)}$.
+On the other hand, assigning $C$ to $g_{8}$ restricts the placement of components that succeed $C$ in $\sigma$ less than the assignment of $C$ to $g_{7}$ in $\Gamma_{7}$.
+Note that there is always space at the right side of $g_{7}$.
+By repeating the above exchange argument, we show that there is a positive integer $k$ such that $\mathcal{G}_{8}=\mathcal{G}_{7}^{(k)}$ and, hence, $\mathcal{G}_{8}$ is constrained level planar due to the fact that the graphs $\mathcal{G}_{7}^{(k-1)},\dots,\mathcal{G}_{7}^{(1)}$ are all constrained level planar.
+We can determine the gap intervals of all enclosed components in quadratic total time because we have at most a quadratic number of constraints in $\mathcal{G}_{7}$.
+Furthermore, we can construct $H_{\mathcal{G}_{7}}^{\prime}$ in quadratic time (there can be a linear number of enclosed components in $\mathcal{G}_{7}$ and a quadratic number of constraints between them).
+We can sort $H_{\mathcal{G}_{7}}^{\prime}$ topologically in time linear in its size, that is, in at most quadratic time.
+∎
+(a) Constrained level planar drawing of $g$ with middle-level vertices $v_{1},\dots,v_{13}$ and its enclosed components.
+Vertices of $A_{g}$ are marked by (green) dotted ovals.
+Constraints are indicated by blue arcs (transitive constraints are omitted).
+(b) The directed auxiliary graph $A_{g}$ has a node for every enclosed component or piece and every non-covered middle-level vertex (transitive edges are omitted).
+(c) Vertex $b$ of the main component (red) is covered by an enclosed component or a piece.
+Figure 20: Arranging the enclosed components $C_{1},\dots,C_{5}$ in a gap $g$ of the main component.
+Constraints between vertices of $\mathcal{G}_{8}$ are indicated by blue arrows; arcs between nodes of $A_{g}$ are indicated by green arrows.
+For clarity, implicit constraints, constraints along the backbone, and transitive arcs are omitted.
+It remains to find a total order of all enclosed components, pieces, and middle-level vertices of the backbone within each gap $g$, and to fix the orientation of each enclosed component.
+If there are constraints of the type $a\prec_{2}b\prec_{2}c$, for a vertex $b$ of the backbone and vertices $a,c$ in an enclosed component or a piece $C$ ($a,b,c$ on the middle level), then $b$ lies in the span of $C$ in every constrained level planar drawing of $\mathcal{G}_{8}$.
+We say that $b$ is covered (by $C$); see Figure 20(c).
+For every gap $g$, we define a directed auxiliary graph $A_{g}$ that has a node for every enclosed component and every piece assigned to $g$ via constraints and for every middle-level vertex in $g$ that is not covered; see Figure 20.
+Note that these middle-level vertices are totally ordered due to Observation 6.
+In $A_{g}$, two nodes $x$ and $y$ are connected by an arc $(x,y)$ if there is a constraint between two vertices $v_{x}$ and $v_{y}$ in $\mathcal{G}_{8}$ that are represented or covered by $x$ and $y$, respectively.
+We sort $A_{g}$ topologically and, for each pair $(x,y)$ of consecutive nodes, we add a constraint for each pair $(v_{x},v_{y})$ of vertices on the same level with the property that $v_{x}$ is represented or covered by $x$ and $v_{y}$ is represented or covered by $y$.
+This results in a total order of the enclosed components, pieces, and uncovered middle-level vertices of the main component.
+Now, for each enclosed component, we fix the orientation of its spine by testing both possibilities.
+(If both orientations are possible, we keep an arbitrary one.) Note that this choice can be made independently for each enclosed component.
+We call the resulting constrained level graph $\mathcal{G}_{9}$.
+Lemma 25.
+The constrained level graph $\mathcal{G}_{9}$ is constrained level planar if and only if $\mathcal{G}_{8}$ is.
+Furthermore, $\mathcal{G}_{9}$ can be constructed from $\mathcal{G}_{8}$ in quadratic time.
+Proof.
+If $\mathcal{G}_{8}$ is not constrained level planar, then $\mathcal{G}_{9}$ is certainly not constrained level planar since the underlying graphs are identical, but the set of constraints of $\mathcal{G}_{9}$ is a superset of that of $\mathcal{G}_{8}$.
+Otherwise, assume that $\mathcal{G}_{8}$ is constrained level planar but $\mathcal{G}_{9}$ is not.
+If we add the new constraints (in an arbitrary order) one by one to $\mathcal{G}_{8}$, then there is a constraint $u\prec_{i}v$ (for some $i\in[3]$) such that adding $u\prec_{i}v$ makes the resulting constrained level graph not constrained level planar.
+Note that $u$ and $v$ cannot both be vertices of the backbone because the backbone vertices are already totally ordered in $\mathcal{G}_{8}$.
+Ler $x$ be the node in $A_{G}$ to which $u$ belongs and let $y$ be the node of $A_{g}$ to which $v$ belongs.
+We consider two cases.
+Case I: $x=y$.
+This means that $x$ is an enclosed component whose orientation has been fixed now because there was no restriction due to constraints.
+This choice of the orientation does not influence the rest of the constrained level graph, so in any constrained level planar drawing of $\mathcal{G}_{8}$, we can flip the orientation and still have a constrained level planar drawing.
+Case II: $x\neq y$. There are three subcases depending on what $x$ and $y$ represent.
+First, we assume that $x$ and $y$ both represent enclosed components or pieces.
+Clearly, there is no arc between $x$ and $y$ in $A_{g}$ (because no constrained level planar drawing of $\mathcal{G}_{8}$ contradicts $A_{g}$ and $\mathcal{G}_{9}$ is based on a topological sorting of $A_{g}$) and the vertices of $x$ and $y$ do not interleave because they share a common band.
+We claim that we can modify any constrained level planar drawing $\Gamma$ of $\mathcal{G}_{8}$ where $y$ precedes $x$ such that $x$ precedes $y$ and the drawing is still constrained level planar.
+Note that at most one of $x$ and $y$ has a covered vertex because otherwise there would be an arc in $A_{g}$.
+Say $y$ may cover some vertices and let $W$ be the set of vertices that are not part of $x$ or $y$ but have a constraint to a vertex of $x$ and lie, within $\Gamma$, between the first vertex of $y$ and the last vertex of $x$.
+Clearly, each node of $A_{g}$ containing a vertex from $W$ has an arc to $x$, but then it does not have an arc from $y$ (as otherwise, there would be an arc from $y$ to $x$ due to transitivity).
+Hence, in $\Gamma$, we can move the vertices in $W$ and $x$ to the left of the vertices of $y$ (while maintaining their internal order) and still have a constrained level planar drawing.
+So, $x$ and $y$ cannot both be enclosed components or pieces.
+Second, we assume that $x$ represents an enclosed component or a piece and that the vertex $v$ (represented by $y$) is a middle-level vertex of the backbone.
+Clearly, $x$ cannot cover a vertex then.
+Consider any constrained level planar drawing $\Gamma$ of $\mathcal{G}_{8}$.
+If $v$ lies in $\operatorname{span}_{\Gamma}(x)$, we can move $v$ immediately to the right of the rightmost vertex of $\operatorname{span}_{\Gamma}(x)$ because there is no constraint between vertices in $x$ and $v$.
+If $v$ lies to the left of all vertices from $x$, we again define $W$ as the set of vertices that are not part of $x$ or $y$ but have a constraint to a vertex of $x$ and lie, within $\Gamma$, between $v$ and the last vertex of $x$.
+By the same argument as before, we can move all vertices in $W$ and $x$ to the left of $v$ and still have a constrained level planar drawing.
+Third, we assume that $y$ represents an enclosed component or a piece and that the vertex $u$ (represented by $x$) is a middle-level vertex of the backbone.
+This subcase is symmetric to the second subcase.
+This finishes the case distinction. In every case we have obtained the desired contradiction.
+It remains to analyze the running time for constructing $\mathcal{G}_{9}$.
+Detecting all covered vertices can be done in at most quadratic time: for each middle-level vertex $b$, first iterate over each incoming constraint $a\prec_{2}b$ and mark the enclosed component or piece that $a$ is part of, afterwards iterate over each outgoing constraint $b\prec_{2}c$ and check whether the enclosed component or piece that $c$ is part of is marked.
+Constructing the auxiliary graphs can then be done in at most quadratic time over all gaps.
+Each auxiliary graph can be sorted topologically in time proportional to its size.
+In total, this takes at most quadratic time.
+Adding the new constraints and fixing the orientation of each enclosed component can both again be done in at most quadratic time.
+∎
+Now we consider the remaining leaves and group them by their neighbors.
+For each such group, we find a topological order and fix it.
+If the group lies on the middle level, then we need to include the backbone vertices in the topological order.
+Now let $\mathcal{G}_{10}$ be the constrained level graph obtained from $\mathcal{G}_{9}$ by fixing the orders of every group of leaves.
+Lemma 26.
+The constrained level graph $\mathcal{G}_{10}$ is constrained level planar if and only if $\mathcal{G}_{9}$ is.
+Furthermore, $\mathcal{G}_{10}$ can be constructed from $\mathcal{G}_{9}$ in linear time.
+Proof.
+If there is a constrained level planar drawing $\Gamma_{10}$ of $\mathcal{G}_{10}$, then it is clearly also a constrained level planar drawing of $\mathcal{G}_{9}$.
+If there is a constrained level planar drawing $\Gamma_{9}$ of $\mathcal{G}_{9}$, we can obtain a constrained level planar drawing of $\mathcal{G}_{10}$ by reordering the leaves of every group according to the topological order they have in $\mathcal{G}_{10}$.
+If this violates a constraint in $\mathcal{G}_{10}$, that constraint must have been in $\mathcal{G}_{9}$ already.
+We claim that groups on the same level are independent.
+Consider the unique neighbors of two such groups.
+Recall that these neighbors are not part of the backbone.
+If both neighbors lie on the same level, the spans of their groups cannot overlap due to planarity constraints.
+If one lies on the top level, the other on the bottom level, their groups are separated by the backbone.
+In other words, if the spans of the two groups overlapped, the backbone would be blocked.
+Finding and grouping all leaves can be done in linear time by going through the graph once. Sorting all groups topologically can be done in linear total time. ∎
+Now, we have added sufficiently many constraints such that, on each level, the corresponding vertices of $\mathcal{G}_{10}$ are totally ordered.
+This yields a polynomial-time algorithm to check whether $\mathcal{G}$ is constrained level planar and, if yes, to compute a constrained level planar drawing of $\mathcal{G}$.
+Our estimation of the running time is most likely not tight.
+Theorem 8.
+$3$-level CLP can be decided in $\mathcal{O}(n^{5})$ time, where $n$ is the number of vertices in the input graph.
+More precisely, given a 3-level constrained level graph $\mathcal{G}$, we can test in $\mathcal{O}(n^{5})$ time whether $\mathcal{G}$ is constrained level planar and, if yes, we can compute a constrained level planar drawing of $\mathcal{G}$ within the same time bound.
+Proof.
+According to the Lemmas 16 to 26, $\mathcal{G}$ is constrained level planar if and only if $\mathcal{G}_{10}$ is.
+We next argue that the constraints of $\mathcal{G}_{10}$ are total orders for all three levels.
+This follows from the following facts.
+  • There are no isolated vertices. (Lemma 7)
+  • There is only one component (the main component) that spans three levels; the other components are enclosed components on two levels. (Lemmas 16 and 20)
+  • The main component consists of the backbone and pieces, which are caterpillars on two levels. (Lemmas 17 and 20)
+  • The backbone vertices are totally ordered. (Observation 6)
+  • There are no leaves adjacent to the backbone and the remaining leaves (including potentially $s$ an $t$) are totally ordered. (Lemmas 21 and 26)
+  • The vertices in every enclosed component and in every piece are totally ordered. (Lemmas 23 and 25)
+  • The backbone vertices, pieces, and enclosed components are totally ordered. (Lemma 25)
+Therefore, for each $i\in[3]$, the constraints of $\mathcal{G}_{10}$ for level $i$ give a total order of the vertices.
+As $\mathcal{G}_{10}$ is proper, we can check in quadratic total time, for each pair of edges, whether it induces a crossing.
+The graph $\mathcal{G}$ is constrained level planar if and only if no pair of edges induces a crossing.
+For computing a constrained level planar drawing of $\mathcal{G}$ in the affirmative case, start with the constrained level planar drawing of $\mathcal{G}_{10}$ and reverse the operations done before to obtain constrained level planar drawings of $\mathcal{G}_{9},\dots,\mathcal{G}_{1},\mathcal{G}$.
+Note that all operations can straight-forwardly be reversed in at most the same asymptotic running time as before.
+It remains to bound the running time.
+We first guess the first and last vertex on the middle level $s$ and $t$ and test all $\mathcal{O}(n^{2})$ pairs.
+This guess yields the first and the last component of a potential hook chain (see Lemma 14).
+For this pair of components, we test four sets of vertices as hook anchors (see Lemma 15); this does not influence the asymptotic running time.
+Given the hook chain and the hook anchors $s$ and $t$, we connect the hook chain in linear time (see Lemma 16).
+Then we add, in linear time, constraints that fix $s$ and $t$ as the first and the last vertex on the middle level, which implicitly fixes the orientation of the backbone (see Lemma 18).
+For each guess of $s$ and $t$, we construct, starting with $\mathcal{G}_{2}$, the sequence $\mathcal{G}_{3},\mathcal{G}_{4},\dots,\mathcal{G}_{10}$, which can be done in $\mathcal{O}(n^{2})$ time each (see Lemmas 19, 20, 21, 22, 23, 24, 25 and 26).
+It remains to consider the running time of checking for and adding implicit constraints (see the beginning of Section 6).
+To this end, note that (i) each of the $O(n^{2})$ constraints can be added at most once and (ii) every time a constraint is added, (ii-a) at most $O(n)$ many other existing constraints must be checked for transitivity constraints and (ii-b) at most $O(n^{2})$ combinations of neighbors must be checked for planarity constraints.
+However, for each pair of edges $ab$ and $cd$ on the same band, note that the constraint $b\prec d$ is only checked if the constraint $a\prec c$ is added, and vice versa.
+Hence, there is only a total of $O(n^{2})$ checks for planarity constraints.
+In summary, checking for and adding implicit constraints takes $\mathcal{O}(n^{3})$ total time over all steps.
+Hence, our algorithm runs in $\mathcal{O}(n^{5})$ total time. ∎
+7 Open Problems and Further Remarks
+First, we propose the following two open problems.
+  1. 1.
+    We have shown that 3-level CLP can be solved in $\mathcal{O}(n^{5})$ time (Theorem 8).
+Our approach has two bottlenecks.
+First, we do not know how to add all implicit constraints (in particular the implicit transitivity constraints) in $o(n^{3})$ time after each step of the algorithm.
+Note that after each step, we could run the algorithm for computing the transitive closure in the same running time as matrix multiplication [16, 29], for which the currently best known algorithm has a running time in $\mathcal{O}(n^{2.37216})$ [32].
+However, if the constraints added by this algorithm lead to new planarity constraints, which in turn imply further transitivity constraints, this could trigger a chain of dependences, which we do not know how to handle in overall $o(n^{3})$ time.
+Second, our algorithm relies on guessing the first and last vertex of the backbone, which means that a quadratic number of guesses need to be checked, each in currently $\mathcal{O}(n^{3})$ time.
+Can we instead directly construct the backbone or guess just one of the two vertices and compute the other one?
+  2. 2.
+    Consider the problem Connected CLP, which is the special case of CLP where the given constrained level graph must be connected.
+It is easy to make the graph in our NP-hardness proof for 4-level CLP connected by adding two new levels, so 6-level Connected CLP is NP-hard.
+What is the complexity of 5-level Connected CLP?
+Finally, we remark that a reviewer pointed us to [12], whose authors study problems where the input is an undirected graph without a level assignment and the task is to find a crossing-free drawing with y-monotone edges that, if interpreted as a level planar drawing, satisfies certain criteria, e.g., being proper and having height at most $h$.
+At the end, they claim an FPT-algorithm for problem variants “where some vertices have been preassigned to layers and some vertices must respect a given partial order” [12, Item 5 in Theorem 5].
+As suggested by the reviewer, this variant appears to be CLP, which would contradict our hardness results.
+However, the statement of [12, Theorem 5] is rather unclear and comes without proof or further explanation.
+In conversation with the authors, we confirmed that they never considered horizontal constraints (“$u$ is to the right of $v$”) and, so, their claim is not related to our results.
+We presume that they refer to vertical constraints (“$u$ is above $v$”) instead.